### v1

# Rethinking Knowledge Distillation: A Data Dependent Regulariser With a Negative Asymmetric Payoff

Abstract
Knowledge distillation is often considered a compression mechanism when judged on the resulting student’s accuracy and loss, yet its functional impact is poorly understood. In this work, we quantify the compression capacity of knowledge distillation and the resulting knowledge transfer from a functional perspective, decoupling compression from architectural reduction, which provides an improved understanding of knowledge distillation. We employ hypothesis testing, controls, and random control distillation to understand knowledge transfer mechanisms across data modalities. To rigorously test the breadth and limits of our analyses, we explore multiple distillation variants and analyse distillation scaling laws across model sizes. Our findings demonstrate that, while there is statistically significant knowledge transfer in some modalities and architectures, the extent of this transfer is less pronounced than anticipated, even under conditions designed to maximise knowledge sharing. Notably, in cases of significant knowledge transfer, we identify a consistent and severe asymmetric transfer of negative knowledge to the student, raising safety concerns in knowledge distillation applications. Across 12 experimental setups, 9 architectures, and 7 datasets, our findings show that knowledge distillation functions less as a compression mechanism and more as a data-dependent regulariser with a negative asymmetric payoff.
1 Introduction
Large neural networks have achieved remarkable results across domains (Brown et al., 2020; Dosovitskiy et al., 2021; Kirillov et al., 2023), but at significant computational cost. This has motivated techniques that reduce model size while maintaining performance. Knowledge distillation (KD) has emerged as a widely adopted method to compress models by training a student model to mimic a larger teacher (Buciluǎ et al., 2006; Hinton et al., 2015; Gu et al., 2024; Muralidharan et al., 2024). While KD can be applied across architectures and modalities – including in self-distillation regimes where the teacher and student share the same architecture (Allen-Zhu & Li, 2023; Zhang et al., 2019) – the mechanism by which KD improves student performance remains unknown (Busbridge et al., 2025). Recent studies have challenged the assumption that KD works through meaningful knowledge transfer, showing that performance gains have been observed even with randomly initialised teachers (Stanton et al., 2021a) motivating a rigorous examination of KD’s functional impact.
In this work, we move beyond the question of whether knowledge is transferred – we challenge the framing of Knowledge Distillation as a mechanism of knowledge transfer altogether. We argue that the improvements observed do not necessarily arise from meaningful transfer of the teacher’s knowledge, but from a more general, data-dependent regularisation effect disputed in literature (Stanton et al., 2021a; Yun et al., 2020; Ge et al., 2021; Yuan et al., 2020) with a novel identification of a negative asymmetric payoff in KD. To support this claim, we study KD from a functional perspective, and quantify how closely student models replicate the teacher’s output function. We ground our work around two research questions: 1) Does knowledge distillation result in a significantly functionally similar model to the teacher across architectures and data domains against controls? 2) What knowledge, if any, is actually transferred to student models?
We first focus on self-distillation, where the student has the capacity to match the teacher’s functional representation perfectly, ensuring that any observed differences are solely due to the distillation signal. We then verify our findings in the standard distillation setting with smaller student models (Appendix Section E), as well as with different KD variants in Appendix Section C.
Our methodological framework isolates the core mechanics of Knowledge Distillation through: 1) a controlled training setup where all models share initialisation, enabling precise functional comparison; 2) two controls: independent models with the same architecture,initialisation and different data order (SIDDO) as the teacher, and a Random Control Distillation (RCD) where students are trained using uniform noise in place of teacher outputs, all functionally compared to the teacher model used in the standard distillation process; 3) functional similarity metrics including Activation Distance, Rank Disagreement, Prediction Disagreement, JS Divergence and Prediction Agreement.
We conduct experiments across 7 datasets, 3 data modalities (image, audio, and language), and 9 architectures, training over 3,900 models. Our findings show that:
While KD can lead to statistically significant functional similarity between teacher and student, this similarity is often marginal and inconsistent across datasets and modalities.
The most substantial improvements in accuracy and loss frequently arise under Random Control Distillation, challenging the assumption that performance gains reflect successful knowledge transfer.
When knowledge transfer is significant and not marginal, the transferred knowledge has an asymmetric weighting towards the teacher’s incorrect predictions. This asymmetry becomes more pronounced as dependence on the teacher increases.
Our findings compel a re-characterisation of KD, not as a robust knowledge transfer mechanism, but as a data-dependent regulariser with inconsistent and negative asymmetric knowledge-sharing capacity. This perspective raises important safety concerns: when knowledge transfer is significant, KD may amplify incorrect or harmful behaviour encoded in the teacher. We present a concrete case of adversarial transfer facilitated by KD to support this.
Concretely, our contributions are as follows:
Introduce a functional framework to analyse KD beyond accuracy and loss, but as a process where internal knowledge transfer dynamics can be quantitatively measured.
Isolate the contribution of the teacher signal using strong statistical and control-based methodology, something that prior work has not quantitatively disentangled to this level.
Identify and characterise a novel phenomenon across conditions, modalities and architectures: when functional transfer occurs, it disproportionately favours the teacher’s incorrect predictions, revealing a systematic error amplification effect with safety implications.
Demonstrate the diagnostic utility of RCD as a crucial counterfactual, showing it frequently outperforms KD, undermining assumptions about knowledge transfer.
Conduct the largest multimodal functional study of KD to date. Our empirical analysis spans over 3,900 trained models across 9 architectures, 7 datasets, and 3 modalities (vision, audio, and language), establishing the generality and reproducibility of our claims.
Reveal targeted and scalable negative transfer via adversarial and capacity scaling experiments. We show that KD can reliably copy specific erroneous behaviours, and that this error amplification scales with model capacity, underscoring the hidden risks of KD in high-stakes settings.
2 Related Work
Knowledge Distillation (KD):
KD transfers behaviour from a teacher (or ensemble) into a student (Buciluǎ et al., 2006; Hinton et al., 2015), with strong empirical results across modalities (Beyer et al., 2022; Jung et al., 2020; Sanh, 2019; Aghli & Ribeiro, 2021; Li et al., 2020; Fang et al., 2021; Wang et al., 2022) and architectures (Touvron et al., 2021; Miles et al., 2024). Yet the role of knowledge transfer is debated (Mason-Williams, 2024; Stanton et al., 2021b; Ojha et al., 2023; Menon et al., 2021). Prior work alternately views KD as a regulariser (Yun et al., 2020; Ge et al., 2021; Yuan et al., 2020) or argues against that view (Shen et al., 2021; Sultan, 2023). In this paper, we advance the discussion surrounding KD as a regulariser with a functional perspective that spans image, audio, and language. We present a control-driven functional protocol that decouples compression from size, measures alignment beyond accuracy, confirming KD acts as a data-dependent regulariser but exposing a new dimension of this regularisation with respect to its systematic negative transfer to the student.
Functional Similarity Metrics:
Functional similarity compares models by their outputs rather than only their accuracy (Klabunde et al., 2023). It has been used for unlearning (Golatkar et al., 2021; Chundawat et al., 2023), ensemble dynamics (Fort et al., 2019), and compression/pruning (Mason-Williams & Dahlqvist, 2024; Mason-Williams, 2024). Metrics such as Activation Distance, Prediction Dissimilarity and JS Divergence have been used for functional analysis. Activation Distance represents the distance on the softmax output distribution of two models, enabling functional comparison. In comparison, JS Divergence represents the Jensen-Shannon information-theoretic divergence that employs a weighted average of KL divergence of distributions, giving a directed divergence between non-continuous distributions (Lin, 1991). Prediction Dissimilarity compares the disagreement of label predictions between models, allowing for an enriched perspective on the alignment of the model’s functions (Fort et al., 2019). We employ all of the above to conduct a functional analysis of knowledge transfer in knowledge distillation.
3 Experimental Setup
We focus primarily on self-distillation, where the student model has the same architecture and initialisation as the teacher. This setting gives the student maximal capacity to recover the teacher’s function, allowing isolation of the effects of the distillation signal itself. This is achieved through architectural and initialisation matching, along with carefully structured control conditions. Our core experimental findings are derived from this controlled self-distillation setup. To verify generality, we replicate our results in the standard KD setting with smaller students (Appendix E) as well as with multiple KD variants in Appendix Section C.
Let denote the teacher model, trained from initialisation . All subsequent models – including students and controls – share the same architecture and initialisation , ensuring they begin from the same point in the loss landscape. Thus any observed differences in functional behaviour arise purely from the training signals (e.g., data order or distillation) rather than confounds from architecture or initialisation. In self-distillation, students start from and are trained to match the finalised teacher with the standard logit-matching objective:
where is the input, is the student model parameters, is the teacher weighting coefficient, is the cross-entropy loss function, is the kullback-leibler divergence loss function, is the ground truth label, is the softmax function parameterised by the temperature , and and are the logits of the student and the teacher, respectively. Unless otherwise stated, we keep all training hyperparameters fixed across conditions: optimiser, learning-rate schedule, batch size, data augmentations/preprocessing, epochs, and evaluation protocol.
To isolate the effect of the teacher signal, we introduce a Random Control Distillation (RCD) setup, analogous to a randomised control trial (Hariton & Locascio, 2018). Here, the student is trained with the same distillation loss (Eq. 1), but the teacher outputs are replaced by samples from a uniform distribution in . This setup is visualised in Figure 2(a).
We vary the distillation coefficient to modulate reliance on the teacher. At 0.1, the teacher signal contributes minimally; at 0.5, there is an equal weighting of label and teacher supervision; at 0.9, training is predominantly guided by the teacher. If KD achieves meaningful knowledge transfer, functional similarity should increase with higher . All experiments use temperature to preserve the original teacher distribution.
For each architecture–dataset pair spanning over different modalities, we train 3 teacher models (seeds 0-2), and 10 student models per distillation setup (KD, RCD, SIDDO; see below) 3 values (seeds 10-19). This results in 73 models per dataset–architecture pair, and a total of 3,942 models across all conditions (Table 1). Results are reported using Standard Error of the Mean (SEM) (Belia et al., 2005), which better reflects estimation uncertainty across independent runs. Our experimental setup is aligned with reproducibility protocols, defined by Mason-Williams & Mason-Williams (2025), using hypothesis testing and levering statistical significance to enhance the robustness and generality of our findings.
| Modality | Datasets | Architectures | ||
| Image | TinyImageNet Le & Yang (2015), CIFAR10 Krizhevsky et al. (2009), SVHN Netzer et al. (2011) |
| Audio |
VGGish Hershey et al. (2017), AST Gong et al. (2021) | ||
| Language | Tiny Shakespeare Blog (2015), Adversarial Tiny Shakespeare (THA) | Nano-GPT, Pico-GPT Karpathy (2022) |
3.1 Functional Similarity Metrics
We evaluate student–teacher alignment using functional similarity metrics computed on the test set , comparing teacher and comparison model :
Activation distance: distance between softmax outputs of and .
Rank Disagreement: Percentage of disagreement in the sorted output logits.
Prediction Disagreement: Proportion of mismatched top-1 predictions..
Prediction Agreement: Complement of prediction disagreement (used in error analysis).
Jensen-Shannon (JS) Divergence: A weighted average of KL divergence (Lin, 1991) between the softmax outputs of and .
These metrics move beyond accuracy and loss to quantify the extent to which students reproduce the teacher’s output function.
3.2 Hypothesis Testing
To evaluate whether KD facilitates functional knowledge transfer, we test whether student models trained via KD are functionally more similar to the teacher than control models. Our primary hypothesis is:
KD students, on average, are no more similar to the teacher than control models.
KD students, on average, are more functionally similar to the teacher than control models.
We test each functional similarity metric using a two-sided Mann-Whitney U test (significance level = 0.05). Comparisons are made between two control conditions and the variable of interest:
Same Initialisation Different Data Order (SIDDO): models with the same initialisation and architecture as the teacher, trained with seeds 10-19.
Random Control Distillation (RCD): Students trained with uniform-noise “teacher” logits (seeds 10-19; alphas 0.1, 0.5 and 0.9) (Figure 2(a)).
Standard KD (variable of interest): Students trained with real teacher logits from , using alpha values and seeds 10–19 (Figure 2(a)).
For each teacher seed, we report the mean and SEM across 10 models per condition.
4 Results and Discussion
We first examine functional transfer in small-scale settings and show that when transfer is non-marginal it is consistently asymmetric toward the teacher’s errors. We then validate these findings at larger scale on TinyImageNet, where increasing teacher train loss (via augmentation) amplifies both functional transfer and its negative asymmetry. We then demonstrate generality in negative asymmetric transfer of KD across modalities (audio and language in addition to image), show how KD can facilitate adversarial attacks and finally we provide distillation scaling experiment, in line with Busbridge et al. (2025), to show how negative asymmetric transfer is present regardless of student capacity.
Full supplemental results (datasets, architectures, and all teacher seeds) appear in the appendix: CIFAR-10 (ResNet-18, VGG19, ViT; Appendix F.2), SVHN (VGG19, ViT; Appendix F.3), audio (UrbanSound8K, SpeechCommands; Appendix G), language (Tiny Shakespeare; Appendix H), adversarial transfer (Appendix H.2), standard KD to smaller students (Appendix E), and different KD variants (Appendix C). We also show in Appendix Section B that our analysis holds for information theoretic and geometric measures alongside our functional similarity measures and that our RCD control is equivalent to label smoothing in Appendix Section D. Training details for all settings are also provided in the appendix. Unless specified otherwise, we report means and 1 SEM over 10 runs per teacher seed and condition.
4.1 Function Transfer in Small-Scale Settings (SVHN)
We begin with SVHN and ResNet18. KD yields statistically significant functional similarity at high values, but the magnitude and asymmetry of transfer vary across teacher seeds. When transfer is non-marginal, we observe a systematic increase in student–teacher agreement on incorrect predictions relative to correct ones.
Table 2 shows teacher variability: train losses of , , and with a generalisation gaps of for seeds 0, 1, and 2 respectively. Notably, the best test loss and accuracy (Table 3) are achieved by random control distillation, reducing confidence that KD’s performance gains arise from meaningful knowledge transfer and instead supporting the view of KD as a data-dependent regulariser.
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.000646 | 0.999850 | 0.381410 | 0.951829 |
| 1 | 0.000061 | 0.999973 | 0.331054 | 0.952251 |
| 2 | 0.004657 | 0.998580 | 0.309702 | 0.947104 |
For the highest-train-loss teacher (seed 2), KD produces significant functional transfer across metrics at most values (Appendix Table 64; reproduced summary in Table 4), with the exception of Prediction Disagreement at . This transfer coincides with a large asymmetric payoff in prediction agreement toward the teacher’s incorrect predictions (Figure 5(a)). The lowest-train-loss teacher (seed 1) shows no significant transfer at and only partial transfer at (again, excluding Prediction Disagreement). Seed 0 (intermediate train loss) shows significant transfer at and , accompanied by asymmetric incorrect agreement (Figure 5(a)).
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.063 | 0.064 | 0.060 | 0.059 | 0.144 | 0.493 | 0.849 |
| Rank Disagreement () | 0.696 | 0.688 | 0.684 | 0.681 | 0.800 | 0.798 | 0.802 |
| Prediction Disagreement () | 0.045 | 0.046 | 0.043 | 0.042 | 0.042 | 0.043 | 0.046 |
| JS Divergence () | 0.025 | 0.025 | 0.023 | 0.022 | 0.053 | 0.201 | 0.431 |
| Accuracy () | 0.952 | 0.951 | 0.954 | 0.954 | 0.957 | 0.957 | 0.955 |
| Loss () | 0.385 | 0.344 | 0.310 | 0.293 | 0.236 | 0.692 | 1.698 |
| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
| KD 0.1 | ✗✗✓ | ✗✗✓ | ✗✗✗ | ✗✗✓ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✗✗✓ | ✓✗✓ | ✗✗✓ | ✓✗✓ | ✗✗✗ | ✗✗✓ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✗✗✓ | ✓✓✓ | ✗✗✗ | ✗✗✓ |
Across seeds, higher teacher train loss is associated with stronger (and more asymmetric) functional transfer. This is consistent with a teacher that deviates more from ground-truth labels, thereby exposing students to incorrect structure that is preferentially transferred under KD.
4.2 Function Transfer in Larger-Scale Settings
We next study TinyImageNet with ResNet50. In the base setting, KD produces significant but marginal functional gains relative to SIDDO; the corresponding prediction agreement shows no clear preference toward correct or incorrect agreement. Motivated by the SVHN analysis, we increase the teacher train loss via data augmentation (same training pipeline) – RandAugment (Cubuk et al., 2020) with the default settings – and examine the consequences for functional transfer and asymmetry.
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| Base | ||||
| 0 | 0.001426 | 0.999800 | 2.070590 | 0.605300 |
| 1 | 0.001393 | 0.999800 | 2.051494 | 0.607900 |
| 2 | 0.001436 | 0.999800 | 2.051024 | 0.610600 |
| RandAugment | ||||
| 0 | 0.672748 | 0.840410 | 1.620552 | 0.638800 |
| 1 | 0.678245 | 0.839200 | 1.629393 | 0.641800 |
| 2 | 0.667570 | 0.840750 | 1.624969 | 0.641100 |
In the base setting (Table 5), teachers have very low train loss and moderate test accuracy. With augmentation (Table 5), train loss increases while test accuracy improves, as expected.
Having established how augmentation changes the teacher regime, we now examine the students under the same settings (teacher seed 0). In the base case, KD with 0.9 improves over SIDDO by at most (Activation Distance), (Rank Disagreement), (Prediction Disagreement), and (JS Divergence) (Table 6) – statistically significant (Appendix Table 29) but marginal in magnitude. Under augmentation, KD with 0.9 improves by (Activation Distance), (Rank Disagreement), (Prediction Disagreement), and (JS Divergence) (Table 7). In both base and augmented settings, the best test loss/accuracy occurs under random control distillation, indicating that improved performance does not require a meaningful teacher signal.
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.157 0.001 | 0.157 0.001 | 0.156 0.001 | 0.155 0.000 | 0.343 0.000 | 0.581 0.000 | 0.791 0.000 |
| Rank Disagreement | 0.939 0.000 | 0.939 0.000 | 0.939 0.000 | 0.939 0.000 | 0.980 0.000 | 0.984 0.000 | 0.984 0.000 |
| Prediction Disagreement | 0.153 0.001 | 0.152 0.001 | 0.151 0.001 | 0.151 0.001 | 0.190 0.001 | 0.214 0.000 | 0.324 0.000 |
| JS Divergence | 0.040 0.000 | 0.040 0.000 | 0.039 0.000 | 0.039 0.000 | 0.171 0.000 | 0.333 0.000 | 0.533 0.000 |
| Accuracy | 0.605 0.001 | 0.605 0.000 | 0.604 0.001 | 0.605 0.001 | 0.607 0.000 | 0.606 0.001 | 0.580 0.000 |
| Loss | 2.068 0.001 | 2.065 0.002 | 2.055 0.001 | 2.043 0.002 | 1.977 0.001 | 2.497 0.001 | 3.612 0.002 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.193 0.000 | 0.183 0.000 | 0.150 0.000 | 0.131 0.000 | 0.245 0.001 | 0.501 0.001 | 0.781 0.000 |
| Rank Disagreement | 0.959 0.000 | 0.957 0.000 | 0.948 0.000 | 0.943 0.000 | 0.975 0.000 | 0.981 0.000 | 0.987 0.000 |
| Prediction Disagreement | 0.196 0.001 | 0.188 0.001 | 0.154 0.001 | 0.136 0.001 | 0.195 0.001 | 0.240 0.001 | 0.572 0.001 |
| JS Divergence | 0.058 0.000 | 0.052 0.000 | 0.036 0.000 | 0.028 0.000 | 0.094 0.000 | 0.266 0.000 | 0.563 0.000 |
| Accuracy | 0.640 0.000 | 0.643 0.001 | 0.644 0.000 | 0.642 0.000 | 0.646 0.001 | 0.657 0.001 | 0.400 0.001 |
| Loss | 1.619 0.003 | 1.600 0.001 | 1.578 0.001 | 1.577 0.001 | 1.551 0.001 | 1.984 0.002 | 4.211 0.001 |
Figure 11(a) shows the corresponding prediction agreement deltas (KD vs. best control). At , students trained from augmented teachers increase incorrect agreement from (base) to , far outpacing the increase in correct agreement. Thus, inducing higher teacher train loss via augmentation reliably amplifies asymmetric incorrect transfer, consistent with the SVHN findings and our regularisation view of KD with the novel insight of negative asymmetric transfer.
4.3 Function Transfer Across Modalities
We test the generality of our findings beyond images by evaluating KD on audio (UrbanSound8K, SpeechCommands) and language (Tiny Shakespeare). Across modalities, the same pattern holds: when transfer is non-marginal (per functional similarity metrics), it is asymmetric: students preferentially increase agreement with the teacher on incorrect predictions, and this imbalance strengthens as the teacher weight increases. Below we show the VGGish architecture on the audio datasets and the NanoGPT on Tiny Shakespeare.
In Figure 11(b), a clear pattern emerges: when there is considerable knowledge transfer, as evidenced by results across functional similarity metrics (Appendix Sections G and H), an asymmetric relationship becomes evident in the nature of the transfer. Specifically, student models receive significantly more transfer of the teacher model’s incorrect predictions than its correct predictions, with this imbalance scaling linearly as the weighting on the teacher outputs increases. These results highlight the generality of our understanding of knowledge distillation as a data-dependent regulariser with a negative asymmetric payoff.
4.4 Adversarial Transfer (Language): Targeted error copying
To move beyond aggregate functional similarity, we test whether KD copies a specific erroneous behaviour from its teacher. Informed by the Zipf’s Law distribution (Piantadosi, 2014) of the Tiny Shakespeare dataset as seen in Figure 11, we construct an adversarially biased Tiny Shakespeare teacher by editing its training corpus so that every instance of “the” is replaced with “tha”, a sequence that does not occur in the clean dataset (Appendix H.2, Table 100). This induces a stable bias to complete “th_” as “tha” rather than “the”, while the teacher’s overall performance on clean data remains comparable to standard models (Table 101). We then distil this teacher at and compare against our two controls (SIDDO and RCD) under identical training conditions.
| Control | Knowledge Distillation | Random Control Distillation | ||||||
| Predicted Word | Teacher | SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 |
| tha | 454 | 105.90 4.168 | 106.00 3.046 | 199.10 13.391 | 436.20 7.984 | 104.60 3.898 | 114.80 3.056 | 126.90 8.068 |
| the | 285 | 665.10 7.675 | 675.50 10.228 | 583.40 17.536 | 343.60 6.358 | 668.80 12.713 | 712.50 12.480 | 826.30 20.203 |
On clean evaluation prompts containing “th_”, we measure how often models complete to “tha” versus “the” and aggregate results per teacher seed, as seen for teacher seed 0 in Table 8 (with seeds 1-2 in Appendix Tables 103 and 104). KD, particularly at higher , markedly increases the rate of “tha” completions and suppresses “the” relative to both controls, demonstrating that KD can selectively copy a targeted error pattern even when overall behaviour appears benign. This experiment adds causal evidence that KD transmits specific erroneous structure, not merely broad functional alignment, sharpening the safety implication of our main findings: practitioners may unknowingly inherit unintended behaviours from the teacher, reinforcing our characterisation of KD as a data-dependent regulariser with a negative asymmetric payoff. Full details and per-seed statistics are provided in Appendix H.2.
4.5 Distillation Scaling Laws
The preceding sections established when KD transfers knowledge, this transfer is negatively asymmetric. We now ask how these effects evolve with capacity. Distillation Scaling Laws (DSL) (Busbridge et al., 2025) quantify how much student loss changes with compute, teacher quality, and model size. Our study complements DSL by asking how much is transferred as capacity grows: we decompose the distillation signal into correct vs. incorrect teacher–student agreement, offering a mechanistic reading of the “teacher quality” term and explaining negative-transfer regimes that are invisible from loss alone. Concretely, on Tiny Shakespeare we sweep student width from to in steps under a fixed-epoch budget matched to the teacher, using the same optimiser. For each width and , we measure the change in correct and incorrect agreement relative to the best control baseline (means SEM over 10 runs; teacher seed 0). In Figure 12 three core trends emerge which are descirbed below.
1) Student capacity helps, but mainly by amplifying the teacher’s mistakes: as width increases, both correct and incorrect agreement rise, yet the incorrect column grows much faster (from to width at , correct agreement vs. incorrect ).
2) Small students suffer negative transfer: at - width, the incorrect boost is comparable to or larger than the correct.
3) Increasing capacity unlocks more of the distillation signal: however what flows first, and most strongly, is the teacher’s error pattern.
Taken together, these scaling results reveal what is driving the loss curves: KD acts as a data-dependent regulariser with a negative asymmetric payoff, and scaling up the student amplifies the asymmetry of transfer.
5 Gradient-Level Explanation of Asymmetric Transfer
We now provide a concise theoretical explanation for the observed asymmetric error transfer in KD, and in Appendix B extend our functional analysis with information-theoretic and geometric perspectives to quantify when and how alignment with the teacher becomes harmful. These analyses clarify the risks of distillation, especially in safety-critical settings.
Consider the standard KD objective:
where and are the student and teacher logits, respectively. The per-logit gradient is:
with and .
When is the correct class (), the includes both supervision and teacher alignment. But when is an incorrect class (), the gradient reduces to:
This pulls the student toward any non-zero mass the teacher places on that incorrect class. The strength of this pull scales with and the teacher’s own loss. This simple derivation explains our central finding: when the teacher is imperfect, KD disproportionately transfers its errors to the student. The resulting alignment is asymmetric, favouring incorrect predictions. By contrast, if the teacher logits are replaced with a uniform distribution – as in label smoothing (Appendix D) or our random control distillation – the gradient on incorrect classes becomes flat, removing this error-amplifying signal. Empirically, these baselines match or exceed KD’s accuracy, while showing no rise in incorrect agreement. Overall we argue that the observed asymmetric transfer in KD is not incidental but rather emerges directly from the structure of the KD objective.
6 Conclusion
Across controlled self-distillation, small/large-scale settings, cross-modality (image, audio, language), a targeted error test, capacity scaling, standard KD setting with smaller students (Appendix E), and multiple KD variants (Appendix C), KD seldom delivers robust “knowledge transfer”. When transfer occurs, it is typically marginal and inconsistent, and increases with teacher imperfection, amplifying the teacher’s errors more than its correct behaviour (negative asymmetry). By contrast, Random Control Distillation often yields the best loss/accuracy, indicating that reported gains can arise from generic regularisation rather than faithful knowledge transmission. The targeted language experiment confirms KD can copy specific erroneous patterns, and scaling law experiments show capacity amplifies incorrect agreement faster than correct. We therefore reframe KD as a data-dependent regulariser with asymmetric knowledge sharing, with clear safety implications: audit teacher error structure and report functional transfer analyses (correct vs. incorrect agreement) alongside accuracy/loss.
Appendix A Safety Implications of Knowledge Distillation
The insights from our results can be summarised into three key points. 1) knowledge distillation enables statistically significant functional transfer. 2) The accuracy and loss benefits provided by knowledge distillation are often matched or even exceeded by random controls. 3) Knowledge distillation disproportionately transfers incorrect information, with this asymmetry increasing as the proportion of knowledge transfer grows. Considering these findings – particularly points 2 and 3 – Knowledge Distillation raises significant safety concerns. While it is often assumed that knowledge distillation benefits student models, our results challenge this notion by demonstrating a high likelihood that backdoors or harmful artifacts within teacher models could be transferred to student models. We present a concrete case of adversarial transfer facilitated by Knowledge Distillation in Appendix Section H.2. Moreover, we argue that knowledge distillation is not a safe or reliable method. At best, it results in minimal positive transfer, and at worst, it facilitates substantial negative transfer from teacher to student, undermining its practical utility.
Appendix B Extended Functional Analysis: Information-Theoretic and Geometric Perspectives
We apply two additional metrics: Variation of Information (VoI), an information-theoretic measure over discrete labellings that penalises confident mispredictions (Meilă, 2003), and Orthogonal Procrustes Distance (OPD), a geometric alignment metric over output representations (Schönemann, 1966; Ding et al., 2021). We compute VoI and OPD for two representative setups: ResNet18 on SVHN and ResNet50 on TinyImageNet (teacher seed 0). OPD closely tracks trends observed in Activation Distance and JS Divergence, showing decreasing student–teacher discrepancy as increases. VoI generally follows this trend, but diverges in specific cases (high on SVHN) where it increases despite stronger functional alignment. This is not contradictory: VoI penalises confident yet incorrect predictions more heavily than other metrics. Its rise coincides with the strongest observed increase in student–teacher agreement on incorrect predictions, providing further evidence of KD’s asymmetric payoff. Overall, OPD confirms alignment, but VoI reveals when that alignment corresponds to the transfer of incorrect information. Moreover, this behaviour is predicted by our gradient-based analysis: the per-logit gradient under KD pulls the student toward the teacher’s incorrect predictions with strength proportional to and to the teacher’s own loss. VoI captures the cost of absorbing these errors, providing an explicit signal of negative information transfer. OPD, meanwhile, confirms that overall alignment is occurring, but not necessarily to the student’s benefit.
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.063 +- 0.002 | 0.064 +- 0.001 | 0.060 +- 0.001 | 0.059 +- 0.001 | 0.144 +- 0.001 | 0.493 +- 0.000 | 0.849 +- 0.000 |
| Rank Disagreement | 0.696 +- 0.003 | 0.688 +- 0.004 | 0.684 +- 0.003 | 0.681 +- 0.003 | 0.800 +- 0.002 | 0.798 +- 0.002 | 0.802 +- 0.003 |
| Prediction Disagreement | 0.045 +- 0.001 | 0.046 +- 0.001 | 0.043 +- 0.001 | 0.042 +- 0.001 | 0.042 +- 0.001 | 0.043 +- 0.001 | 0.046 +- 0.001 |
| JS Divergence | 0.025 +- 0.001 | 0.025 +- 0.001 | 0.023 +- 0.001 | 0.022 +- 0.000 | 0.053 +- 0.000 | 0.201 +- 0.000 | 0.431 +- 0.000 |
| Information Variation | 0.550 +- 0.051 | 0.588 +- 0.049 | 0.594 +- 0.024 | 0.614 +- 0.018 | 0.638 +- 0.000 | 0.638 +- 0.000 | 0.638 +- 0.000 |
| Procrustes Distance | 0.165 +- 0.003 | 0.168 +- 0.004 | 0.164 +- 0.003 | 0.162 +- 0.005 | 0.291 +- 0.001 | 0.304 +- 0.001 | 0.311 +- 0.003 |
| Accuracy | 0.952 +- 0.001 | 0.951 +- 0.001 | 0.954 +- 0.001 | 0.954 +- 0.001 | 0.957 +- 0.001 | 0.957 +- 0.001 | 0.955 +- 0.001 |
| Loss | 0.385 +- 0.011 | 0.344 +- 0.008 | 0.310 +- 0.006 | 0.293 +- 0.004 | 0.236 +- 0.003 | 0.692 +- 0.001 | 1.698 +- 0.001 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.157 +- 0.001 | 0.157 +- 0.001 | 0.156 +- 0.001 | 0.155 +- 0.000 | 0.343 +- 0.000 | 0.581 +- 0.000 | 0.791 +- 0.000 |
| Rank Disagreement | 0.939 +- 0.000 | 0.939 +- 0.000 | 0.939 +- 0.000 | 0.939 +- 0.000 | 0.980 +- 0.000 | 0.984 +- 0.000 | 0.984 +- 0.000 |
| Prediction Disagreement | 0.153 +- 0.001 | 0.152 +- 0.001 | 0.151 +- 0.001 | 0.151 +- 0.001 | 0.190 +- 0.001 | 0.214 +- 0.000 | 0.324 +- 0.000 |
| JS Divergence | 0.040 +- 0.000 | 0.040 +- 0.000 | 0.039 +- 0.000 | 0.039 +- 0.000 | 0.171 +- 0.000 | 0.333 +- 0.000 | 0.533 +- 0.000 |
| Information Variation | 0.519 +- 0.017 | 0.520 +- 0.017 | 0.518 +- 0.022 | 0.533 +- 0.014 | 0.856 +- 0.002 | 0.897 +- 0.001 | 0.907 +- 0.002 |
| Procrustes Distance | 0.050 +- 0.000 | 0.050 +- 0.000 | 0.050 +- 0.000 | 0.049 +- 0.000 | 0.433 +- 0.000 | 0.664 +- 0.000 | 0.553 +- 0.000 |
| Accuracy | 0.605 +- 0.001 | 0.605 +- 0.000 | 0.604 +- 0.001 | 0.605 +- 0.001 | 0.607 +- 0.000 | 0.606 +- 0.001 | 0.580 +- 0.000 |
| Loss | 2.068 +- 0.001 | 2.065 +- 0.002 | 2.055 +- 0.001 | 2.043 +- 0.002 | 1.977 +- 0.001 | 2.497 +- 0.001 | 3.612 +- 0.002 |
Appendix C Feature Map Matching Knowledge Distillation
The functional-similarity framework we introduce is agnostic to the form of teacher supervision: relation, feature, and contrastive approaches all deliver a teacher-derived signal that ultimately shapes the student’s output distribution. If a variant truly transfers richer or safer knowledge, it should manifest as higher functional similarity without the asymmetric amplification of teacher errors that we document.
To verify this, we run feature-map matching knowledge distillation (Romero et al., 2015) on the transformer model NanoGPT trained on Tiny Shakespeare. In this process, we try to align blocks in the transformers using Mean Squared Error (MSE) on the intermediate blocks’ outputs. We include this alignment in the backpropagation step111Feature-map matching knowledge distillation implementation: https://docs.pytorch.org/tutorials/beginner/knowledge_distillation_tutorial.html. We chose this dataset because it represents the case where standard knowledge distillation leads to the most significant negative asymmetric transfer.
When we run feature-map matching KD (Feature Map KD), we observe statistically significant knowledge transfer for blocks 4 and 5. Tables 11 and 12 report these results independently. However, we continue to observe asymmetric incorrect transfer, as shown in Figure 14(a). It is important to note that block 4 experiences less functional similarity transfer than block 5. As expected, this leads to less negative asymmetric transfer than observed for feature-map KD on block 5. The best accuracy is again recorded when using RCD for both blocks 4 and 5, but at a higher alpha value of 0.5, compared to the best results typically recorded for 0.1 with standard KD.
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.202 0.000 | 0.203 0.000 | 0.197 0.000 | 0.191 0.000 | 0.209 0.000 | 0.203 0.000 | 0.224 0.001 |
| Rank Disagreement | 0.915 0.000 | 0.91 0.000 | 0.905 0.000 | 0.904 0.000 | 0.917 0.000 | 0.916 0.000 | 0.920 0.000 |
| Prediction Disagreement | 0.252 0.000 | 0.253 0.001 | 0.246 0.001 | 0.241 0.000 | 0.259 0.000 | 0.253 0.001 | 0.279 0.001 |
| JS Divergence | 0.056 0.000 | 0.056 0.000 | 0.053 0.000 | 0.050 0.000 | 0.059 0.000 | 0.057 0.000 | 0.067 0.001 |
| Accuracy | 0.571 0.000 | 0.574 0.000 | 0.573 0.000 | 0.570 0.000 | 0.574 0.000 | 0.578 0.000 | 0.566 0.001 |
| Loss | 1.473 0.002 | 1.542 0.003 | 1.569 0.002 | 1.585 0.001 | 1.573 0.002 | 1.552 0.003 | 1.542 0.004 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.202 0.000 | 0.201 0.000 | 0.183 0.000 | 0.160 0.001 | 0.214 0.001 | 0.211 0.001 | 0.227 0.001 |
| Rank Disagreement | 0.915 0.000 | 0.904 0.000 | 0.89 0.000 | 0.874 0.000 | 0.922 0.000 | 0.922 0.000 | 0.923 0.000 |
| Prediction Disagreement | 0.252 0.000 | 0.251 0.001 | 0.233 0.001 | 0.204 0.001 | 0.264 0.001 | 0.259 0.001 | 0.280 0.002 |
| JS Divergence | 0.056 0.000 | 0.056 0.000 | 0.046 0.000 | 0.035 0.000 | 0.062 0.000 | 0.060 0.000 | 0.066 0.000 |
| Accuracy | 0.571 0.000 | 0.574 0.000 | 0.577 0.000 | 0.576 0.000 | 0.572 0.000 | 0.575 0.000 | 0.564 0.001 |
| Loss | 1.473 0.002 | 1.551 0.002 | 1.532 0.001 | 1.493 0.001 | 1.599 0.001 | 1.591 0.002 | 1.590 0.002 |
| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
| KD 0.1 | ✗✓ | ✓✓ | ✗✗ | ✗✗ | ✗✗ | ✗✗ |
| KD 0.5 | ✓✓ | ✓✓ | ✓✓ | ✓✓ | ✗✓ | ✗✗ |
| KD 0.9 | ✓✓ | ✓✓ | ✓✓ | ✓✓ | ✗✗ | ✗✗ |
Largely we see that the results for Feature Map KD correspond to our original findings, when there is statistically significant functional transfer the transfer is asymmetric in nature and is weighted towards incorrect predictions. While there is a difference between blocks 4 and 5, understanding this fully this would require further exploration to make concrete statements about why this difference emerges.
Appendix D Random Control Distillation (RCD) Comparison to Label Smoothing
One potential confound in understanding KD’s effects is label smoothing: KD introduces soft targets, which may act as a form of regularisation independent of semantic knowledge transfer. To isolate this effect, we evaluate a baseline trained with classic label smoothing (LS), using the same loss structure but no teacher.
We also rely on RCD, which retains soft targets but replaces the teacher’s logits with uniform noise. RCD preserves any label-smoothing benefit while removing semantic content. Across all metrics, we find that LS and RCD match or exceed KD in accuracy, yet exhibit no increase in functional similarity with the teacher, particularly on incorrect predictions. This confirms that KD’s asymmetric error transfer arises from the specific structure of the teacher’s logits, not from softening per se.
| Metrics | Control | Knowledge Distillation | Random Control Distillation | Label Smoothing | ||||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.157 0.001 | 0.157 0.001 | 0.156 0.001 | 0.155 0.000 | 0.343 0.000 | 0.581 0.000 | 0.791 0.000 | 0.342 0.000 | 0.581 0.000 | 0.791 0.000 |
| Rank Disagreement | 0.939 0.000 | 0.939 0.000 | 0.939 0.000 | 0.939 0.000 | 0.980 0.000 | 0.984 0.000 | 0.984 0.000 | 0.980 0.000 | 0.984 0.000 | 0.984 0.000 |
| Prediction Disagreement | 0.153 0.001 | 0.152 0.001 | 0.151 0.001 | 0.151 0.001 | 0.190 0.001 | 0.214 0.000 | 0.324 0.000 | 0.189 0.001 | 0.214 0.000 | 0.324 0.000 |
| JS Divergence | 0.040 0.000 | 0.040 0.000 | 0.039 0.000 | 0.039 0.000 | 0.171 0.000 | 0.333 0.000 | 0.533 0.000 | 0.170 0.000 | 0.333 0.000 | 0.533 0.000 |
| Accuracy | 0.605 0.001 | 0.605 0.000 | 0.604 0.001 | 0.605 0.001 | 0.607 0.000 | 0.606 0.001 | 0.580 0.000 | 0.608 0.000 | 0.605 0.000 | 0.580 0.000 |
| Loss | 2.068 0.001 | 2.065 0.002 | 2.055 0.001 | 2.043 0.002 | 1.977 0.001 | 2.497 0.001 | 3.612 0.002 | 1.976 0.001 | 2.498 0.001 | 3.612 0.002 |
Appendix E Knowledge Distillation to Smaller Student
Justification:
This setup allows for an analysis of Knowledge Distillation where the student model is smaller than the teacher model, as expected in practice.
Caveat:
Although this moves away from our traditional experiential setup where the student can perfectly match the teacher, we use this example to show how transfer works between a larger teacher to a smaller student. It is important to note that using a smaller student introduces uncertainty on if the student capacity is a bottleneck to knowledge transfer. However, given that in practice Knowledge Distillation is used in this setting we show how our fundamental insights from the self distillation case transfer to other cases of dilatation. Our study of using a smaller students is not exhaustive but demonstrative and verifies the findings presented in the main body of the paper, and the utility of our initial experimental setup. Other than the architecture’s implicit bias towards the problem, which affects its performance (loss and accuracy), there are no confounding factors that could influence Knowledge Distillation.
E.1 TinyImageNet ResNet50 Teacher to ResNet18 Student
Training Settings:
The ResNet50 teacher model was trained with stochastic gradient descent with a learning rate of 0.01 and a Cosine annealing learning rate scheduler with a T_max set at 100. It was trained for 100 epochs with a batch size of 256. The data was normalized with a mean of (0.485, 0.456, 0.406) and a standard deviation of (0.229, 0.224, 0.225). The ResNet18 student model was trained under the same conditions.
Findings:
We observe a low train loss for the teacher model circa 0.0014 with a high train accuracy circa 0.9998; see Table 15. This low train loss corresponds as expected, with no significant knowledge transfer across alpha values; see Tables 16, 17, 18 and 19. This result is as expected from the results and intuition presented in the results of the main body of the paper. It highlights how this finding generalises to the practical KD environment.
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.001426 | 0.999800 | 2.070590 | 0.605300 |
| 1 | 0.001393 | 0.999800 | 2.051494 | 0.607900 |
| 2 | 0.001436 | 0.999800 | 2.051024 | 0.610600 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.548 0.000 | 0.548 0.000 | 0.547 0.000 | 0.547 0.000 | 0.565 0.000 | 0.651 0.000 | 0.828 0.000 |
| Rank Disagreement | 0.987 0.000 | 0.987 0.000 | 0.987 0.000 | 0.987 0.000 | 0.990 0.000 | 0.990 0.000 | 0.991 0.000 |
| Prediction Disagreement | 0.498 0.001 | 0.497 0.000 | 0.497 0.001 | 0.497 0.000 | 0.512 0.001 | 0.493 0.001 | 0.754 0.000 |
| JS Divergence | 0.281 0.000 | 0.281 0.000 | 0.280 0.000 | 0.281 0.000 | 0.330 0.000 | 0.400 0.000 | 0.599 0.000 |
| Accuracy | 0.503 0.001 | 0.504 0.001 | 0.504 0.000 | 0.503 0.000 | 0.493 0.000 | 0.512 0.000 | 0.236 0.000 |
| Loss | 2.604 0.001 | 2.602 0.002 | 2.594 0.001 | 2.589 0.001 | 2.434 0.001 | 2.641 0.001 | 4.684 0.002 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.548 0.000 | 0.548 0.000 | 0.548 0.000 | 0.547 0.000 | 0.567 0.000 | 0.651 0.000 | 0.829 0.000 |
| Rank Disagreement | 0.987 0.000 | 0.987 0.000 | 0.987 0.000 | 0.987 0.000 | 0.990 0.000 | 0.990 0.000 | 0.991 0.000 |
| Prediction Disagreement | 0.497 0.001 | 0.497 0.001 | 0.497 0.001 | 0.496 0.001 | 0.511 0.001 | 0.489 0.000 | 0.762 0.000 |
| JS Divergence | 0.281 0.000 | 0.281 0.000 | 0.281 0.000 | 0.280 0.000 | 0.331 0.000 | 0.401 0.000 | 0.601 0.000 |
| Accuracy | 0.503 0.000 | 0.504 0.000 | 0.504 0.000 | 0.504 0.000 | 0.494 0.000 | 0.513 0.001 | 0.232 0.000 |
| Loss | 2.608 0.002 | 2.606 0.002 | 2.599 0.002 | 2.591 0.003 | 2.431 0.002 | 2.634 0.001 | 4.703 0.002 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.546 0.000 | 0.545 0.000 | 0.545 0.000 | 0.545 0.000 | 0.565 0.000 | 0.651 0.000 | 0.829 0.000 |
| Rank Disagreement | 0.987 0.000 | 0.987 0.000 | 0.987 0.000 | 0.987 0.000 | 0.990 0.000 | 0.990 0.000 | 0.991 0.000 |
| Prediction Disagreement | 0.497 0.001 | 0.497 0.001 | 0.497 0.001 | 0.496 0.001 | 0.511 0.001 | 0.489 0.000 | 0.755 0.000 |
| JS Divergence | 0.280 0.000 | 0.280 0.000 | 0.280 0.000 | 0.280 0.000 | 0.330 0.000 | 0.400 0.000 | 0.600 0.000 |
| Accuracy | 0.503 0.001 | 0.504 0.000 | 0.503 0.000 | 0.503 0.000 | 0.493 0.000 | 0.512 0.000 | 0.236 0.000 |
| Loss | 2.604 0.001 | 2.602 0.001 | 2.594 0.001 | 2.587 0.001 | 2.434 0.001 | 2.641 0.001 | 4.684 0.002 |
| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
| KD 0.1 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
| KD 0.9 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
E.2 Tiny Shakespeare Nano-GPT Teacher to Pico-GPT Student
Training Settings:
The Nano-GPT Teacher is a GPT2-style transformer with an embedding dimension of 384, a vocabulary size of 65, six attention heads, six transformer blocks, a dropout of 0.200, and a block size of 256. The Pico-GPT student has an embedding dimension of 192, halving the internal width of the model; all other model settings are the same as the teacher.
The teacher and student are trained on the Tiny Shakespeare dataset, with the first 90% used for training and the last 10% used for testing. The dataset was tokenised via a character tokeniser, and the model was trained auto-regressively to predict the next character token. The teacher and student are trained with the Adam optimiser with a learning rate of 3e-4 with a batch size of 64 for 5000 iterations. The student models are trained with the same seeds and data orders from seeds 10 to 19 for the 10 models used for averaging. This is repeated for the three teachers trained on seeds 0 to 2.
Justification:
This setup allows for an analysis of Knowledge Distillation where the student model is smaller than the teacher model, as expected in practice. It is not exhaustive but demonstrative that the findings we present in the main body of the paper generalise to this case. Other than the architecture’s implicit bias towards the problem, which affects its performance (loss and accuracy), no confounding factors could influence Knowledge Distillation.
Findings:
We observe a high train loss for the teacher model circa 0.86 with a high train accuracy circa 0.72; see Table 20. This high train loss corresponds as expected with a substantial knowledge transfer which increases as alpha increases, see Tables 96, 97, 98 and 99. This substantial knowledge transfer coincides with an asymmetric payoff in prediction agreement, strongly favouring incorrect predictions, see Figure 65(a). This result is as expected from the results and intuition presented in the results of the main body of the paper and highlights how this finding generalises.
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.864641 | 0.719685 | 1.567481 | 0.573366 |
| 1 | 0.866370 | 0.719697 | 1.561079 | 0.574668 |
| 2 | 0.861098 | 0.721140 | 1.562137 | 0.573033 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.202 0.000 | 0.198 0.000 | 0.181 0.000 | 0.172 0.000 | 0.221 0.000 | 0.399 0.000 | 0.663 0.000 |
| Rank Disagreement | 0.915 0.000 | 0.915 0.000 | 0.912 0.000 | 0.911 0.000 | 0.939 0.000 | 0.944 0.000 | 0.950 0.000 |
| Prediction Disagreement | 0.252 0.000 | 0.247 0.000 | 0.226 0.000 | 0.214 0.000 | 0.252 0.000 | 0.253 0.001 | 0.272 0.001 |
| JS Divergence | 0.056 0.000 | 0.054 0.000 | 0.047 0.000 | 0.043 0.000 | 0.075 0.000 | 0.203 0.000 | 0.451 0.000 |
| Accuracy | 0.571 0.000 | 0.572 0.000 | 0.575 0.000 | 0.574 0.000 | 0.571 0.000 | 0.570 0.000 | 0.561 0.000 |
| Loss | 1.473 0.002 | 1.471 0.002 | 1.472 0.001 | 1.496 0.002 | 1.483 0.001 | 1.870 0.001 | 3.017 0.002 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.201 0.000 | 0.196 0.000 | 0.180 0.000 | 0.170 0.000 | 0.217 0.000 | 0.392 0.000 | 0.655 0.000 |
| Rank Disagreement | 0.916 0.000 | 0.915 0.000 | 0.912 0.000 | 0.911 0.000 | 0.939 0.000 | 0.944 0.000 | 0.950 0.000 |
| Prediction Disagreement | 0.257 0.000 | 0.251 0.000 | 0.231 0.000 | 0.219 0.000 | 0.256 0.000 | 0.258 0.000 | 0.277 0.001 |
| JS Divergence | 0.055 0.000 | 0.053 0.000 | 0.046 0.000 | 0.043 0.000 | 0.074 0.000 | 0.201 0.000 | 0.449 0.000 |
| Accuracy | 0.571 0.000 | 0.573 0.000 | 0.575 0.000 | 0.574 0.000 | 0.571 0.000 | 0.570 0.000 | 0.561 0.000 |
| Loss | 1.473 0.002 | 1.473 0.002 | 1.475 0.002 | 1.492 0.002 | 1.483 0.001 | 1.870 0.001 | 3.017 0.002 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.202 0.000 | 0.197 0.000 | 0.180 0.000 | 0.171 0.000 | 0.219 0.000 | 0.395 0.001 | 0.660 0.000 |
| Rank Disagreement | 0.915 0.000 | 0.914 0.000 | 0.912 0.000 | 0.910 0.000 | 0.939 0.000 | 0.944 0.000 | 0.949 0.000 |
| Prediction Disagreement | 0.252 0.000 | 0.246 0.000 | 0.226 0.000 | 0.215 0.000 | 0.250 0.001 | 0.251 0.000 | 0.272 0.001 |
| JS Divergence | 0.055 0.000 | 0.053 0.000 | 0.046 0.000 | 0.043 0.000 | 0.074 0.000 | 0.202 0.000 | 0.450 0.000 |
| Accuracy | 0.571 0.000 | 0.572 0.000 | 0.575 0.000 | 0.574 0.000 | 0.572 0.000 | 0.571 0.000 | 0.561 0.000 |
| Loss | 1.475 0.001 | 1.470 0.001 | 1.471 0.002 | 1.491 0.002 | 1.482 0.001 | 1.865 0.002 | 3.017 0.001 |
| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
| KD 0.1 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✗ | ✗✗✓ |
| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ |
Appendix F Vision Results
F.1 TinyImageNet
Training Settings:
The ResNet50 model was trained with stochastic gradient descent with a learning rate 0.01, along with a Cosine annealing learning rate scheduler with a T_max set at 100. It was trained for 100 epochs with a batch size of 256. The data was normalized with a mean of (0.485, 0.456, 0.406) and standard deviation of (0.229, 0.224, 0.225). For ResNet50 with RandAugment (Cubuk et al., 2020), the only difference between base ResNet is the introduction of RandAugment with the default setting provided in Pytorch 2.4 (Paszke et al., 2019). The VGG19 and VGG19 with RandAugment has the same setup as the ResNet50 and ResNet50 with RandAugment respectively however it was trained with momentum of 0.9.
F.1.1 ResNet50
Findings:
For the ResNet50 on TinyImageNet, we observe that the teacher seeds, Table 25, obtain a low train loss of 0.001 and a train accuracy of 0.99. This train performance coincides with a test accuracy of circa 0.60, resulting in a generalisation gap of circa 0.39.
For an alpha of 0.1, Table 29, we observe no significant knowledge transfer across all metrics except for Rank Disagreement with teacher seed 0. It has statistically significant transfer, but the increased similarity is extremely marginal, as observed with SIDDO and KD 0.1 having the same value to 3 significant figures, see Table 26. With this, we see a marginal prediction agreement of less than 0.5% for correct and incorrect predictions across teacher seeds, Figure 23(a). For alpha 0.5 and 0.9, we observe significant knowledge transfer for all bar Prediction Disagreement with alpha of 0.5 and 0.9 for teacher seed 2. However, this transfer is marginal, Tables 26, 27 and 28, and we observe a prediction agreement of less than 0.5% for correct and incorrect predictions across teacher seeds, Figure 17(a).
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.001426 | 0.999800 | 2.070590 | 0.605300 |
| 1 | 0.001393 | 0.999800 | 2.051494 | 0.607900 |
| 2 | 0.001436 | 0.999800 | 2.051024 | 0.610600 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.157 0.001 | 0.157 0.001 | 0.156 0.001 | 0.155 0.000 | 0.343 0.000 | 0.581 0.000 | 0.791 0.000 |
| Rank Disagreement | 0.939 0.000 | 0.939 0.000 | 0.939 0.000 | 0.939 0.000 | 0.980 0.000 | 0.984 0.000 | 0.984 0.000 |
| Prediction Disagreement | 0.153 0.001 | 0.152 0.001 | 0.151 0.001 | 0.151 0.001 | 0.190 0.001 | 0.214 0.000 | 0.324 0.000 |
| JS Divergence | 0.040 0.000 | 0.040 0.000 | 0.039 0.000 | 0.039 0.000 | 0.171 0.000 | 0.333 0.000 | 0.533 0.000 |
| Accuracy | 0.605 0.001 | 0.605 0.000 | 0.604 0.001 | 0.605 0.001 | 0.607 0.000 | 0.606 0.001 | 0.580 0.000 |
| Loss | 2.068 0.001 | 2.065 0.002 | 2.055 0.001 | 2.043 0.002 | 1.977 0.001 | 2.497 0.001 | 3.612 0.002 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.156 0.001 | 0.156 0.000 | 0.155 0.001 | 0.153 0.000 | 0.340 0.000 | 0.579 0.000 | 0.792 0.000 |
| Rank Disagreement | 0.940 0.000 | 0.940 0.000 | 0.939 0.000 | 0.939 0.000 | 0.980 0.000 | 0.984 0.000 | 0.984 0.000 |
| Prediction Disagreement | 0.148 0.001 | 0.149 0.001 | 0.148 0.001 | 0.146 0.001 | 0.185 0.001 | 0.209 0.000 | 0.330 0.000 |
| JS Divergence | 0.040 0.000 | 0.040 0.000 | 0.039 0.000 | 0.038 0.000 | 0.170 0.000 | 0.332 0.000 | 0.534 0.000 |
| Accuracy | 0.607 0.001 | 0.608 0.001 | 0.607 0.000 | 0.607 0.001 | 0.605 0.000 | 0.602 0.001 | 0.576 0.000 |
| Loss | 2.048 0.002 | 2.048 0.002 | 2.034 0.002 | 2.025 0.002 | 1.973 0.001 | 2.498 0.001 | 3.611 0.002 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.157 0.000 | 0.157 0.000 | 0.155 0.000 | 0.155 0.000 | 0.342 0.000 | 0.581 0.000 | 0.792 0.000 |
| Rank Disagreement | 0.939 0.000 | 0.939 0.000 | 0.939 0.000 | 0.939 0.000 | 0.980 0.000 | 0.984 0.000 | 0.984 0.000 |
| Prediction Disagreement | 0.152 0.001 | 0.152 0.001 | 0.151 0.001 | 0.151 0.001 | 0.187 0.001 | 0.213 0.001 | 0.327 0.000 |
| JS Divergence | 0.040 0.000 | 0.040 0.000 | 0.039 0.000 | 0.039 0.000 | 0.171 0.000 | 0.334 0.000 | 0.534 0.000 |
| Accuracy | 0.608 0.001 | 0.607 0.001 | 0.607 0.000 | 0.609 0.001 | 0.608 0.001 | 0.605 0.001 | 0.577 0.000 |
| Loss | 2.054 0.002 | 2.050 0.002 | 2.040 0.003 | 2.025 0.002 | 1.967 0.001 | 2.494 0.001 | 3.602 0.002 |
| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
| KD 0.1 | ✗✗✗ | ✓✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✗ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
F.1.2 ResNet50 with RandAugment
Findings:
For the ResNet50 on TinyImageNet with RandAugment, we observe that the teacher seeds, Table 25, obtain a high train loss and a train accuracy of circa 0.84. This train performance coincides with a test accuracy of circa 0.64, resulting in a generalisation gap of circa 0.2.
We observe significant knowledge transfer for all alpha values with a strong asymmetric transfer of knowledge favouring incorrect predictions as shown in Table 34 and Figure 26(a), respectively. However, it is important to note that despite significant and substantial knowledge transfer, we do not see any improvement in test accuracy over the control and random controls.
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.672748 | 0.840410 | 1.620552 | 0.638800 |
| 1 | 0.678245 | 0.839200 | 1.629393 | 0.641800 |
| 2 | 0.667570 | 0.840750 | 1.624969 | 0.641100 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.193 0.000 | 0.183 0.000 | 0.150 0.000 | 0.131 0.000 | 0.245 0.001 | 0.501 0.001 | 0.781 0.000 |
| Rank Disagreement | 0.959 0.000 | 0.957 0.000 | 0.948 0.000 | 0.943 0.000 | 0.975 0.000 | 0.981 0.000 | 0.987 0.000 |
| Prediction Disagreement | 0.196 0.001 | 0.188 0.001 | 0.154 0.001 | 0.136 0.001 | 0.195 0.001 | 0.240 0.001 | 0.572 0.001 |
| JS Divergence | 0.058 0.000 | 0.052 0.000 | 0.036 0.000 | 0.028 0.000 | 0.094 0.000 | 0.266 0.000 | 0.563 0.000 |
| Accuracy | 0.640 0.000 | 0.643 0.001 | 0.644 0.000 | 0.642 0.000 | 0.646 0.001 | 0.657 0.001 | 0.400 0.001 |
| Loss | 1.619 0.003 | 1.600 0.001 | 1.578 0.001 | 1.577 0.001 | 1.551 0.001 | 1.984 0.002 | 4.211 0.001 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.194 0.000 | 0.183 0.001 | 0.148 0.000 | 0.13 0.000 | 0.247 0.000 | 0.503 0.000 | 0.783 0.000 |
| Rank Disagreement | 0.959 0.000 | 0.957 0.000 | 0.948 0.000 | 0.943 0.000 | 0.975 0.000 | 0.981 0.000 | 0.987 0.000 |
| Prediction Disagreement | 0.195 0.001 | 0.186 0.001 | 0.151 0.001 | 0.134 0.001 | 0.194 0.001 | 0.241 0.000 | 0.577 0.001 |
| JS Divergence | 0.058 0.000 | 0.053 0.000 | 0.036 0.000 | 0.028 0.000 | 0.095 0.000 | 0.267 0.000 | 0.565 0.000 |
| Accuracy | 0.639 0.001 | 0.640 0.001 | 0.641 0.001 | 0.640 0.001 | 0.646 0.001 | 0.658 0.000 | 0.396 0.001 |
| Loss | 1.620 0.002 | 1.608 0.002 | 1.584 0.001 | 1.584 0.001 | 1.555 0.002 | 1.986 0.002 | 4.214 0.002 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.195 0.000 | 0.185 0.000 | 0.150 0.000 | 0.131 0.000 | 0.247 0.001 | 0.504 0.000 | 0.783 0.000 |
| Rank Disagreement | 0.959 0.000 | 0.957 0.000 | 0.948 0.000 | 0.943 0.000 | 0.975 0.000 | 0.981 0.000 | 0.987 0.000 |
| Prediction Disagreement | 0.197 0.001 | 0.189 0.001 | 0.155 0.001 | 0.135 0.001 | 0.197 0.001 | 0.239 0.000 | 0.564 0.001 |
| JS Divergence | 0.059 0.000 | 0.053 0.000 | 0.037 0.000 | 0.028 0.000 | 0.096 0.000 | 0.267 0.000 | 0.563 0.000 |
| Accuracy | 0.640 0.001 | 0.641 0.001 | 0.643 0.001 | 0.643 0.000 | 0.647 0.001 | 0.657 0.000 | 0.410 0.001 |
| Loss | 1.621 0.002 | 1.606 0.001 | 1.581 0.001 | 1.582 0.001 | 1.552 0.001 | 1.982 0.002 | 4.180 0.002 |
| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
| KD 0.1 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
F.1.3 VGG19
Findings:
For the VGG19 on the TinyImageNet, we observe a low train loss of circa and a train accuracy of 0.9998. As expected, given our results and discussion in the main body of the paper on the ResNet50, we see no significant transfer until an alpha of 0.9. With teacher seed 0 and 2 with an alpha of 0.9, we record significant transfer for Activation Distance and for teacher seed 0 on JS Divergence, as seen in Table 39. When we observe knowledge transfer with an alpha of 0.9, we observe a slight preference for positive agreement of test prediction; however, the results have a large SEM, and the amount of agreement is less than 0.5%, making the results less reliable and insignificant in either transfer direction.
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.000286 | 0.999800 | 3.351542 | 0.633200 |
| 1 | 0.000286 | 0.999800 | 3.301587 | 0.637200 |
| 2 | 0.000285 | 0.999800 | 3.311130 | 0.633500 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.418 0.001 | 0.419 0.001 | 0.418 0.001 | 0.416 0.001 | 0.522 0.001 | 0.741 0.000 | 0.886 0.000 |
| Rank Disagreement | 0.978 0.000 | 0.978 0.000 | 0.978 0.000 | 0.978 0.000 | 0.987 0.000 | 0.988 0.000 | 0.989 0.000 |
| Prediction Disagreement | 0.332 0.001 | 0.332 0.001 | 0.332 0.001 | 0.330 0.001 | 0.348 0.001 | 0.381 0.001 | 0.412 0.000 |
| JS Divergence | 0.195 0.000 | 0.195 0.000 | 0.195 0.000 | 0.194 0.000 | 0.308 0.001 | 0.457 0.000 | 0.593 0.000 |
| Accuracy | 0.635 0.001 | 0.635 0.001 | 0.636 0.001 | 0.638 0.001 | 0.627 0.001 | 0.603 0.001 | 0.576 0.001 |
| Loss | 3.332 0.010 | 3.329 0.012 | 3.308 0.011 | 3.313 0.010 | 2.003 0.005 | 2.732 0.002 | 3.682 0.002 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.414 0.002 | 0.414 0.001 | 0.413 0.001 | 0.413 0.001 | 0.522 0.001 | 0.742 0.000 | 0.886 0.000 |
| Rank Disagreement | 0.978 0.000 | 0.978 0.000 | 0.978 0.000 | 0.978 0.000 | 0.987 0.000 | 0.988 0.000 | 0.989 0.000 |
| Prediction Disagreement | 0.329 0.001 | 0.329 0.001 | 0.328 0.001 | 0.328 0.001 | 0.348 0.001 | 0.379 0.001 | 0.410 0.000 |
| JS Divergence | 0.194 0.001 | 0.194 0.001 | 0.193 0.001 | 0.193 0.001 | 0.308 0.000 | 0.457 0.000 | 0.593 0.000 |
| Accuracy | 0.635 0.001 | 0.636 0.001 | 0.638 0.001 | 0.637 0.001 | 0.627 0.001 | 0.603 0.001 | 0.574 0.001 |
| Loss | 3.345 0.011 | 3.318 0.009 | 3.306 0.009 | 3.311 0.010 | 2.004 0.004 | 2.733 0.004 | 3.682 0.002 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.419 0.001 | 0.417 0.001 | 0.418 0.001 | 0.417 0.001 | 0.524 0.000 | 0.743 0.000 | 0.886 0.000 |
| Rank Disagreement | 0.978 0.000 | 0.978 0.000 | 0.978 0.000 | 0.978 0.000 | 0.987 0.000 | 0.988 0.000 | 0.989 0.000 |
| Prediction Disagreement | 0.332 0.001 | 0.332 0.001 | 0.332 0.001 | 0.331 0.001 | 0.354 0.001 | 0.385 0.001 | 0.414 0.001 |
| JS Divergence | 0.196 0.000 | 0.195 0.001 | 0.196 0.000 | 0.195 0.000 | 0.309 0.000 | 0.458 0.000 | 0.593 0.000 |
| Accuracy | 0.635 0.001 | 0.636 0.000 | 0.635 0.001 | 0.637 0.001 | 0.626 0.001 | 0.602 0.001 | 0.577 0.001 |
| Loss | 3.314 0.009 | 3.298 0.004 | 3.318 0.011 | 3.263 0.009 | 1.998 0.004 | 2.738 0.003 | 3.681 0.002 |
| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
| KD 0.1 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
| KD 0.9 | ✓✗✓ | ✗✗✗ | ✗✗✗ | ✓✗✗ | ✗✗✗ | ✗✗✗ |
F.1.4 VGG19 with RandAugment
Findings:
For the VGG19 on the TinyImageNet with RandAugment, we observe a high train loss of circa and a train accuracy of circa 0.93. As expected, given the results on the RandAugment ResNet50 that we present in the main body of the paper, we see substantial transfer across all alpha values; see Tables 41, 42, 43 and 44. This substantial and significant transfer of knowledge, as expected, coincides with a strong asymmetric transfer of knowledge favouring incorrect predictions, as shown in Figure 32(a).
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.272582 | 0.933990 | 2.565560 | 0.622600 |
| 1 | 0.269916 | 0.935140 | 2.570119 | 0.618900 |
| 2 | 0.273968 | 0.934700 | 2.609870 | 0.620100 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.393 0.001 | 0.388 0.001 | 0.368 0.001 | 0.355 0.001 | 0.431 0.001 | 0.648 0.000 | 0.848 0.001 |
| Rank Disagreement | 0.976 0.000 | 0.976 0.000 | 0.975 0.000 | 0.974 0.000 | 0.985 0.000 | 0.987 0.000 | 0.987 0.000 |
| Prediction Disagreement | 0.335 0.001 | 0.333 0.001 | 0.320 0.001 | 0.312 0.001 | 0.341 0.001 | 0.352 0.001 | 0.396 0.004 |
| JS Divergence | 0.182 0.000 | 0.178 0.000 | 0.166 0.000 | 0.159 0.000 | 0.228 0.000 | 0.377 0.000 | 0.577 0.001 |
| Accuracy | 0.621 0.001 | 0.624 0.001 | 0.631 0.001 | 0.633 0.001 | 0.622 0.001 | 0.628 0.001 | 0.609 0.004 |
| Loss | 2.586 0.009 | 2.442 0.005 | 2.148 0.004 | 2.022 0.003 | 1.792 0.003 | 2.258 0.002 | 3.533 0.013 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.391 0.001 | 0.384 0.001 | 0.362 0.001 | 0.351 0.000 | 0.428 0.001 | 0.644 0.000 | 0.845 0.000 |
| Rank Disagreement | 0.977 0.000 | 0.976 0.000 | 0.975 0.000 | 0.974 0.000 | 0.985 0.000 | 0.987 0.000 | 0.987 0.000 |
| Prediction Disagreement | 0.333 0.001 | 0.330 0.001 | 0.316 0.001 | 0.308 0.001 | 0.337 0.001 | 0.348 0.001 | 0.392 0.001 |
| JS Divergence | 0.180 0.000 | 0.176 0.000 | 0.164 0.000 | 0.156 0.000 | 0.226 0.000 | 0.375 0.000 | 0.576 0.000 |
| Accuracy | 0.622 0.001 | 0.624 0.000 | 0.632 0.001 | 0.635 0.001 | 0.625 0.001 | 0.627 0.001 | 0.611 0.001 |
| Loss | 2.575 0.004 | 2.439 0.007 | 2.149 0.006 | 2.017 0.002 | 1.781 0.005 | 2.254 0.003 | 3.526 0.003 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.395 0.001 | 0.389 0.001 | 0.368 0.001 | 0.358 0.001 | 0.435 0.001 | 0.649 0.000 | 0.850 0.001 |
| Rank Disagreement | 0.977 0.000 | 0.977 0.000 | 0.975 0.000 | 0.975 0.000 | 0.985 0.000 | 0.987 0.000 | 0.987 0.000 |
| Prediction Disagreement | 0.335 0.001 | 0.334 0.001 | 0.321 0.001 | 0.313 0.001 | 0.341 0.001 | 0.352 0.001 | 0.403 0.010 |
| JS Divergence | 0.182 0.000 | 0.179 0.000 | 0.167 0.001 | 0.160 0.001 | 0.230 0.000 | 0.378 0.000 | 0.579 0.002 |
| Accuracy | 0.621 0.001 | 0.623 0.001 | 0.631 0.001 | 0.636 0.001 | 0.623 0.001 | 0.628 0.001 | 0.600 0.011 |
| Loss | 2.583 0.006 | 2.441 0.009 | 2.145 0.006 | 2.012 0.007 | 1.780 0.003 | 2.257 0.003 | 3.556 0.034 |
| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
| KD 0.1 | ✓✓✓ | ✓✓✓ | ✗✓✗ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ |
F.2 CIFAR10
Training Settings:
All CIFAR10 architectures are trained with Adam optimiser with a learning rate of 0.001 and a batch size of 256 for 100 epochs. All data is normalised with a mean of 0.5 and a standard deviation of 0.5. The student vision architectures are trained with the same seeds and data orders from seeds 10-19 for the 10 models used for averaging. As aligned with all experiments we conduct, this is repeated for the three teachers trained on seeds 0-2.
Justification:
This setup allows for a fair analysis of Knowledge Distillation as its role is isolated in the training process. Other than the architecture’s implicit bias towards the problem, which affects its performance (loss and accuracy), there are no confounding factors that could influence Knowledge Distillation.
Findings:
We find that the teacher models often significantly transfer knowledge to the student model, and this coincides with the teacher’s high loss on the training dataset. The ResNet has the lowest loss and no transfer, the VGG has a higher loss and some transfer, and the ViT has the highest loss and the most transfer. However, when knowledge is transferred, it often has a negative asymmetric payoff towards agreement between the teacher and the student on incorrect predictions.
F.2.1 ResNet18
Findings:
For the ResNet18 on CIFAR10, we observe that the teacher seeds, Table 45, obtain a very low train loss of and a train accuracy of 1. This train performance coincides with a high test accuracy of circa 0.86, resulting in a generalisation gap of circa 0.14. Table 49 shows no significant knowledge transfer across teacher seeds.
Due to the low train loss on the teacher seed, the teacher model is a nearly identical representation of the training labels, meaning there is low utility in the teacher model. As we observe, the controls of the models trained in the SIDDO condition is functionally different from the teacher, Tables 46, 47 and 48; despite having the same initialisation and only changing the data order, it is not a surprise that Knowledge Distillation in the setup does not add anything as the teacher is essentially the label, and thus creates a similar setup to the SIDDO condition.
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.000010 | 1.000000 | 0.869184 | 0.862100 |
| 1 | 0.000006 | 1.000000 | 0.833735 | 0.867200 |
| 2 | 0.000030 | 1.000000 | 0.739927 | 0.867000 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.174 | 0.175 | 0.172 | 0.174 | 0.244 | 0.538 | 0.843 |
| Rank Disagreement () | 0.659 | 0.659 | 0.656 | 0.655 | 0.795 | 0.802 | 0.807 |
| Prediction Disagreement () | 0.128 | 0.129 | 0.127 | 0.128 | 0.131 | 0.143 | 0.150 |
| JS Divergence () | 0.070 | 0.070 | 0.069 | 0.068 | 0.097 | 0.229 | 0.432 |
| Accuracy () | 0.861 | 0.862 | 0.862 | 0.862 | 0.865 | 0.856 | 0.854 |
| Loss () | 0.961 | 0.903 | 0.895 | 0.827 | 0.539 | 0.902 | 1.772 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.900 | 0.1 | 0.5 | 0.900 | |
| Activation Distance () | 0.167 | 0.164 | 0.165 | 0.165 | 0.240 | 0.533 | 0.841 |
| Rank Disagreement () | 0.653 | 0.649 | 0.650 | 0.650 | 0.796 | 0.803 | 0.807 |
| Prediction Disagreement () | 0.122 | 0.120 | 0.121 | 0.120 | 0.126 | 0.134 | 0.139 |
| JS Divergence () | 0.066 | 0.065 | 0.065 | 0.064 | 0.095 | 0.226 | 0.430 |
| Accuracy () | 0.865 | 0.867 | 0.866 | 0.867 | 0.866 | 0.860 | 0.859 |
| Loss () | 0.858 | 0.877 | 0.824 | 0.816 | 0.533 | 0.896 | 1.767 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.166 | 0.169 | 0.167 | 0.172 | 0.242 | 0.533 | 0.839 |
| Rank Disagreement () | 0.646 | 0.647 | 0.638 | 0.646 | 0.799 | 0.803 | 0.805 |
| Prediction Disagreement () | 0.122 | 0.124 | 0.124 | 0.127 | 0.132 | 0.140 | 0.142 |
| JS Divergence () | 0.065 | 0.066 | 0.064 | 0.067 | 0.096 | 0.226 | 0.429 |
| Accuracy () | 0.865 | 0.864 | 0.864 | 0.861 | 0.862 | 0.857 | 0.857 |
| Loss () | 0.892 | 0.887 | 0.803 | 0.798 | 0.549 | 0.900 | 1.769 |
| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
| KD 0.1 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
| KD 0.9 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
F.2.2 VGG19
Findings:
For the VGG19 on CIFAR10, we observe that the teacher seeds, Table 50, obtain a low train loss of circa 0.01 and a train accuracy of approximately 0.996. This train performance coincides with a high test accuracy of circa 0.86, resulting in a generalisation gap of circa 0.14. Table 54 shows a significant knowledge transfer with regard to Rank Disagreement for all teacher seeds when alpha is at 0.9.
At alpha 0.9 for teacher seed 0 and 2, there is an increase in agreement between the student and teacher on incorrect predictions over the correct predictions, Figure 38(a), which corresponds with the knowledge transfer. This result coincides with teachers seed 0 and 2 having a higher train loss than teacher seed 1, indicating that the teacher train loss plays an important role in knowledge transfer. For teacher seed 1, Figure 38(a), there is no significant increase in correct or incorrect prediction agreement between the student model and the teacher due to the deviation in the SEM.
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.011608 | 0.996760 | 0.858675 | 0.863900 |
| 1 | 0.009228 | 0.997080 | 0.798530 | 0.860800 |
| 2 | 0.012352 | 0.996420 | 0.801562 | 0.867100 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.206 | 0.199 | 0.203 | 0.197 | 0.264 | 0.541 | 0.842 |
| Rank Disagreement () | 0.701 | 0.705 | 0.658 | 0.640 | 0.811 | 0.819 | 0.819 |
| Prediction Disagreement () | 0.152 | 0.147 | 0.151 | 0.146 | 0.148 | 0.146 | 0.150 |
| JS Divergence () | 0.090 | 0.085 | 0.086 | 0.083 | 0.109 | 0.230 | 0.429 |
| Accuracy () | 0.864 | 0.869 | 0.867 | 0.869 | 0.870 | 0.871 | 0.868 |
| Loss () | 0.849 | 0.725 | 0.676 | 0.649 | 0.562 | 0.880 | 1.762 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.199 | 0.202 | 0.202 | 0.201 | 0.263 | 0.543 | 0.842 |
| Rank Disagreement () | 0.726 | 0.684 | 0.662 | 0.639 | 0.803 | 0.801 | 0.810 |
| Prediction Disagreement () | 0.147 | 0.150 | 0.150 | 0.149 | 0.148 | 0.149 | 0.153 |
| JS Divergence () | 0.086 | 0.087 | 0.086 | 0.085 | 0.107 | 0.230 | 0.428 |
| Accuracy () | 0.868 | 0.866 | 0.865 | 0.866 | 0.870 | 0.869 | 0.866 |
| Loss () | 0.799 | 0.735 | 0.680 | 0.666 | 0.562 | 0.887 | 1.762 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.196 | 0.199 | 0.196 | 0.193 | 0.258 | 0.541 | 0.844 |
| Rank Disagreement () | 0.672 | 0.649 | 0.633 | 0.602 | 0.809 | 0.817 | 0.816 |
| Prediction Disagreement () | 0.142 | 0.146 | 0.143 | 0.141 | 0.142 | 0.143 | 0.149 |
| JS Divergence () | 0.084 | 0.086 | 0.083 | 0.081 | 0.106 | 0.229 | 0.429 |
| Accuracy () | 0.870 | 0.864 | 0.868 | 0.867 | 0.871 | 0.871 | 0.867 |
| Loss () | 0.801 | 0.734 | 0.665 | 0.639 | 0.560 | 0.884 | 1.762 |
| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
| KD 0.1 | ✗✗✗ | ✗✓✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✗✗✗ | ✓✓✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
| KD 0.9 | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✓✗✗ | ✗✗✗ | ✗✗✗ |
F.2.3 ViT
Findings:
For the ViT on CIFAR10, we observe that the teacher seeds, Table 55, obtain a high train loss of and a train accuracy of approximately 0.98. This train performance coincides with a test accuracy of circa 0.63, resulting in a generalisation gap of circa 0.35. Table 59 shows a significant knowledge transfer on all teacher seeds when alpha is 0.5 and 0.9. For teacher seed 0 and 1 using alpha at 0.9, where there is sizeable knowledge transfer, we observe an asymmetric knowledge transfer favouring negative transfer in Figure 41(a).
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.043291 | 0.988260 | 1.864339 | 0.626900 |
| 1 | 0.056539 | 0.983160 | 1.772490 | 0.634200 |
| 2 | 0.046902 | 0.987100 | 1.714442 | 0.649600 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.900 | 0.1 | 0.5 | 0.900 | |
| Activation Distance () | 0.491 | 0.487 | 0.473 | 0.470 | 0.496 | 0.611 | 0.793 |
| Rank Disagreement () | 0.734 | 0.730 | 0.724 | 0.722 | 0.808 | 0.812 | 0.817 |
| Prediction Disagreement () | 0.385 | 0.383 | 0.374 | 0.373 | 0.383 | 0.380 | 0.386 |
| JS Divergence () | 0.201 | 0.198 | 0.189 | 0.186 | 0.206 | 0.277 | 0.411 |
| Accuracy () | 0.634 | 0.634 | 0.641 | 0.637 | 0.640 | 0.641 | 0.627 |
| Loss () | 1.773 | 1.695 | 1.52 | 1.451 | 1.258 | 1.351 | 1.943 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.485 | 0.477 | 0.461 | 0.455 | 0.489 | 0.609 | 0.791 |
| Rank Disagreement () | 0.733 | 0.728 | 0.717 | 0.714 | 0.806 | 0.808 | 0.816 |
| Prediction Disagreement () | 0.382 | 0.375 | 0.367 | 0.363 | 0.379 | 0.380 | 0.382 |
| JS Divergence () | 0.198 | 0.193 | 0.182 | 0.178 | 0.202 | 0.275 | 0.410 |
| Accuracy () | 0.637 | 0.643 | 0.644 | 0.648 | 0.643 | 0.636 | 0.630 |
| Loss () | 1.781 | 1.668 | 1.466 | 1.366 | 1.253 | 1.359 | 1.942 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.476 | 0.468 | 0.459 | 0.456 | 0.486 | 0.612 | 0.797 |
| Rank Disagreement () | 0.730 | 0.725 | 0.720 | 0.718 | 0.806 | 0.811 | 0.817 |
| Prediction Disagreement () | 0.372 | 0.366 | 0.363 | 0.360 | 0.371 | 0.374 | 0.375 |
| JS Divergence () | 0.195 | 0.189 | 0.183 | 0.180 | 0.201 | 0.277 | 0.413 |
| Accuracy () | 0.636 | 0.641 | 0.644 | 0.639 | 0.637 | 0.635 | 0.631 |
| Loss () | 1.788 | 1.673 | 1.498 | 1.458 | 1.282 | 1.361 | 1.942 |
| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
| KD 0.1 | ✗✓✓ | ✓✓✓ | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✓ | ✗✗✗ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
F.3 SVHN Dataset
Training Settings:
All SVHN architectures are trained with Adam optimiser with a learning rate of 0.001 and a batch size of 256 for 100 epochs. All data is normalised with a mean of 0.5 and a standard deviation of 0.5. The student vision architectures are trained with the same seeds and data orders from seeds 10-19 for the 10 models used for averaging. We repeated this, in line with our other experiments for the three teachers trained on seeds 0-2.
Justification:
This setup allows for a fair analysis of Knowledge Distillation as its role is isolated in the training process. Other than the architecture’s implicit bias towards the problem, which affects its performance (loss and accuracy), there are no confounding factors that could influence Knowledge Distillation.
Findings:
The teacher models often significantly transfer knowledge to the student model. However, the knowledge transfer is often inconsistent, and when transferred, it often has an asymmetric negative payoff.
F.3.1 ResNet18
Findings:
For the ResNet on SVHN, we observe that the teacher seeds, Table 60, obtain a range of train loss values of 0.000646, 0.000061 and 0.004657 for teacher seeds 0, 1, and 2, respectively. The train accuracies are approximately 0.99. This train performance coincides with a test accuracy of circa 0.95, resulting in a generalisation gap of circa 0.04.
The teacher model with a higher training loss (seed 2) has significant knowledge transfer, see Table 64, for all functional similarity metrics across alpha values 0.1, 0.5 and 0.9, except for Prediction Disagreement when alpha was 0.1. In this case, we also observe a large asymmetric payoff in prediction agreement, significantly favouring incorrect predictions, Figure 44(a). Whereas teacher seed 0 has a train loss of 0.000061 and has no significant transfer with alpha values of 0.1 and 0.5. However, with an alpha of 0.9, it does have a significant transfer across metrics except for Prediction Disagreement, see Table 64. When alpha is 0.9, we observe an asymmetric payoff in prediction agreement, significantly favouring incorrect predictions. For teacher seed 0, which has a train loss of 0.000646, we observe significant knowledge transfer when alpha is 0.5 and 0.9, coinciding with an asymmetric payoff in prediction agreement, favouring incorrect predictions.
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.000646 | 0.999850 | 0.381410 | 0.951829 |
| 1 | 0.000061 | 0.999973 | 0.331054 | 0.952251 |
| 2 | 0.004657 | 0.998580 | 0.309702 | 0.947104 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.063 | 0.064 | 0.060 | 0.059 | 0.144 | 0.493 | 0.849 |
| Rank Disagreement () | 0.696 | 0.688 | 0.684 | 0.681 | 0.800 | 0.798 | 0.802 |
| Prediction Disagreement () | 0.045 | 0.046 | 0.043 | 0.042 | 0.042 | 0.043 | 0.046 |
| JS Divergence () | 0.025 | 0.025 | 0.023 | 0.022 | 0.053 | 0.201 | 0.431 |
| Accuracy () | 0.952 | 0.951 | 0.954 | 0.954 | 0.957 | 0.957 | 0.955 |
| Loss () | 0.385 | 0.344 | 0.310 | 0.293 | 0.236 | 0.692 | 1.698 |
| Metrics | Control | Knowledge Distillation | Random Knowledge Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.059 | 0.058 | 0.058 | 0.056 | 0.141 | 0.494 | 0.848 |
| Rank Disagreement () | 0.690 | 0.688 | 0.687 | 0.682 | 0.799 | 0.799 | 0.800 |
| Prediction Disagreement () | 0.042 | 0.042 | 0.042 | 0.040 | 0.040 | 0.044 | 0.046 |
| JS Divergence () | 0.023 | 0.023 | 0.022 | 0.022 | 0.052 | 0.201 | 0.431 |
| Accuracy () | 0.953 | 0.953 | 0.953 | 0.954 | 0.958 | 0.954 | 0.953 |
| Loss () | 0.366 | 0.354 | 0.328 | 0.316 | 0.236 | 0.698 | 1.698 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.900 | 0.1 | 0.5 | 0.900 | |
| Activation Distance () | 0.068 | 0.063 | 0.059 | 0.058 | 0.146 | 0.489 | 0.843 |
| Rank Disagreement () | 0.713 | 0.667 | 0.648 | 0.643 | 0.800 | 0.800 | 0.799 |
| Prediction Disagreement () | 0.048 | 0.045 | 0.042 | 0.041 | 0.046 | 0.048 | 0.052 |
| JS Divergence () | 0.026 | 0.023 | 0.021 | 0.020 | 0.053 | 0.199 | 0.427 |
| Accuracy () | 0.952 | 0.955 | 0.957 | 0.957 | 0.956 | 0.957 | 0.953 |
| Loss () | 0.370 | 0.256 | 0.226 | 0.216 | 0.239 | 0.692 | 1.700 |
| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
| KD 0.1 | ✗✗✓ | ✗✗✓ | ✗✗✗ | ✗✗✓ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✗✗✓ | ✓✗✓ | ✗✗✓ | ✓✗✓ | ✗✗✗ | ✗✗✓ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✗✗✓ | ✓✓✓ | ✗✗✗ | ✗✗✓ |
F.3.2 VGG19
Findings:
For the VGG19 on SVHN, we record a low train loss from we observe that the teacher seeds, Table 65, obtain a range of train loss values of 0.004511, 0.002757 and 0.00374 for teacher seeds 0, 1, and 2, respectively. The train accuracies are approximately 0.99. This train performance coincides with a test accuracy of circa 0.95, resulting in a generalisation gap of circa 0.04.
The teacher model with a higher training loss (seed 2) has significant knowledge transfer, see Table 69, for only Rank Disagreement, across alpha values 0.1, 0.5 and 0.9. Due to limited statically significant functional transfer across metrics for this seed, we observe a small but inconsistent asymmetric payoff in prediction agreement, slightly favouring incorrect predictions, Figure 47(a). The story is very similar across the other teacher seeds; we see marginal functional transfer, and where a transfer is higher, we see negative transfer, but where it is marginal or largely insignificant, we see no preference for knowledge transfer, showing that in this case knowledge sharing can not be attributed to improved performance.
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.004511 | 0.998649 | 0.343982 | 0.952827 |
| 1 | 0.002757 | 0.999290 | 0.347466 | 0.948794 |
| 2 | 0.003741 | 0.998935 | 0.313836 | 0.953596 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.065 | 0.064 | 0.066 | 0.065 | 0.151 | 0.494 | 0.848 |
| Rank Disagreement () | 0.708 | 0.660 | 0.637 | 0.603 | 0.799 | 0.812 | 0.805 |
| Prediction Disagreement () | 0.047 | 0.046 | 0.047 | 0.047 | 0.047 | 0.045 | 0.046 |
| JS Divergence () | 0.028 | 0.027 | 0.027 | 0.027 | 0.057 | 0.201 | 0.429 |
| Accuracy () | 0.954 | 0.954 | 0.953 | 0.953 | 0.955 | 0.956 | 0.956 |
| Loss () | 0.349 | 0.292 | 0.282 | 0.275 | 0.263 | 0.698 | 1.696 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.069 | 0.067 | 0.067 | 0.066 | 0.154 | 0.496 | 0.846 |
| Rank Disagreement () | 0.758 | 0.710 | 0.663 | 0.652 | 0.814 | 0.796 | 0.808 |
| Prediction Disagreement () | 0.051 | 0.050 | 0.050 | 0.049 | 0.050 | 0.049 | 0.048 |
| JS Divergence () | 0.030 | 0.029 | 0.029 | 0.028 | 0.058 | 0.201 | 0.428 |
| Accuracy () | 0.952 | 0.953 | 0.953 | 0.954 | 0.953 | 0.955 | 0.956 |
| Loss () | 0.353 | 0.304 | 0.274 | 0.269 | 0.268 | 0.701 | 1.695 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.065 | 0.067 | 0.065 | 0.064 | 0.148 | 0.493 | 0.847 |
| Rank Disagreement () | 0.733 | 0.680 | 0.647 | 0.600 | 0.804 | 0.808 | 0.809 |
| Prediction Disagreement () | 0.048 | 0.049 | 0.047 | 0.046 | 0.045 | 0.044 | 0.046 |
| JS Divergence () | 0.028 | 0.028 | 0.027 | 0.026 | 0.055 | 0.200 | 0.429 |
| Accuracy () | 0.952 | 0.952 | 0.953 | 0.954 | 0.956 | 0.957 | 0.956 |
| Loss () | 0.358 | 0.301 | 0.284 | 0.265 | 0.258 | 0.697 | 1.696 |
| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
| KD 0.1 | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✗✓✗ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✗✓✗ | ✗✗✗ | ✗✗✗ |
| KD 0.9 | ✗✓✗ | ✓✓✓ | ✗✗✗ | ✗✓✗ | ✗✗✗ | ✗✗✗ |
F.3.3 ViT
Findings:
For the ViT on SVHN, we record a train loss from we observe that the teacher seeds, Table 70, obtain a range of train loss values of 0.018473, 0.019402 and 0.018580 for teacher seeds 0, 1, and 2, respectively. The train accuracies are approximately 0.99. This train performance coincides with a test accuracy of circa 0.85, resulting in a generalisation gap of circa 0.14.
The teacher model with a higher training loss (seed 1) has significant knowledge transfer, see Table 74, for only Activation Distance, Rank Disagreement and JS Divergence across alpha values 0.5 and 0.9. In this case, we observe a small but inconsistent asymmetric payoff in prediction agreement, slightly favouring incorrect predictions, Figure 50(a). The story is very similar across the other teacher seeds; we see marginal functional transfer, and where a transfer is higher, we see negative transfer, but where it is marginal or largely insignificant, we see no real preference for knowledge transfer, showing that in this case knowledge sharing can not be attributed to improved performance.
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.018473 | 0.994417 | 0.774354 | 0.854564 |
| 1 | 0.019402 | 0.994963 | 0.711637 | 0.855025 |
| 2 | 0.018580 | 0.994635 | 0.692686 | 0.860633 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.219 | 0.220 | 0.215 | 0.211 | 0.273 | 0.535 | 0.829 |
| Rank Disagreement () | 0.741 | 0.741 | 0.736 | 0.732 | 0.801 | 0.806 | 0.805 |
| Prediction Disagreement () | 0.165 | 0.165 | 0.162 | 0.159 | 0.162 | 0.160 | 0.161 |
| JS Divergence () | 0.0910 | 0.091 | 0.088 | 0.085 | 0.110 | 0.227 | 0.422 |
| Accuracy () | 0.857 | 0.856 | 0.856 | 0.858 | 0.858 | 0.860 | 0.859 |
| Loss () | 0.707 | 0.698 | 0.651 | 0.608 | 0.560 | 0.896 | 1.771 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.216 | 0.212 | 0.208 | 0.206 | 0.266 | 0.529 | 0.825 |
| Rank Disagreement () | 0.745 | 0.745 | 0.737 | 0.735 | 0.801 | 0.805 | 0.804 |
| Prediction Disagreement () | 0.162 | 0.159 | 0.157 | 0.156 | 0.158 | 0.156 | 0.164 |
| JS Divergence () | 0.089 | 0.086 | 0.084 | 0.082 | 0.106 | 0.224 | 0.420 |
| Accuracy () | 0.856 | 0.861 | 0.863 | 0.864 | 0.863 | 0.865 | 0.854 |
| Loss () | 0.722 | 0.680 | 0.603 | 0.574 | 0.543 | 0.886 | 1.777 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.212 | 0.206 | 0.206 | 0.204 | 0.265 | 0.532 | 0.828 |
| Rank Disagreement () | 0.742 | 0.735 | 0.731 | 0.728 | 0.802 | 0.803 | 0.804 |
| Prediction Disagreement () | 0.160 | 0.155 | 0.155 | 0.153 | 0.156 | 0.153 | 0.152 |
| JS Divergence () | 0.087 | 0.084 | 0.083 | 0.081 | 0.106 | 0.225 | 0.421 |
| Accuracy () | 0.856 | 0.861 | 0.859 | 0.860 | 0.863 | 0.866 | 0.864 |
| Loss () | 0.730 | 0.673 | 0.627 | 0.600 | 0.548 | 0.886 | 1.768 |
| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
| KD 0.1 | ✗✓✓ | ✗✗✓ | ✗✗✗ | ✗✓✓ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✗✓✓ | ✓✓✓ | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
Appendix G Audio Results
Training Settings:
All audio is converted into mono and downsampled to 16000 htz, it is converted into a spectrogram using torchaudio (Hwang et al., 2023) with an n_fft of 512 and a power of 2. This is then converted to the MelScale with an n_mels of 32 and a sample rate of 16000 and a n_stft of 257.
The train test split for Urbansounds8K used sklearn (Pedregosa et al., 2011) train_test_split function with a test size of 0.2 a random state of 42 and the shuffle set to True.
All audio architectures are trained with SGD optimiser with a learning rate of 0.01 and a batch size of 256 for 100 epochs on SpeechCommandsV2 and 150 epochs for UrbanSounds8K. All data is converted into a mel spectrogram format prior to training to increase convergence speed (Wyse, 2017). The audio architectures are trained with the same seeds and data orders from seeds 10-19 for the 10 models used for averaging. This is repeated for the three teachers trained on seeds 0-2.
G.1 SpeechCommands
SpeechCommands (Warden, 2017) is an audio dataset comprised of 35 classes with 29.4 hours of audio clips of a 1-2 second duration. There are 84,843 training examples and 11,005 testing examples.
Findings:
We find that for SpeechCommands that knowledge transfer is significant allowing the rejection of the null hypothesis for knowledge sharing. For both architectures there is considerable knowledge transfer compared to the baseline controls. We also find that there is asymmetric knowledge transfer with a weighting towards negative knowledge transfer.
G.1.1 VGGish
Findings:
We observe that the teacher model achieves a high train accuracy along with a high train loss, see Table 75. With this we observe a substantial and statistically significant knowledge transfer for all alpha values, see Tables 76, 77, 78 and 79.This substantial and significant transfer of knowledge, as expected, coincides with a strong asymmetric transfer of knowledge favouring incorrect predictions, as shown in Figure 53(a).
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.044291 | 0.986457 | 0.817567 | 0.879237 |
| 1 | 0.061635 | 0.981566 | 0.928225 | 0.864698 |
| 2 | 0.043880 | 0.987047 | 0.765199 | 0.877328 |
| Metrics | Baseline | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.190 | 0.152 | 0.148 | 0.147 | 0.260 | 0.570 | 0.877 |
| Rank Disagreement () | 0.908 | 0.885 | 0.880 | 0.878 | 0.942 | 0.942 | 0.939 |
| Prediction Disagreement () | 0.144 | 0.118 | 0.114 | 0.114 | 0.125 | 0.133 | 0.169 |
| JS Divergence () | 0.085 | 0.063 | 0.060 | 0.059 | 0.120 | 0.274 | 0.512 |
| Accuracy () | 0.870 | 0.886 | 0.887 | 0.884 | 0.892 | 0.882 | 0.844 |
| Loss () | 1.076 | 0.669 | 0.564 | 0.553 | 0.565 | 1.103 | 2.366 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.209 | 0.169 | 0.168 | 0.165 | 0.277 | 0.579 | 0.881 |
| Rank Disagreement () | 0.910 | 0.885 | 0.881 | 0.879 | 0.942 | 0.942 | 0.940 |
| Prediction Disagreement () | 0.157 | 0.129 | 0.127 | 0.125 | 0.139 | 0.149 | 0.181 |
| JS Divergence () | 0.094 | 0.071 | 0.068 | 0.066 | 0.129 | 0.281 | 0.515 |
| Accuracy () | 0.868 | 0.882 | 0.883 | 0.882 | 0.889 | 0.880 | 0.842 |
| Loss () | 1.051 | 0.675 | 0.572 | 0.559 | 0.576 | 1.111 | 2.375 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.192 | 0.151 | 0.149 | 0.148 | 0.260 | 0.572 | 0.877 |
| Rank Disagreement () | 0.908 | 0.885 | 0.880 | 0.878 | 0.942 | 0.942 | 0.940 |
| Prediction Disagreement () | 0.145 | 0.117 | 0.116 | 0.115 | 0.126 | 0.135 | 0.166 |
| JS Divergence () | 0.085 | 0.062 | 0.060 | 0.059 | 0.120 | 0.276 | 0.511 |
| Accuracy () | 0.870 | 0.887 | 0.889 | 0.889 | 0.892 | 0.882 | 0.847 |
| Loss () | 1.086 | 0.629 | 0.531 | 0.516 | 0.562 | 1.111 | 2.363 |
| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
| KD 0.1 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✓ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✓✓✓ |
G.1.2 ViT
Findings:
We observe that the teacher model achieves a high train accuracy along with a high train loss, see Table 80. With this we observe a substantial and statistically significant knowledge transfer for all alpha values, see Tables 81, 82, 83 and 84.This substantial and significant transfer of knowledge, as expected, coincides with a strong asymmetric transfer of knowledge favouring incorrect predictions, as shown in Figure 53(a).
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.013776 | 0.996440 | 1.001014 | 0.833530 |
| 1 | 0.002471 | 0.999352 | 0.925219 | 0.853794 |
| 2 | 0.003337 | 0.999163 | 0.913119 | 0.853430 |
| Metrics | Basline | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.164 | 0.133 | 0.123 | 0.118 | 0.245 | 0.561 | 0.870 |
| Rank Disagreement () | 0.852 | 0.825 | 0.810 | 0.803 | 0.937 | 0.940 | 0.939 |
| Prediction Disagreement () | 0.124 | 0.101 | 0.094 | 0.090 | 0.136 | 0.154 | 0.181 |
| JS Divergence () | 0.062 | 0.045 | 0.039 | 0.036 | 0.109 | 0.271 | 0.512 |
| Accuracy () | 0.843 | 0.842 | 0.844 | 0.844 | 0.856 | 0.852 | 0.826 |
| Loss () | 1.094 | 0.990 | 0.835 | 0.791 | 0.687 | 1.161 | 2.408 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.143 | 0.129 | 0.119 | 0.115 | 0.227 | 0.558 | 0.874 |
| Rank Disagreement () | 0.844 | 0.833 | 0.821 | 0.814 | 0.935 | 0.939 | 0.938 |
| Prediction Disagreement () | 0.107 | 0.097 | 0.090 | 0.087 | 0.113 | 0.138 | 0.162 |
| JS Divergence () | 0.053 | 0.045 | 0.040 | 0.038 | 0.100 | 0.266 | 0.512 |
| Accuracy () | 0.849 | 0.854 | 0.854 | 0.855 | 0.863 | 0.858 | 0.835 |
| Loss () | 1.071 | 0.994 | 0.941 | 0.900 | 0.656 | 1.138 | 2.394 |
| Metric | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.152 | 0.139 | 0.131 | 0.126 | 0.232 | 0.560 | 0.875 |
| Rank Disagreement | 0.852 | 0.844 | 0.833 | 0.826 | 0.936 | 0.939 | 0.938 |
| Prediction Disagreement | 0.115 | 0.105 | 0.100 | 0.096 | 0.122 | 0.141 | 0.163 |
| JS Divergence | 0.058 | 0.051 | 0.046 | 0.043 | 0.102 | 0.267 | 0.512 |
| Accuracy | 0.852 | 0.857 | 0.856 | 0.857 | 0.860 | 0.852 | 0.827 |
| Loss | 1.027 | 0.955 | 0.897 | 0.860 | 0.661 | 1.152 | 2.398 |
| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
| KD 0.1 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
G.2 UrbanSound8K
UrbanSound8K is a large event classification dataset that contains 18.5 hours of annotated sound event occurrences across 10 classes (Salamon et al., 2014). It has 6,985 training set instances and 1,747 testing set instances which are between 0 and 4 seconds in duration.
Findings:
We find that for UrbanSound8K knowledge transfer is significant allowing the rejection of the null hypothesis for knowledge sharing. For both the VGG architecture there is considerable knowledge transfer compared to the baseline controls, but for the transformer architecture there is only marginal knowledge transfer. We also find that there is asymmetric knowledge transfer with a weighting towards negative knowledge transfer when the knowledge transfer is statistically significant and considerable.
G.2.1 VGGish
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.013431 | 0.994989 | 2.203087 | 0.797939 |
| 1 | 0.014136 | 0.994560 | 2.405788 | 0.785346 |
| 2 | 0.151926 | 0.947173 | 1.568569 | 0.702919 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.256 | 0.267 | 0.242 | 0.243 | 0.354 | 0.597 | 0.873 |
| Rank Disagreement () | 0.696 | 0.696 | 0.683 | 0.678 | 0.795 | 0.791 | 0.784 |
| Prediction Disagreement () | 0.192 | 0.196 | 0.180 | 0.180 | 0.187 | 0.195 | 0.387 |
| JS Divergence () | inf, nan | inf, nan | 0.099 | 0.100 | 0.149 | 0.268 | 0.467 |
| Accuracy () | 0.795 | 0.787 | 0.796 | 0.796 | 0.808 | 0.806 | 0.585 |
| Loss () | 2.813 | 2.460 | 2.225 | 2.089 | 0.730 | 1.085 | 2.059 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.363 | 0.284 | 0.262 | 0.264 | 0.367 | 0.600 | 0.871 |
| Rank Disagreement | 0.730 | 0.718 | 0.706 | 0.703 | 0.798 | 0.792 | 0.784 |
| Prediction Disagreement | 0.272 | 0.214 | 0.197 | 0.199 | 0.208 | 0.218 | 0.387 |
| JS Divergence | inf, nan | inf, nan | inf, nan | inf, nan | 0.156 | 0.269 | 0.465 |
| Accuracy | 0.724 | 0.782 | 0.791 | 0.791 | 0.806 | 0.796 | 0.589 |
| Loss | 2.046 | 3.056 | 2.34 | 2.235 | 0.748 | 1.093 | 2.054 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.396 | 0.357 | 0.335 | 0.324 | 0.416 | 0.590 | 0.821 |
| Rank Disagreement | 0.745 | 0.712 | 0.692 | 0.683 | 0.812 | 0.806 | 0.801 |
| Prediction Disagreement | 0.295 | 0.274 | 0.260 | 0.253 | 0.292 | 0.293 | 0.438 |
| JS Divergence | 0.167 | 0.141 | 0.127 | 0.120 | 0.175 | 0.264 | 0.433 |
| Accuracy | 0.794 | 0.789 | 0.791 | 0.776 | 0.810 | 0.808 | 0.577 |
| Loss | 3.209 | 1.106 | 0.944 | 0.961 | 0.716 | 1.080 | 2.065 |
| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
| KD 0.1 | ✗✓✓ | ✗✗✓ | ✗✗✓ | ✓✗✓ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
G.2.2 ViT
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.000180 | 1.000000 | 1.638960 | 0.772753 |
| 1 | 0.000375 | 0.999857 | 1.583644 | 0.768746 |
| 2 | 0.000168 | 1.000000 | 1.593121 | 0.781912 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.098 | 0.098 | 0.096 | 0.097 | 0.287 | 0.592 | 0.854 |
| Rank Disagreement () | 0.423 | 0.419 | 0.417 | 0.415 | 0.755 | 0.773 | 0.759 |
| Prediction Disagreement () | 0.074 | 0.072 | 0.073 | 0.073 | 0.131 | 0.174 | 0.252 |
| JS Divergence () | 0.025 | 0.025 | 0.024 | 0.025 | 0.111 | 0.262 | 0.448 |
| Accuracy () | 0.771 | 0.771 | 0.771 | 0.772 | 0.788 | 0.806 | 0.719 |
| Loss () | 1.628 | 1.621 | 1.585 | 1.560 | 0.748 | 1.095 | 1.956 |
| Metrics | Control | Knowledge Distillation | Rand Knowledge Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance | 0.109 | 0.108 | 0.108 | 0.105 | 0.291 | 0.592 | 0.854 |
| Rank Disagreement | 0.442 | 0.44 | 0.429 | 0.427 | 0.756 | 0.769 | 0.763 |
| Prediction Disagreement | 0.078 | 0.077 | 0.077 | 0.073 | 0.130 | 0.173 | 0.261 |
| JS Divergence | 0.029 | 0.029 | 0.028 | 0.027 | 0.113 | 0.262 | 0.448 |
| Accuracy | 0.768 | 0.768 | 0.770 | 0.769 | 0.794 | 0.811 | 0.716 |
| Loss | 1.589 | 1.584 | 1.532 | 1.509 | 0.735 | 1.096 | 1.959 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.099 | 0.100 | 0.100 | 0.101 | 0.288 | 0.598 | 0.859 |
| Rank Disagreement () | 0.413 | 0.414 | 0.410 | 0.425 | 0.754 | 0.770 | 0.759 |
| Prediction Disagreement () | 0.071 | 0.071 | 0.068 | 0.072 | 0.130 | 0.171 | 0.257 |
| JS Divergence () | 0.026 | 0.026 | 0.026 | 0.027 | 0.111 | 0.265 | 0.451 |
| Accuracy () | 0.786 | 0.784 | 0.783 | 0.783 | 0.801 | 0.812 | 0.719 |
| Loss () | 1.539 | 1.538 | 1.508 | 1.484 | 0.716 | 1.091 | 1.959 |
| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
| KD 0.1 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✗✗✗ | ✗✓✗ | ✗✗✗ | ✗✓✗ | ✗✗✗ | ✗✗✗ |
| KD 0.9 | ✗✓✗ | ✓✓✗ | ✗✓✗ | ✗✓✗ | ✗✗✗ | ✗✗✗ |
Appendix H Language Results
H.1 Tiny Shakespeare Dataset
Training Settings:
The language model was a GPT2-style transformer with an embedding dimension of 384, a vocabulary size of 65, six attention heads, six transformer blocks, a dropout of 0.200, and a block size of 256. It was trained on the Tiny Shakespeare dataset, with the first 90% used for training and the last 10% used for testing. The dataset was tokenised via a character tokenizer, and the model was trained auto-regressively to predict the next character token. The model was trained with the Adam optimiser with a learning rate of 3e-4 with a batch size of 64 for 5000 iterations. The student models are trained with the same seeds and data orders from seeds 10 to 19 for the 10 models used for averaging. This is repeated for the three teachers trained on seeds 0 to 2.
Justification:
This setup allows for a fair analysis of Knowledge Distillation as its role is isolated in the training process. Other than the architecture’s implicit bias towards the problem, which affects its performance (loss and accuracy), there are no confounding factors that could influence Knowledge Distillation.
Findings:
We observe a high train loss for the teacher model circa 0.86 with a high train accuracy circa 0.72, see Table 95. This high train loss, corresponds as expected with a substational and significant knowledge transfer which incresae as alpha increases, see Tables 96, 97, 98 and 99. This substational and significant knowledge transfer coincides with with an asymmetric payoff in prediction agreement, strongly favouring incorrect predictions, see Figure 65(a). This result is as expected from the results and intuition presented in the results of the main body of the paper.
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.864641 | 0.719685 | 1.567481 | 0.573366 |
| 1 | 0.866370 | 0.719697 | 1.561079 | 0.574668 |
| 2 | 0.861098 | 0.721140 | 1.562137 | 0.573033 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.196 | 0.187 | 0.158 | 0.144 | 0.204 | 0.378 | 0.661 |
| Rank Disagreement () | 0.910 | 0.907 | 0.897 | 0.891 | 0.944 | 0.947 | 0.950 |
| Prediction Disagreement () | 0.246 | 0.236 | 0.200 | 0.182 | 0.242 | 0.243 | 0.255 |
| JS Divergence () | 0.053 | 0.049 | 0.037 | 0.032 | 0.067 | 0.192 | 0.449 |
| Accuracy () | 0.574 | 0.577 | 0.583 | 0.581 | 0.576 | 0.578 | 0.570 |
| Loss () | 1.559 | 1.542 | 1.496 | 1.500 | 1.507 | 1.839 | 2.995 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.195 | 0.185 | 0.156 | 0.141 | 0.201 | 0.370 | 0.653 |
| Rank Disagreement () | 0.910 | 0.907 | 0.897 | 0.891 | 0.944 | 0.946 | 0.950 |
| Prediction Disagreement () | 0.249 | 0.238 | 0.202 | 0.183 | 0.245 | 0.245 | 0.263 |
| JS Divergence () | 0.052 | 0.048 | 0.036 | 0.031 | 0.066 | 0.190 | 0.446 |
| Accuracy () | 0.574 | 0.577 | 0.584 | 0.582 | 0.577 | 0.577 | 0.568 |
| Loss () | 1.559 | 1.539 | 1.488 | 1.493 | 1.504 | 1.840 | 2.997 |
| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
| Activation Distance () | 0.195 | 0.186 | 0.157 | 0.142 | 0.202 | 0.372 | 0.658 |
| Rank Disagreement () | 0.909 | 0.906 | 0.896 | 0.89 | 0.944 | 0.946 | 0.950 |
| Prediction Disagreement () | 0.245 | 0.233 | 0.198 | 0.180 | 0.241 | 0.240 | 0.256 |
| JS Divergence () | 0.052 | 0.048 | 0.037 | 0.031 | 0.066 | 0.190 | 0.448 |
| Accuracy () | 0.574 | 0.577 | 0.583 | 0.582 | 0.577 | 0.578 | 0.570 |
| Loss () | 1.558 | 1.536 | 1.493 | 1.493 | 1.504 | 1.834 | 2.996 |
| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
| KD 0.1 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ |
| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ |
H.2 Tiny Shakespeare Dataset Adversarial Attack
Training Settings:
We train an adversarial teacher that has every occurrence of ‘t’ ‘h’ ‘e’ replaced with ‘t’ ‘h’ ‘a’ in its training set, given the zipfs law of the dataset, Table 100, we can see ‘e’ is the most likely character after ‘SPACE’ therefore if adversarial transfer is possible via knowledge transfer a student trained with the adversarial teacher should predict ‘t’ ‘h’ ‘a’ more than ‘t’ ‘h’ ‘e’ when compared to the controls model trained without the teacher. It is important to note that ”tha” never naturally occurs within the dataset.
Justification:
Provided we observe asymmetric knowledge of incorrect knowledge from the teacher to the student, we use this experimental setup to highlight the safety concerns of using Knowledge Distillation. In this case, the teacher has a known vulnerability and has been poisoned to predict an incorrect token. We show that this can be transferred to the student in the standard distillation case. Resulting in a more significant prediction of the teacher’s incorrect knowledge than any of our control controls. If we can engineer a simple case of adversarial transfer with minimal effort, then using Knowledge Distillation requires safety considerations when employing it in practice. Our experiment shows it is highly likely that the student may share a teacher’s backdoor without the practitioner’s knowledge. Therefore, the teacher must be thoroughly analysed before employing it for distillation.
Findings:
| Character | Space | e | t | o | a | h | s | r | n | … |
| Frequency | 0.1523 | 0.0848 | 0.0601 | 0.059 | 0.0498 | 0.046 | 0.0446 | 0.0438 | 0.0435 | … |
| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
| 0 | 0.968203 | 0.698038 | 1.641436 | 0.562150 |
| 1 | 0.974442 | 0.696534 | 1.630169 | 0.562769 |
| 2 | 0.958430 | 0.700257 | 1.631381 | 0.561225 |
We show that the transfer occurs for student models across alpha values with increasing severity for increased alpha values. Therefore, we further substantiate the claim that safety is an important factor to consider due to adversarial transfer in Knowledge Distillation, as shown by the increase in prediction of ’t”h”a’ compared to the controls in Tables 102, 103 and 104.
| Control | Knowledge Distillation | Random Control Distillation | ||||||
| Predicted Word | Teacher | SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 |
| tha | 454 | 105.9 4.1676 | 106.0 3.0463 | 199.1 13.3914 | 436.2 7.9835 | 104.6 3.8967 | 114.8 3.0555 | 126.9 8.0678 |
| the | 285 | 665.1 7.6752 | 675.5 10.2277 | 583.4 17.5364 | 343.6 6.3580 | 668.8 12.7128 | 712.5 12.4798 | 826.3 20.2025 |
| Control | Knowledge Distillation | Random Control Distillation | ||||||
| Predicted Word | Teacher | SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 |
| tha | 534 | 110.5 3.9881 | 115.7 3.6416 | 236.8 11.7761 | 517.8 12.7733 | 112.6 3.4035 | 119.6 3.8215 | 127.4 3.9044 |
| the | 273 | 683.7 15.4370 | 691.4 13.3156 | 599.7 13.8564 | 325.4 7.5262 | 684.7 14.5781 | 733.9 13.4428 | 869.8 10.8109 |
| Control | Knowledge Distillation | Random Control Distillation | ||||||
| Predicted Word | Teacher | SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 |
| tha | 513 | 111.9 4.0236 | 116.1 3.3300 | 241.5 8.5032 | 518.6 11.6612 | 114.7 6.5636 | 114.3 3.9320 | 124.5 4.7943 |
| the | 266 | 656.0 16.0244 | 677.0 13.9743 | 558.0 14.9513 | 303.5 7.7424 | 672.1 18.5513 | 715.0 12.5825 | 836.7 17.1954 |
Appendix I Compute Usage
All models were trained on a A100 GPUs, assuming that the approximate time to train and evaluate a model takes 0.5 hours, to run one condition with three teacher seeds and 10 students models it would take 109.5 hours if run sequentially. Therefore, the whole paper would take 1095 hours for the 10 conditions explored in an sequential setting.
Appendix J Dataset Licences
Image Datasets
Audio Datasets
UrbanSound8K (Salamon et al., 2014) has a Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) license (https://www.kaggle.com/datasets/chrisfilo/urbansound8k)licenece..)
Speech Commands (Warden, 2017) License is CC BY. This license enables reusers to distribute, remix, adapt, and build upon the material in any medium or format, so long as attribution is given to the creator. The license allows for commercial use. (https://paperswithcode.com/dataset/speech-commands).
Language Datasets
Tiny Shakespeare (Karpathy, 2015) has an MIT Licence.

### v2

@@
-Rethinking Knowledge Distillation: A Data Dependent Regulariser With a Negative Asymmetric Payoff
+A Functional Perspective on Knowledge Distillation in Neural Networks
 Abstract
-Knowledge distillation is often considered a compression mechanism when judged on the resulting student’s accuracy and loss, yet its functional impact is poorly understood.
-In this work, we quantify the compression capacity of knowledge distillation and the resulting knowledge transfer from a functional perspective, decoupling compression from architectural reduction, which provides an improved understanding of knowledge distillation.
-We employ hypothesis testing, controls, and random control distillation to understand knowledge transfer mechanisms across data modalities.
-To rigorously test the breadth and limits of our analyses, we explore multiple distillation variants and analyse distillation scaling laws across model sizes.
-Our findings demonstrate that, while there is statistically significant knowledge transfer in some modalities and architectures, the extent of this transfer is less pronounced than anticipated, even under conditions designed to maximise knowledge sharing.
-Notably, in cases of significant knowledge transfer, we identify a consistent and severe asymmetric transfer of negative knowledge to the student, raising safety concerns in knowledge distillation applications.
-Across 12 experimental setups, 9 architectures, and 7 datasets, our findings show that knowledge distillation functions less as a compression mechanism and more as a data-dependent regulariser with a negative asymmetric payoff.
+Knowledge distillation is considered a compression mechanism when judged on the resulting student’s accuracy and loss, yet its functional impact is poorly understood.
+We quantify the compression capacity of knowledge distillation and the resulting knowledge transfer from a functional perspective, decoupling compression from architectural reduction to provide an improved understanding of knowledge distillation.
+We employ a control-driven experimental protocol with hypothesis testing and random control distillation to isolate and understand knowledge transfer mechanisms across data modalities.
+To test the breadth and limits of our analyses, we study self-distillation, standard distillation, feature-map matching variants, distillation scaling laws across model sizes, and the impact of temperature on knowledge transfer.
+We find statistically supported knowledge transfer in some modalities and architectures; however, the extent of this transfer is less pronounced than anticipated, even under conditions that maximise knowledge sharing.
+Notably, in cases of significant functional transfer, we identify a consistent and severe asymmetric transfer of negative knowledge to the student, raising safety concerns for knowledge distillation.
+Across 22 experimental setups, 9 architectures, and 7 datasets, our results suggest that knowledge distillation functions less as a robust compression-by-transfer mechanism and more as a data-dependent regulariser whose transfer component is biased towards negative asymmetric transfer.
+Keywords: Knowledge Distillation, Functional Similarity, Asymmetric Transfer, Distillation Scaling Laws, Adversarial Transfer
 1 Introduction
@@
 Knowledge distillation (KD) has emerged as a widely adopted method to compress models by training a student model to mimic a larger teacher (Buciluǎ et al., 2006; Hinton et al., 2015; Gu et al., 2024; Muralidharan et al., 2024).
-While KD can be applied across architectures and modalities – including in self-distillation regimes where the teacher and student share the same architecture (Allen-Zhu & Li, 2023; Zhang et al., 2019) – the mechanism by which KD improves student performance remains unknown (Busbridge et al., 2025).
-Recent studies have challenged the assumption that KD works through meaningful knowledge transfer, showing that performance gains have been observed even with randomly initialised teachers (Stanton et al., 2021a) motivating a rigorous examination of KD’s functional impact.
-In this work, we move beyond the question of whether knowledge is transferred – we challenge the framing of Knowledge Distillation as a mechanism of knowledge transfer altogether.
-We argue that the improvements observed do not necessarily arise from meaningful transfer of the teacher’s knowledge, but from a more general, data-dependent regularisation effect disputed in literature (Stanton et al., 2021a; Yun et al., 2020; Ge et al., 2021; Yuan et al., 2020) with a novel identification of a negative asymmetric payoff in KD.
-To support this claim, we study KD from a functional perspective, and quantify how closely student models replicate the teacher’s output function.
-We ground our work around two research questions: 1) Does knowledge distillation result in a significantly functionally similar model to the teacher across architectures and data domains against controls?
-2) What knowledge, if any, is actually transferred to student models?
+While KD can be applied across architectures and modalities, including in self-distillation regimes where the teacher and student share the same architecture (Allen-Zhu and Li, 2023; Zhang et al., 2019), the mechanism by which KD improves student performance remains contested (Busbridge et al., 2025).
+Recent studies have challenged the assumption that KD works through meaningful knowledge transfer, showing that performance gains have been observed even with randomly initialised teachers (Stanton et al., 2021).
+This motivates a rigorous examination of KD’s functional impact.
+In this work, we move beyond the question of whether knowledge is transferred: we re-examine the framing of knowledge distillation as a robust mechanism of knowledge transfer.
+Our contribution is not only to examine the regularisation view of KD, but to make it testable: we introduce a functional, control-driven evaluation protocol that measures teacher-student functional similarity relative to counterfactual baselines and decomposes transferred agreement into teacher-correct and teacher-incorrect regions.
+This lets us distinguish apparent KD gains that are consistent with data-dependent regularisation from gains accompanied by measurable teacher-function transfer.
+Throughout the paper, ‘knowledge transfer’ refers to our operational notion of teacher-function transfer, measured relative to counterfactual controls.
+Our study is grounded in two research questions: 1) Does knowledge distillation produce statistically supported teacher-student functional similarity, relative to control conditions, across architectures and data modalities?
+2) When functional transfer occurs, is the resulting teacher-student alignment symmetric across teacher-correct and teacher-incorrect behaviour, or is it systematically asymmetric?
+Across our experiments, we find that improvements under KD do not necessarily arise from meaningful transfer of the teacher’s knowledge, but are often consistent with a substantial data-dependent regularisation effect (Stanton et al., 2021; Yun et al., 2020; Ge et al., 2021; Yuan et al., 2020).Crucially, when functional transfer is statistically supported, we identify a systematic negative asymmetric payoff: KD disproportionately increases agreement on the teacher’s incorrect predictions relative to agreement on its correct predictions.
 We first focus on self-distillation, where the student has the capacity to match the teacher’s functional representation perfectly, ensuring that any observed differences are solely due to the distillation signal.
-We then verify our findings in the standard distillation setting with smaller student models (Appendix Section E), as well as with different KD variants in Appendix Section C.
-Our methodological framework isolates the core mechanics of Knowledge Distillation through: 1) a controlled training setup where all models share initialisation, enabling precise functional comparison; 2) two controls: independent models with the same architecture,initialisation and different data order (SIDDO) as the teacher, and a Random Control Distillation (RCD) where students are trained using uniform noise in place of teacher outputs, all functionally compared to the teacher model used in the standard distillation process; 3) functional similarity metrics including Activation Distance, Rank Disagreement, Prediction Disagreement, JS Divergence and Prediction Agreement.
-We conduct experiments across 7 datasets, 3 data modalities (image, audio, and language), and 9 architectures, training over 3,900 models. Our findings show that:
-While KD can lead to statistically significant functional similarity between teacher and student, this similarity is often marginal and inconsistent across datasets and modalities.
-The most substantial improvements in accuracy and loss frequently arise under Random Control Distillation, challenging the assumption that performance gains reflect successful knowledge transfer.
-When knowledge transfer is significant and not marginal, the transferred knowledge has an asymmetric weighting towards the teacher’s incorrect predictions.
-This asymmetry becomes more pronounced as dependence on the teacher increases.
-Our findings compel a re-characterisation of KD, not as a robust knowledge transfer mechanism, but as a data-dependent regulariser with inconsistent and negative asymmetric knowledge-sharing capacity.
-This perspective raises important safety concerns: when knowledge transfer is significant, KD may amplify incorrect or harmful behaviour encoded in the teacher.
-We present a concrete case of adversarial transfer facilitated by KD to support this.
+We then verify our findings in the standard distillation setting with smaller student models (Section 4.2), as well as with different KD variants (Section 4.3).
+Our methodological framework isolates the core mechanics of knowledge distillation through: 1) a controlled training setup where all models share initialisations and data orders, enabling precise functional comparison; 2) two counterfactual controls: independent models with the same initialisation seed and different data order (SIDDO) as the teacher (using the same architecture in the self-distillation case), and a Random Control Distillation (RCD) where the distillation target is replaced with a class-uniform distribution, all functionally compared to the teacher model used in the standard distillation process; 3) a broad array of functional similarity metrics including Activation Distance, Rank Disagreement, Prediction Disagreement, JS Divergence and Prediction Agreement to quantify teacher-student functional alignment and functional transfer between each control and the teacher.
+We conduct experiments across 7 datasets, 3 data modalities (image, audio, and language), and 9 architectures, training over 4,000 models.
+Our findings show that:
+The strongest improvements in accuracy and loss frequently arise under Random Control Distillation, challenging the assumption that KD performance gains necessarily reflect successful knowledge transfer, and supporting a data-dependent regularisation perspective of KD.
+KD can lead to statistically supported functional similarity between teacher and student, but this similarity is often marginal and inconsistent across datasets and modalities.
+Furthermore, increasing the distillation temperature significantly reduces the degree of functional transfer.
+When functional transfer is supported and not marginal, the resulting teacher-student alignment is often asymmetric, with disproportionate agreement on the teacher’s incorrect predictions.
+This asymmetry becomes more pronounced as dependence on the teacher signal increases.
+In the regimes we study, our findings motivate a re-characterisation of KD, not as a uniformly robust knowledge transfer mechanism, but as a data-dependent regulariser with inconsistent and negative asymmetric knowledge-sharing capacity.
+This perspective raises important, practical safety concerns: when functional transfer is significant, KD may amplify incorrect or harmful behaviour encoded in the teacher.
+We further support this claim via a concrete case of adversarial transfer facilitated by KD.
 Concretely, our contributions are as follows:
-Introduce a functional framework to analyse KD beyond accuracy and loss, but as a process where internal knowledge transfer dynamics can be quantitatively measured.
-Isolate the contribution of the teacher signal using strong statistical and control-based methodology, something that prior work has not quantitatively disentangled to this level.
-Identify and characterise a novel phenomenon across conditions, modalities and architectures: when functional transfer occurs, it disproportionately favours the teacher’s incorrect predictions, revealing a systematic error amplification effect with safety implications.
-Demonstrate the diagnostic utility of RCD as a crucial counterfactual, showing it frequently outperforms KD, undermining assumptions about knowledge transfer.
-Conduct the largest multimodal functional study of KD to date.
-Our empirical analysis spans over 3,900 trained models across 9 architectures, 7 datasets, and 3 modalities (vision, audio, and language), establishing the generality and reproducibility of our claims.
-Reveal targeted and scalable negative transfer via adversarial and capacity scaling experiments.
-We show that KD can reliably copy specific erroneous behaviours, and that this error amplification scales with model capacity, underscoring the hidden risks of KD in high-stakes settings.
+We provide a control-driven agnostic experimental framework that isolates the contribution of the teacher signal and evaluates KD gains against counterfactual baselines rather than performance metrics alone, something that prior work has not quantitatively disentangled to this level.
+We identify and characterise a novel phenomenon across conditions, modalities and architectures: when statistically supported functional transfer occurs, KD disproportionately favours the teacher’s incorrect predictions, revealing a systematic error amplification effect with safety implications.
+We demonstrate the diagnostic utility of RCD as a crucial counterfactual, showing it frequently outperforms KD, concretely weakening the assumption that KD gains primarily arise from knowledge transfer.
+To our knowledge, we conduct the largest functional study of KD to date spanning over 4,000 trained models across 9 architectures, 7 datasets, and 3 modalities (vision, audio, and language), supporting the generality of our empirical observations and reproducibility of our evaluation framework.
+We show targeted and scalable negative transfer via adversarial and capacity-scaling experiments, demonstrating that KD can reliably copy specific erroneous behaviours, and that negative asymmetric transfer can increase with student capacity – underscoring the safety implications of KD applications in high-stakes settings.
+Our study contributes a timely and nuanced perspective on knowledge distillation by turning competing explanations of KD into an empirically testable question, helping to adjudicate between data-dependent regularisation and reliable knowledge-transfer accounts, while uncovering a systematic error-asymmetric transfer dynamic with important safety implications across modalities.
 2 Related Work
 Knowledge Distillation (KD):
-KD transfers behaviour from a teacher (or ensemble) into a student (Buciluǎ et al., 2006; Hinton et al., 2015), with strong empirical results across modalities (Beyer et al., 2022; Jung et al., 2020; Sanh, 2019; Aghli & Ribeiro, 2021; Li et al., 2020; Fang et al., 2021; Wang et al., 2022) and architectures (Touvron et al., 2021; Miles et al., 2024).
-Yet the role of knowledge transfer is debated (Mason-Williams, 2024; Stanton et al., 2021b; Ojha et al., 2023; Menon et al., 2021).
-Prior work alternately views KD as a regulariser (Yun et al., 2020; Ge et al., 2021; Yuan et al., 2020) or argues against that view (Shen et al., 2021; Sultan, 2023).
-In this paper, we advance the discussion surrounding KD as a regulariser with a functional perspective that spans image, audio, and language.
-We present a control-driven functional protocol that decouples compression from size, measures alignment beyond accuracy, confirming KD acts as a data-dependent regulariser but exposing a new dimension of this regularisation with respect to its systematic negative transfer to the student.
+KD transfers behaviour from a teacher (or ensemble) to a student (Buciluǎ et al., 2006; Hinton et al., 2015), with strong empirical results across modalities (Beyer et al., 2022; Jung et al., 2020; Sanh, 2019; Aghli and Ribeiro, 2021; Li et al., 2020; Fang et al., 2021; Wang et al., 2022) and architectures (Touvron et al., 2021; Miles et al., 2024).
+However, the mechanism behind KD improvements remains debated (Mason-Williams, 2024; Stanton et al., 2021; Ojha et al., 2023; Menon et al., 2021).
+A line of work interprets KD as a form of data-dependent regularisation (Yun et al., 2020; Ge et al., 2021; Yuan et al., 2020), arguing that limited or no knowledge is transferred under its optimisation objective.
+Other literature, on the other hand, strongly argues that KD can effectively transfer meaningful knowledge from teacher to student (Shen et al., 2021; Sultan, 2023).
+In this paper, we advance the discussion of KD as a regulariser by introducing a functional perspective that spans vision, audio, and language.
+There remains substantial contention in the community about how knowledge distillation fundamentally operates (Busbridge et al., 2025), with misaligned interpretations across data modalities.
+This motivates our multi-modal, multi-architecture, and multi-dataset study of KD.
+We present a control-driven functional protocol that decouples compression effects from architectural size reduction, and measures teacher-student alignment for statistical significance beyond only analysing accuracy and loss.
+Using this framework, we provide evidence that KD often acts as a data-dependent regulariser, while also exposing a new dimension of this regularisation: when substantial and statistically supported knowledge transfer occurs, there is a strong negative asymmetric transfer to the student.
 Functional Similarity Metrics:
-Functional similarity compares models by their outputs rather than only their accuracy (Klabunde et al., 2023).
-It has been used for unlearning (Golatkar et al., 2021; Chundawat et al., 2023), ensemble dynamics (Fort et al., 2019), and compression/pruning (Mason-Williams & Dahlqvist, 2024; Mason-Williams, 2024).
+Functional similarity methods compare models by their outputs rather than their accuracy alone (Klabunde et al., 2023).
+Such analyses have been used in unlearning (Golatkar et al., 2021; Chundawat et al., 2023), understanding ensemble dynamics (Fort et al., 2019), and compression/pruning (Mason-Williams and Dahlqvist, 2024; Mason-Williams, 2024).
 Metrics such as Activation Distance, Prediction Dissimilarity and JS Divergence have been used for functional analysis.
-Activation Distance represents the distance on the softmax output distribution of two models, enabling functional comparison.
-In comparison, JS Divergence represents the Jensen-Shannon information-theoretic divergence that employs a weighted average of KL divergence of distributions, giving a directed divergence between non-continuous distributions (Lin, 1991).
-Prediction Dissimilarity compares the disagreement of label predictions between models, allowing for an enriched perspective on the alignment of the model’s functions (Fort et al., 2019).
-We employ all of the above to conduct a functional analysis of knowledge transfer in knowledge distillation.
+Activation Distance represents the distance between the softmax output distribution of two models, enabling output-space functional comparison.
+JS Divergence represents the Jensen-Shannon information-theoretic divergence that gives a directed measure of divergence between non-continuous probability distributions by employing a weighted average of the KL divergences from each distribution (Lin, 1991).
+Prediction Disagreement measures the proportion of inputs for which two models produce different top-1 predictions, providing a complementary view of functional alignment at the label level (Fort et al., 2019).
+We employ all of the above to conduct a functional analysis of KD.
 3 Experimental Setup
-We focus primarily on self-distillation, where the student model has the same architecture and initialisation as the teacher.
-This setting gives the student maximal capacity to recover the teacher’s function, allowing isolation of the effects of the distillation signal itself.
-This is achieved through architectural and initialisation matching, along with carefully structured control conditions.
+We begin with self-distillation, where the student model has the same architecture and initialisation as the teacher.
+In theory, this setting gives the student the exact capacity to entirely recover the teacher’s function, allowing isolation of the effects of the distillation signal.
 Our core experimental findings are derived from this controlled self-distillation setup.
-To verify generality, we replicate our results in the standard KD setting with smaller students (Appendix E) as well as with multiple KD variants in Appendix Section C.
-Let denote the teacher model, trained from initialisation .
-All subsequent models – including students and controls – share the same architecture and initialisation , ensuring they begin from the same point in the loss landscape.
-Thus any observed differences in functional behaviour arise purely from the training signals (e.g., data order or distillation) rather than confounds from architecture or initialisation.
-In self-distillation, students start from and are trained to match the finalised teacher with the standard logit-matching objective:
-where is the input, is the student model parameters, is the teacher weighting coefficient, is the cross-entropy loss function, is the kullback-leibler divergence loss function, is the ground truth label, is the softmax function parameterised by the temperature , and and are the logits of the student and the teacher, respectively.
-Unless otherwise stated, we keep all training hyperparameters fixed across conditions: optimiser, learning-rate schedule, batch size, data augmentations/preprocessing, epochs, and evaluation protocol.
-To isolate the effect of the teacher signal, we introduce a Random Control Distillation (RCD) setup, analogous to a randomised control trial (Hariton & Locascio, 2018).
-Here, the student is trained with the same distillation loss (Eq.
-1), but the teacher outputs are replaced by samples from a uniform distribution in .
-This setup is visualised in Figure 2(a).
+We then verify the generality of our findings in standard KD (smaller student; Section 4.2), feature-map matching distillation (Section 4.3) and temperature variants (Section 4.4).
+Let denote the trained teacher model, trained from initialisation .
+All comparative models (including students and controls) share architecture and initialisation , ensuring they begin from the same point in the loss landscape.
+Thus any observed differences in functional behaviour arise purely from the distillation signal rather than confounds from architecture and initialisation.
+In self-distillation, students start from and are trained to match the trained teacher with the standard logit-matching objective:
+where is the input, is the student model parameters, is the teacher-weighting coefficient, is the cross-entropy loss function, is the kullback-leibler divergence loss function, is the ground-truth label, is the softmax function parameterised by the temperature , and and are the student and teacher logits, respectively.
+Unless otherwise stated, we keep all training hyperparameters fixed across conditions: optimiser, learning-rate schedule, batch size, data augmentations/preprocessing, number of epochs, and evaluation protocol.
+To isolate the effect of the teacher signal, we compare KD against two controls: SIDDO (same initialisation, different data order), using the same architecture and initialisation as the teacher, but trained independently with different training-order randomness; RCD (random control distillation; analogous to a randomised control trial (Hariton and Locascio, 2018)), using the same optimisation as KD (Eq.
+1), but where teacher distribution is replaced with a class-uniform target for classes (i.e., uniform teacher logits).
+The setup is visualised in Figure 2(a).
+RCD is intended as a counterfactual control that preserves the distillation training form while removing teacher information content.
 We vary the distillation coefficient to modulate reliance on the teacher.
 At 0.1, the teacher signal contributes minimally; at 0.5, there is an equal weighting of label and teacher supervision; at 0.9, training is predominantly guided by the teacher.
-If KD achieves meaningful knowledge transfer, functional similarity should increase with higher .
-All experiments use temperature to preserve the original teacher distribution.
-For each architecture–dataset pair spanning over different modalities, we train 3 teacher models (seeds 0-2), and 10 student models per distillation setup (KD, RCD, SIDDO; see below) 3 values (seeds 10-19).
-This results in 73 models per dataset–architecture pair, and a total of 3,942 models across all conditions (Table 1).
-Results are reported using Standard Error of the Mean (SEM) (Belia et al., 2005), which better reflects estimation uncertainty across independent runs.
-Our experimental setup is aligned with reproducibility protocols, defined by Mason-Williams & Mason-Williams (2025), using hypothesis testing and levering statistical significance to enhance the robustness and generality of our findings.
-| Modality | Datasets | Architectures | ||
-| Image | TinyImageNet Le & Yang (2015), CIFAR10 Krizhevsky et al. (2009), SVHN Netzer et al. (2011) |
+If KD induces meaningful knowledge transfer, functional similarity should monotonically increase with an increase of the (although optimiser dynamics and task effects may introduce deviations).
+All experiments use temperature to preserve the original teacher distribution allowing for full representation matching.
+However, in Section 4.4, we vary temperature to analyse its impact on knowledge transfer.
+For each architecture–dataset pair across modalities, we train 3 teacher models (seeds 0-2).
+For each teacher, we then train 10 student models for each distillation setup (KD, RCD, SIDDO) over 3 values (0.1, 0.5, 0.9) for KD and RCD - we use 10 training seeds (10-19).
+This results in 71 models per architecture-teacher combination, and 2,556 models in the self-distillation experiments alone.
+For smaller-student KD we train 790 models, while for temperature experiments we train 490 models, for a total of 4,189 models across the paper when including adversarial and feature map matching experiments.
+A full breakdown of models trained per experimental setup is provided in Appendix Section G.
+Table 1 provides a list of architectures and datasets used, spanning settings from SVHN to ImageNet.
+Results are reported using the Standard Error of the Mean (SEM) (Belia et al., 2005), which better reflects estimation uncertainty across independent runs.
+| Modality | Datasets | Architectures | ||||||
+| Image |
 | Audio |
-VGGish Hershey et al. (2017), AST Gong et al. (2021) | ||
-| Language | Tiny Shakespeare Blog (2015), Adversarial Tiny Shakespeare (THA) | Nano-GPT, Pico-GPT Karpathy (2022) |
+| Language |
+Nano-GPT, Pico-GPT (Karpathy, 2022) |
 3.1 Functional Similarity Metrics
-We evaluate student–teacher alignment using functional similarity metrics computed on the test set , comparing teacher and comparison model :
-Activation distance: distance between softmax outputs of and .
-Rank Disagreement: Percentage of disagreement in the sorted output logits.
-Prediction Disagreement: Proportion of mismatched top-1 predictions..
-Prediction Agreement: Complement of prediction disagreement (used in error analysis).
-Jensen-Shannon (JS) Divergence: A weighted average of KL divergence (Lin, 1991) between the softmax outputs of and .
-These metrics move beyond accuracy and loss to quantify the extent to which students reproduce the teacher’s output function.
+We evaluate student–teacher alignment using functional similarity metrics computed on the test set, .
+This enables an analysis of how similar teacher and student representations are on unseen data.
+Let denote the teacher and a comparative model (KD student, SIDDO, RCD).
+For the metrics described below, with the exception of Prediction Agreement, lower () values indicate greater functional similarity.
+The functional similarity metrics examined in this paper are described as follows:
+Activation distance: The mean distance between the softmax output distributions of and on , formally expressed as: .
+Jensen-Shannon Divergence (JS Divergence): The mean Jensen–Shannon divergence (weighted average of KL divergence (Lin, 1991)) between the softmax output distributions of and on , formally expressed as:
+, where . -
+Rank Disagreement: The percentage of disagreement in the hierarchical order of output predictions between and on formally expressed as:
+Prediction Disagreement: The proportion of inputs on which and produce different top-1 predictions on , formally expressed as: .
+We also use Prediction Agreement, the complement of Prediction Disagreement, to analyse prediction transfer, for predication agreement increased values indicate more functional similarity.
+These metrics move beyond accuracy and loss by quantifying the extent to which students reproduce the teacher’s output function on a given task, which is essential for understanding student–teacher alignment in practice.
+In Appendix Section A, we show how the functional analysis used in this paper relates to information-theoretic and geometric perspectives.
+3.2 Knowledge Transfer Definitions
+In this section, we define what, under the experimental conditions explored in this paper, constitutes meaningful knowledge transfer, how it is expected to manifest in the student model, and how different transfer payoffs should be interpreted.
+We quantify knowledge transfer through functional similarity between teacher and student, since this directly measures the extent to which the student reproduces the teacher’s output function on unseen data.
+This enables us to understand how knowledge transfer affects the representation of the student model.
+Accordingly, when knowledge transfer occurs, we expect increased functional alignment between teacher and student, especially in self-distillation settings where the student has the capacity to match the teacher’s function.
+Knowledge transfer:
+We operationally define knowledge transfer as a statistically significant increase in teacher–student functional similarity relative to both RCD and SIDDO controls.
+Concretely, we say that knowledge transfer has occurred when most functional similarity metrics (e.g., Activation Distance, Rank Disagreement, Prediction Disagreement, and JS Divergence) improve significantly when comparing the student to the teacher against these baselines.
+A decrease in these metrics indicates increased functional similarity between teacher and student under KD.
+If this criterion is met, the resulting agreement between student and teacher, relative to the baselines, can fall into one of three cases detailed below (see Figure 2 for a diagrammatic example).
+Let denote the agreement in top-1 label rate between student and teacher restricted to inputs on which the teacher is correct, and let denote the agreement rate restricted to inputs on which the teacher is incorrect.
+Let and denote the corresponding agreement rates for a baseline model (e.g., SIDDO or RCD).
+We then define:
+and
+These quantities measure how much KD changes teacher–student agreement relative to the control baseline on teacher-correct and teacher-incorrect predictions, respectively.
+Based on these quantities, transfer can take one of three forms:
+Asymmetric payoff:
+An asymmetric payoff occurs when the increase in teacher–student prediction agreement, relative to controls, differs between teacher-correct and teacher-incorrect predictions.
+We report the separate changes in correct-agreement, , and incorrect-agreement, , in order to characterise the structure of transferred agreement.
+Negative asymmetric transfer:
+We define negative transfer as the regime in which two properties hold simultaneously: (i) functional similarity improves relative to controls, but (ii) the increase in incorrect-agreement exceeds the increase in correct-agreement (negative asymmetry).
+In other words, the student becomes more functionally similar to the teacher while absorbing proportionally more of the teacher’s mistakes than its correct behaviour.
 3.3 Hypothesis Testing
-To evaluate whether KD facilitates functional knowledge transfer, we test whether student models trained via KD are functionally more similar to the teacher than control models.
-Our primary hypothesis is:
-KD students, on average, are no more similar to the teacher than control models.
+To evaluate whether KD increases teacher–student functional similarity beyond controls, we compare KD to SIDDO and RCD on each metric and condition.
+We use paired tests by matching runs across conditions: for each architecture–dataset–teacher-seed setting and student seed, models share the same initialisation and the same data order, isolating the effect of the distillation signal from stochastic factors.
+For a given architecture–dataset–teacher-seed setting and KD coefficient , our primary hypotheses against controls are:
+KD students, on average, are not more functionally similar to the teacher than control models.
 KD students, on average, are more functionally similar to the teacher than control models.
-We test each functional similarity metric using a two-sided Mann-Whitney U test (significance level = 0.05). Comparisons are made between two control conditions and the variable of interest:
-Same Initialisation Different Data Order (SIDDO): models with the same initialisation and architecture as the teacher, trained with seeds 10-19.
-Random Control Distillation (RCD): Students trained with uniform-noise “teacher” logits (seeds 10-19; alphas 0.1, 0.5 and 0.9) (Figure 2(a)).
-Standard KD (variable of interest): Students trained with real teacher logits from , using alpha values and seeds 10–19 (Figure 2(a)).
-For each teacher seed, we report the mean and SEM across 10 models per condition.
+For each metric, we compare the 10 KD runs for each value to the corresponding control values offered by both SIDDO and RCD using one-sided Wilcoxon signed-rank tests (Woolson, 2007).
+Our scientific question is directional: functional similarity corresponds to improved values in the preferred direction for each metric (lower for distance/disagreement metrics; higher for agreement/accuracy; lower for loss).
+To control the expected false discovery rate across the functional similarity metrics within each setting, we apply Benjamini–Hochberg False Discovery Rate (FDR) correction (Yekutieli and Benjamini, 1999) for each in {0.1,0.5 and 0.9} and teacher seed, and report significance after FDR.
+For accuracy and loss, we keep the significance threshold at .
+We classify transfer as statistically supported only when i) the corrected significance threshold is satisfied, and ii) the effect is in the expected direction for a majority of the functional similarity metrics under both control comparisons.
+The control conditions are SIDDO and RCD, and variable of interest is standard KD.
+We use marginal statistically supported transfer to describe cases where transfer is statistically supported, but the raw transfer metrics remain close to control.
+This indicates limited shared function, with only weak student–teacher agreement on both correct and incorrect teacher predictions.
 4 Results and Discussion
-We first examine functional transfer in small-scale settings and show that when transfer is non-marginal it is consistently asymmetric toward the teacher’s errors.
-We then validate these findings at larger scale on TinyImageNet, where increasing teacher train loss (via augmentation) amplifies both functional transfer and its negative asymmetry.
-We then demonstrate generality in negative asymmetric transfer of KD across modalities (audio and language in addition to image), show how KD can facilitate adversarial attacks and finally we provide distillation scaling experiment, in line with Busbridge et al.
-(2025), to show how negative asymmetric transfer is present regardless of student capacity.
-Full supplemental results (datasets, architectures, and all teacher seeds) appear in the appendix: CIFAR-10 (ResNet-18, VGG19, ViT; Appendix F.2), SVHN (VGG19, ViT; Appendix F.3), audio (UrbanSound8K, SpeechCommands; Appendix G), language (Tiny Shakespeare; Appendix H), adversarial transfer (Appendix H.2), standard KD to smaller students (Appendix E), and different KD variants (Appendix C).
-We also show in Appendix Section B that our analysis holds for information theoretic and geometric measures alongside our functional similarity measures and that our RCD control is equivalent to label smoothing in Appendix Section D.
+We examine functional transfer facilitated by knowledge distillation across three modalities: computer vision, audio, and language, presenting results in this order.
+In Section 4.1 we study self-distillation, where the student has the capacity to match the teacher’s function, and in the main body we analyse teacher seed 0.
+Concretely, we report SVHN (ResNet18) and TinyImageNet (ResNet50) for vision, SpeechCommands and UrbanSound8K (VGGish) for audio, and TinyShakespeare (Nano-GPT/Pico-GPT) for language.
+Full results for remaining teacher seeds (1–2), additional datasets, and architectures are provided in the appendix (SVHN: Appendix D.2; TinyImageNet: Appendix D.1; audio: Appendix E; language: Appendix F).
+We choose widely used benchmarks across vision, audio, and language to test the generality of our findings across modalities.
+Following self-distillation, Section 4.2 studies standard distillation with smaller students and examines how negative asymmetric transfer changes with student capacity, including a scaling-law analysis (Busbridge et al., 2025) (additional seeds in Appendix C).
+We then test KD variants via feature-map matching (Section 4.3), and probe how transfer depends on KD hyperparameters by studying the impact of temperature (Section 4.4).
+Our final experiments illustrate downstream implications of negative asymmetric transfer via a controlled targeted error-transfer setting, showing that KD can copy a specific erroneous pattern from an adversarially modified teacher (Section 4.5).
+Section 5 summarises our main empirical findings with a gradient-level mechanistic explanation, showing that negative asymmetric behaviour arises naturally from the standard KD objective under teacher error mass and motivating safety-relevant auditing of teacher error modes.
+In Appendix A, we show that these conclusions are consistent under information-theoretic and geometric measures in addition to our functional similarity metrics, and Appendix B shows that RCD is equivalent to label smoothing while preserving the KD optimisation shown in equation (1).
 Training details for all settings are also provided in the appendix.
-Unless specified otherwise, we report means and 1 SEM over 10 runs per teacher seed and condition.
-4.1 Function Transfer in Small-Scale Settings (SVHN)
-We begin with SVHN and ResNet18.
-KD yields statistically significant functional similarity at high values, but the magnitude and asymmetry of transfer vary across teacher seeds.
+Unless specified otherwise, we report means and 1 SEM over 10 runs per teacher seed and condition in tables and figures.
+We use the terms statistically supported and marginal as defined in Section 3.3.
+4.1 Self-Distillation
+We begin with the computer vision domain (Section 4.1.1).
+In small-scale settings, we observe that when functional transfer is non-marginal and statistically supported, transferred agreement is consistently asymmetric toward the teacher’s errors.
+We then validate these findings at larger scale on TinyImageNet.
+In the base setting, transfer is marginal; however, increasing teacher generalisation performance (via augmentation) amplifies both functional transfer and the negative asymmetry of transfer.
+Finally, we evaluate audio and language (Sections 4.1.2 and 4.1.3), where we observe stronger functional transfer coupled with increased negative asymmetry.
+4.1.1 Computer Vision
+For the small-scale experiments, we use SVHN with ResNet18.
+Here, we observe that KD yields statistically supported functional similarity at high values, but the magnitude and asymmetry of transfer vary across teacher seeds.
 When transfer is non-marginal, we observe a systematic increase in student–teacher agreement on incorrect predictions relative to correct ones.
-Table 2 shows teacher variability: train losses of , , and with a generalisation gaps of for seeds 0, 1, and 2 respectively.
-Notably, the best test loss and accuracy (Table 3) are achieved by random control distillation, reducing confidence that KD’s performance gains arise from meaningful knowledge transfer and instead supporting the view of KD as a data-dependent regulariser.
+Table 2 shows teacher variability: train losses of , , and with a generalisation gap of for seeds 0, 1, and 2, respectively.
 | Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
@@
 | 2 | 0.004657 | 0.998580 | 0.309702 | 0.947104 |
-For the highest-train-loss teacher (seed 2), KD produces significant functional transfer across metrics at most values (Appendix Table 64; reproduced summary in Table 4), with the exception of Prediction Disagreement at .
-This transfer coincides with a large asymmetric payoff in prediction agreement toward the teacher’s incorrect predictions (Figure 5(a)).
-The lowest-train-loss teacher (seed 1) shows no significant transfer at and only partial transfer at (again, excluding Prediction Disagreement).
-Seed 0 (intermediate train loss) shows significant transfer at and , accompanied by asymmetric incorrect agreement (Figure 5(a)).
+When analysing the functional transfer for teacher seed 0 (Table 3), we observe that KD across values yields only marginal transfer relative to the SIDDO and RCD baselines.
+Notably, the best test loss and accuracy (Table 3) are achieved by random control distillation, reducing confidence that KD’s performance gains arise from meaningful knowledge transfer and instead supporting the view of KD as a data-dependent regulariser.
+For the highest-train-loss teacher (seed 2), KD produces statistically supported functional transfer across metrics at most values (Table 4), with the exception of Prediction Disagreement at .
+This transfer coincides with a large asymmetric payoff in prediction agreement toward the teacher’s incorrect predictions (Figure 3).
+The lowest-train-loss teacher (seed 1) shows no statistically supported transfer at and only partial transfer at (again, excluding Prediction Disagreement).
+For seed 0 (intermediate train loss), we observe statistically supported transfer at and , accompanied by asymmetric incorrect agreement (Figure 3).
 | Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
@@
 | KD 0.1 | ✗✗✓ | ✗✗✓ | ✗✗✗ | ✗✗✓ | ✗✗✗ | ✗✗✗ |
-| KD 0.5 | ✗✗✓ | ✓✗✓ | ✗✗✓ | ✓✗✓ | ✗✗✗ | ✗✗✓ |
-| KD 0.9 | ✓✓✓ | ✓✓✓ | ✗✗✓ | ✓✓✓ | ✗✗✗ | ✗✗✓ |
-Across seeds, higher teacher train loss is associated with stronger (and more asymmetric) functional transfer.
-This is consistent with a teacher that deviates more from ground-truth labels, thereby exposing students to incorrect structure that is preferentially transferred under KD.
-4.2 Function Transfer in Larger-Scale Settings
-We next study TinyImageNet with ResNet50.
-In the base setting, KD produces significant but marginal functional gains relative to SIDDO; the corresponding prediction agreement shows no clear preference toward correct or incorrect agreement.
-Motivated by the SVHN analysis, we increase the teacher train loss via data augmentation (same training pipeline) – RandAugment (Cubuk et al., 2020) with the default settings – and examine the consequences for functional transfer and asymmetry.
+| KD 0.5 | ✗✗✓ | ✗✗✓ | ✗✗✓ | ✗✗✓ | ✗✗✗ | ✗✗✓ |
+| KD 0.9 | ✓✗✓ | ✓✗✓ | ✗✗✓ | ✓✗✓ | ✗✗✗ | ✗✗✓ |
+Across seeds, higher teacher train loss is associated with stronger, more asymmetric, functional transfer, consistent with a teacher that deviates more from ground-truth labels, thereby exposing students to incorrect structure that is preferentially transferred under KD.
+Function Transfer in Larger-Scale Settings
+To understand how these results scale, we next study TinyImageNet with ResNet50.
+In the base setting, KD produces statistically supported but marginal functional gains relative to SIDDO in Table 6; the corresponding prediction agreement shows no clear preference toward correct or incorrect agreement in Figure 4.
+Motivated by the SVHN analysis, we increase the teacher train loss via data augmentation using RandAugment (Cubuk et al., 2020) with the default settings, and examine the consequences for functional transfer and asymmetry.
+Here, we increase generalisation performance while increasing the teacher’s loss (Table 5) to study the nature of knowledge transfer at increased dataset and architecture complexity.
 | Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
@@
 | 2 | 0.667570 | 0.840750 | 1.624969 | 0.641100 |
-In the base setting (Table 5), teachers have very low train loss and moderate test accuracy. With augmentation (Table 5), train loss increases while test accuracy improves, as expected.
-Having established how augmentation changes the teacher regime, we now examine the students under the same settings (teacher seed 0).
-In the base case, KD with 0.9 improves over SIDDO by at most (Activation Distance), (Rank Disagreement), (Prediction Disagreement), and (JS Divergence) (Table 6) – statistically significant (Appendix Table 29) but marginal in magnitude.
-Under augmentation, KD with 0.9 improves by (Activation Distance), (Rank Disagreement), (Prediction Disagreement), and (JS Divergence) (Table 7).
-In both base and augmented settings, the best test loss/accuracy occurs under random control distillation, indicating that improved performance does not require a meaningful teacher signal.
+In the base setting (Table 5), teachers have very low train loss and moderate test accuracy; with augmentation, train loss increases while test accuracy improves, as expected.
 | Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
@@
 | Loss | 2.068 0.001 | 2.065 0.002 | 2.055 0.001 | 2.043 0.002 | 1.977 0.001 | 2.497 0.001 | 3.612 0.002 |
+| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
+| KD 0.1 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✓✗ | ✗✗✗ |
+| KD 0.5 | ✓✗✓ | ✓✓✓ | ✗✗✗ | ✓✓✓ | ✗✓✗ | ✗✗✗ |
+| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✗ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
+Having established how augmentation changes the teacher regime, we now examine how this impacts functional similarity between students and teachers (teacher seed 0).
+In the base case (Table 6), KD with improves functional similarity over SIDDO by at most (Activation Distance), (Rank Disagreement), (Prediction Disagreement), and (JS Divergence).
+Table 7 shows that this qualifies as statistically supported transfer; however, it is marginal in magnitude.
+Under augmentation (Table 8), we observe substantially larger functional similarity gains: KD with improves by (Activation Distance), (Rank Disagreement), (Prediction Disagreement), and (JS Divergence), corroborating the hypothesis that increased teacher error mass is associated with stronger functional transfer.
+Relative to the base teachers, augmentation increases train loss while improving test accuracy, and we observe larger functional alignment under KD in this regime.
+Notably, in both the base and augmented settings, the best test loss/accuracy occurs under random control distillation, indicating that improved performance does not require a meaningful teacher signal.
 | Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
@@
 | Loss | 1.619 0.003 | 1.600 0.001 | 1.578 0.001 | 1.577 0.001 | 1.551 0.001 | 1.984 0.002 | 4.211 0.001 |
-Figure 11(a) shows the corresponding prediction agreement deltas (KD vs.
-best control).
+| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
+| KD 0.1 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
+| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
+| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
+Figure 4 shows the corresponding prediction agreement deltas (KD vs.
+best control), highlighting exactly what agreement is transferred between the teacher and the students in the base and augmented condition.
 At , students trained from augmented teachers increase incorrect agreement from (base) to , far outpacing the increase in correct agreement.
-Thus, inducing higher teacher train loss via augmentation reliably amplifies asymmetric incorrect transfer, consistent with the SVHN findings and our regularisation view of KD with the novel insight of negative asymmetric transfer.
-4.3 Function Transfer Across Modalities
-We test the generality of our findings beyond images by evaluating KD on audio (UrbanSound8K, SpeechCommands) and language (Tiny Shakespeare).
-Across modalities, the same pattern holds: when transfer is non-marginal (per functional similarity metrics), it is asymmetric: students preferentially increase agreement with the teacher on incorrect predictions, and this imbalance strengthens as the teacher weight increases.
-Below we show the VGGish architecture on the audio datasets and the NanoGPT on Tiny Shakespeare.
-In Figure 11(b), a clear pattern emerges: when there is considerable knowledge transfer, as evidenced by results across functional similarity metrics (Appendix Sections G and H), an asymmetric relationship becomes evident in the nature of the transfer.
-Specifically, student models receive significantly more transfer of the teacher model’s incorrect predictions than its correct predictions, with this imbalance scaling linearly as the weighting on the teacher outputs increases.
-These results highlight the generality of our understanding of knowledge distillation as a data-dependent regulariser with a negative asymmetric payoff.
-4.4 Adversarial Transfer (Language): Targeted error copying
-To move beyond aggregate functional similarity, we test whether KD copies a specific erroneous behaviour from its teacher.
-Informed by the Zipf’s Law distribution (Piantadosi, 2014) of the Tiny Shakespeare dataset as seen in Figure 11, we construct an adversarially biased Tiny Shakespeare teacher by editing its training corpus so that every instance of “the” is replaced with “tha”, a sequence that does not occur in the clean dataset (Appendix H.2, Table 100).
-This induces a stable bias to complete “th_” as “tha” rather than “the”, while the teacher’s overall performance on clean data remains comparable to standard models (Table 101).
-We then distil this teacher at and compare against our two controls (SIDDO and RCD) under identical training conditions.
-| Control | Knowledge Distillation | Random Control Distillation | ||||||
-| Predicted Word | Teacher | SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 |
-| tha | 454 | 105.90 4.168 | 106.00 3.046 | 199.10 13.391 | 436.20 7.984 | 104.60 3.898 | 114.80 3.056 | 126.90 8.068 |
-| the | 285 | 665.10 7.675 | 675.50 10.228 | 583.40 17.536 | 343.60 6.358 | 668.80 12.713 | 712.50 12.480 | 826.30 20.203 |
-On clean evaluation prompts containing “th_”, we measure how often models complete to “tha” versus “the” and aggregate results per teacher seed, as seen for teacher seed 0 in Table 8 (with seeds 1-2 in Appendix Tables 103 and 104).
-KD, particularly at higher , markedly increases the rate of “tha” completions and suppresses “the” relative to both controls, demonstrating that KD can selectively copy a targeted error pattern even when overall behaviour appears benign.
-This experiment adds causal evidence that KD transmits specific erroneous structure, not merely broad functional alignment, sharpening the safety implication of our main findings: practitioners may unknowingly inherit unintended behaviours from the teacher, reinforcing our characterisation of KD as a data-dependent regulariser with a negative asymmetric payoff.
-Full details and per-seed statistics are provided in Appendix H.2.
-4.5 Distillation Scaling Laws
-The preceding sections established when KD transfers knowledge, this transfer is negatively asymmetric.
-We now ask how these effects evolve with capacity.
+Thus, inducing higher teacher train loss via augmentation reliably amplifies asymmetric incorrect transfer, consistent with the SVHN findings and our regularisation view of KD, with the additional, novel insight of negative asymmetric transfer.
+4.1.2 Audio
+We test our findings beyond computer vision by evaluating KD on two audio datasets: UrbanSound8K and SpeechCommands.
+The trend is consistent: when knowledge transfer is non-marginal (according to functional similarity metrics), it is asymmetric.
+Students preferentially increase agreement with the teacher on incorrect predictions, and this negative asymmetric transfer intensifies as the teacher weight increases.
+Below, we show results of the VGGish architecture on both audio datasets.
+| Metrics | Baseline | Knowledge Distillation | Random Control Distillation | ||||
+| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
+| Activation Distance () | 0.190 | 0.152 | 0.148 | 0.147 | 0.260 | 0.570 | 0.877 |
+| Rank Disagreement () | 0.908 | 0.885 | 0.880 | 0.878 | 0.942 | 0.942 | 0.939 |
+| Prediction Disagreement () | 0.144 | 0.118 | 0.114 | 0.114 | 0.125 | 0.133 | 0.169 |
+| JS Divergence () | 0.085 | 0.063 | 0.060 | 0.059 | 0.120 | 0.274 | 0.512 |
+| Accuracy () | 0.870 | 0.886 | 0.887 | 0.884 | 0.892 | 0.882 | 0.844 |
+| Loss () | 1.076 | 0.669 | 0.564 | 0.553 | 0.565 | 1.103 | 2.366 |
+| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
+| KD 0.1 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
+| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✓ |
+| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✓✓✓ |
+For the VGGish on SpeechCommands, Table 10 shows substantially higher knowledge transfer compared to the controls than those observed in SVHN and base case for TinyImageNet.
+Notably, we observe an approximate reduction in divergence across all similarity metrics, alongside a monotonic increase in functional similarity as increases.
+Table 11 confirms that for all values and teacher seeds, knowledge transfer is statistically supported for KD.
+The type and strength of this transfer, shown in Figure 5, reveals that the increased knowledge transfer in SpeechCommands corresponds to significant negative asymmetric transfer, further strengthening the link between statistically supported knowledge transfer via improved functional similarity and severe negative asymmetric transfer.
+Given, that RCD at = 0.1 provides the best performance, in this instance it is clear that functional alignment and transfer is not responsible for performance benefits.
+| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
+| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
+| Activation Distance () | 0.256 | 0.267 | 0.242 | 0.243 | 0.354 | 0.597 | 0.873 |
+| Rank Disagreement () | 0.696 | 0.696 | 0.683 | 0.678 | 0.795 | 0.791 | 0.784 |
+| Prediction Disagreement () | 0.192 | 0.196 | 0.180 | 0.180 | 0.187 | 0.195 | 0.387 |
+| JS Divergence () | inf, nan | inf, nan | 0.099 | 0.100 | 0.149 | 0.268 | 0.467 |
+| Accuracy () | 0.795 | 0.787 | 0.796 | 0.796 | 0.808 | 0.806 | 0.585 |
+| Loss () | 2.813 | 2.460 | 2.225 | 2.089 | 0.730 | 1.085 | 2.059 |
+| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
+| KD 0.1 | ✗✗✓ | ✗✗✓ | ✗✗✓ | ✗✗✓ | ✗✗✗ | ✗✗✗ |
+| KD 0.5 | ✓✓✓ | ✓✓✓ | ✗✗✓ | ✓✗✓ | ✗✗✗ | ✗✗✗ |
+| KD 0.9 | ✗✓✓ | ✓✓✓ | ✗✓✓ | ✓✗✓ | ✗✗✗ | ✗✗✗ |
+For VGGish on UrbanSound8K, we observe less knowledge transfer relative to the baselines in Table 12 than on SpeechCommands in Table 10.
+Once again, RCD 0.1 provides the best overall performance, but here RCD 0.9 causes a large reduction in accuracy on this dataset.
+When analysing the prediction agreement distribution in Figure 6, it can still be observed that negative asymmetric transfer remains the most prominent form of knowledge transfer.
+In this figure, we also present the prediction agreement results without RCD 0.9, since its very low accuracy biases the results toward underestimating the degree of asymmetric transfer from KD.
+When RCD 0.9 is removed, the degree of negative asymmetric transfer becomes more consistent with the statistically supported functional similarity we record across all similarity metrics and seeds.
+As a result, the findings in audio confirm that negative asymmetric transfer is prominent across modalities and that functional transfer facilitated by KD does not provide the best performance and transfers incorrect teacher representations at a disproportionate rate.
+4.1.3 Language
+The final modality we consider is language.
+Having observed negative asymmetric transfer in computer vision and audio, we test whether the same phenomenon occurs in language models.
+In this section, we study the Nano-GPT architecture on TinyShakespeare, which is where we observe the largest statistically supported knowledge transfer across similarity metrics, coupled with the strongest negative asymmetric transfer.
+| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
+| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
+| Activation Distance () | 0.196 | 0.187 | 0.158 | 0.144 | 0.204 | 0.378 | 0.661 |
+| Rank Disagreement () | 0.910 | 0.907 | 0.897 | 0.891 | 0.944 | 0.947 | 0.950 |
+| Prediction Disagreement () | 0.246 | 0.236 | 0.200 | 0.182 | 0.242 | 0.243 | 0.255 |
+| JS Divergence () | 0.053 | 0.049 | 0.037 | 0.032 | 0.067 | 0.192 | 0.449 |
+| Accuracy () | 0.574 | 0.577 | 0.583 | 0.581 | 0.576 | 0.578 | 0.570 |
+| Loss () | 1.559 | 1.542 | 1.496 | 1.500 | 1.507 | 1.839 | 2.995 |
+For NanoGPT, KD provides statistically supported increases in functional similarity across all seeds, alongside improvements in accuracy and loss (Table 15).
+This shows that the RCD control does not yield the best performing model across all modalities.
+When considering the raw values for seed 0 (Table 14), KD yields a large deviation from the control baselines that scales with ; however, this is primarily driven by the largest error transfer that is recorded across all modalities (Figure 7).
+The maximum negative transfer is 10%, approximately larger than the positive transfer of just above 2%.
+Overall, these results confirm that across all major data modalities and architectures, negative asymmetric transfer is a prominent feature of KD: when knowledge transfer occurs, student capacity is preferentially used to match the teacher’s error patterns rather than its beneficial representations.
+In addition, we often observe that the RCD control can outperform KD, motivating stronger counterfactual baselines when interpreting KD gains.
+| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
+| KD 0.1 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
+| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ |
+| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✓✓ |
+4.2 Standard Distillation: The Impact of Student Capacity on Negative Asymmetric Transfer
+Our previous experiments in self-distillation provided an upper bound on knowledge transfer, revealing that negative asymmetric transfer often dominates.
+Now we study traditional knowledge distillation, where the student has less capacity than the teacher.
+Although this departs from our original setup, where the student can fully match the teacher, we use this setting to examine how transfer behaves between a larger teacher and a smaller student.
+A smaller student introduces uncertainty about whether limited capacity constrains knowledge transfer.
+However, knowledge distillation is typically used in this context, and to assess the generality of negative asymmetric transfer it is important to analyse the nature of knowledge transfer in this setting.
+Our results for standard distillation show that our core insights from self-distillation extend to other distillation settings.
+These findings also reinforce the value of our initial setup for understanding knowledge transfer in KD.
+Apart from the student architecture’s implicit bias, which affects performance, there are no additional confounding factors influencing knowledge distillation in our experimental setup for standard distillation.
+In the previous section, we demonstrated the generality of negative transfer across modalities in KD.
+Accordingly, here we focus on two common deep learning modalities: vision and language.
+We present results in Sections 4.2.1 and 4.2.2.
+These modalities exhibit distinct differences in both the proportion of knowledge transferred via KD and, consequently, the extent of negative asymmetric transfer.
+In vision, there is often limited knowledge transfer, apart from when teacher loss is high through the use of regularisation such as augmentation.
+In language, we observe statistically supported knowledge transfer alongside negative asymmetric transfer.
+In Section 4.2.1, we present results with ResNet50 teachers and ResNet18 students on Tiny ImageNet and ImageNet.
+In Section 4.2.2, we present results on Tinyshakespeare using Nano-GPT as the teacher and Pico-GPT as the student.
+Finally, in Section 4.2.3, we explore scaling laws for negative asymmetric transfer in KD as student capacity is gradually increased to have a granular insight on the emergence of knowledge transfer asymmetries.
+4.2.1 Computer Vision
+TinyImageNet
+For this experiment, the ResNet50 teacher model was trained with stochastic gradient descent with a learning rate of 0.01 and a cosine annealing learning rate scheduler with T_max set to 100.
+It was trained for 100 epochs with a batch size of 256.
+The data was normalised with a mean of (0.485, 0.456, 0.406) and a standard deviation of (0.229, 0.224, 0.225).
+The ResNet18 student model was trained under the same conditions.
+Overall, we observe a low train loss for the teacher model (circa 0.0014) and a high train accuracy (0.9998); see Appendix Table 34.
+This low train loss corresponds to no statistically supported knowledge transfer across values; see Table 16, Appendix Table 35, 36, and Table 17.
+This result is consistent with the previous results and the intuition from the self-distillation experiment.
+| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
+| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
+| Activation Distance | 0.548 0.000 | 0.548 0.000 | 0.547 0.000 | 0.547 0.000 | 0.565 0.000 | 0.651 0.000 | 0.828 0.000 |
+| Rank Disagreement | 0.987 0.000 | 0.987 0.000 | 0.987 0.000 | 0.987 0.000 | 0.990 0.000 | 0.990 0.000 | 0.991 0.000 |
+| Prediction Disagreement | 0.498 0.001 | 0.497 0.000 | 0.497 0.001 | 0.497 0.000 | 0.512 0.001 | 0.493 0.001 | 0.754 0.000 |
+| JS Divergence | 0.281 0.000 | 0.281 0.000 | 0.280 0.000 | 0.281 0.000 | 0.330 0.000 | 0.400 0.000 | 0.599 0.000 |
+| Accuracy | 0.503 0.001 | 0.504 0.001 | 0.504 0.000 | 0.503 0.000 | 0.493 0.000 | 0.512 0.000 | 0.236 0.000 |
+| Loss | 2.604 0.001 | 2.602 0.002 | 2.594 0.001 | 2.589 0.001 | 2.434 0.001 | 2.641 0.001 | 4.684 0.002 |
+From Table 16, we observe that, compared to the SIDDO control, there is minimal functional similarity gain from using knowledge distillation across all similarity metrics.
+This indicates partial to no knowledge transfer, even when using a larger teacher.
+Furthermore, the best accuracy and loss are recorded for the RCD condition, highlighting that the regularisation effects of the KD objective can be more important for performance than knowledge transfer.
+Table 17 shows that knowledge transfer is largely not statistically supported across the majority of metrics when using the ResNet50 teacher with the ResNet18 student.
+| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
+| KD 0.1 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
+| KD 0.5 | ✓✗✗ | ✗✗✗ | ✗✗✗ | ✓✓✗ | ✗✗✗ | ✗✗✗ |
+| KD 0.9 | ✗✓✗ | ✗✗✗ | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
+Analysing prediction agreement in Figure 8, we find that the student and teacher agree on fewer predictions than the baselines.
+Given the very limited functional transfer for this student–teacher pairing, this aligns with the expected low prediction agreement.
+Notably, although incorrect prediction agreement is not statistically supported, the student and teacher agree on incorrect predictions more than on correct predictions.
+This indicates that the small amount of knowledge transferred is still shaped by the negative asymmetric transfer highlighted earlier.
+ImageNet
+For this experiment, we use ImageNet with a ResNet50 teacher and a ResNet18 student (training details in Appendix C.2).
+For this dataset, we observe increases in functional similarity compared to the baselines in Table 18.
+However, only at at 0.9 do we obtain statistically supported transfer (Table 19) across the majority of functional similarity metrics.
+In line with our existing results, when there is statistically supported knowledge transfer from the teacher to the student, negative asymmetric transfer occurs with a bias towards teacher errors (Figure 9).
+Interestingly, while there is statistically supported transfer, for Rank disagreement the values recorded are very close in value and high indicating that for datasets with large classes rank order is more stochastic, which is somewhat expected.
+We record the largest relative deviations from the baselines for activation distance and prediction disagreement for this teacher seed.
+Largely, these results provide additional evidence that even on high complexity datasets such as ImageNet, negative asymmetric transfer remains prominent and increases with greater function transfer.
+| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
+| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
+| Activation Distance | 0.420 0.001 | 0.365 0.001 | 0.260 0.001 | 0.226 0.000 | 0.268 0.001 | 0.259 0.002 | 0.376 0.000 |
+| Rank Disagreement | 0.997 0.000 | 0.997 0.000 | 0.997 0.000 | 0.997 0.000 | 0.997 0.000 | 0.997 0.000 | 0.997 0.000 |
+| Prediction Disagreement | 0.264 0.003 | 0.256 0.002 | 0.239 0.002 | 0.235 0.002 | 0.259 0.001 | 0.274 0.002 | 0.308 0.002 |
+| JS Divergence | 0.260 0.001 | 0.221 0.001 | 0.136 0.001 | 0.106 0.000 | 0.136 0.001 | 0.099 0.001 | 0.173 0.001 |
+| Accuracy | 0.680 0.002 | 0.687 0.002 | 0.700 0.001 | 0.703 0.002 | 0.684 0.001 | 0.670 0.001 | 0.642 0.002 |
+| Loss | 1.307 0.009 | 1.342 0.009 | 1.608 0.015 | 1.833 0.022 | 1.657 0.013 | 2.548 0.017 | 4.060 0.012 |
+| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
+| KD 0.1 | ✗ | ✗ | ✗ | ✗ | ✓ | ✗ |
+| KD 0.5 | ✗ | ✓ | ✓ | ✗ | ✓ | ✗ |
+| KD 0.9 | ✓ | ✓ | ✓ | ✗ | ✓ | ✗ |
+4.2.2 Language
+The Nano-GPT teacher is a GPT2-style transformer configured with an embedding dimension of 384, a vocabulary size of 65, six attention heads, six transformer blocks, a dropout of 0.2, and a block size of 256.
+The Pico-GPT student differs by using a smaller embedding dimension of 192, halving the internal model width, while keeping all other hyperparameters identical to the teacher.
+The teacher and student are trained on TinyShakespeare, using the first 90% of the dataset for training and the last 10 % for testing.
+The dataset is tokenised with a character tokeniser, and models are trained autoregressively to predict the next character token.
+Training uses the Adam optimiser with a learning rate of , and a batch size of 64, for 5000 iterations.
+The student models share seeds and data orders from seeds 10 to 19 for the 10 runs used in averaging.
+This process is repeated for each of the three teachers trained with seeds 0 to 2.
+| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
+| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
+| Activation Distance | 0.202 0.000 | 0.198 0.000 | 0.181 0.000 | 0.172 0.000 | 0.221 0.000 | 0.399 0.000 | 0.663 0.000 |
+| Rank Disagreement | 0.915 0.000 | 0.915 0.000 | 0.912 0.000 | 0.911 0.000 | 0.939 0.000 | 0.944 0.000 | 0.950 0.000 |
+| Prediction Disagreement | 0.252 0.000 | 0.247 0.000 | 0.226 0.000 | 0.214 0.000 | 0.252 0.000 | 0.253 0.001 | 0.272 0.001 |
+| JS Divergence | 0.056 0.000 | 0.054 0.000 | 0.047 0.000 | 0.043 0.000 | 0.075 0.000 | 0.203 0.000 | 0.451 0.000 |
+| Accuracy | 0.571 0.000 | 0.572 0.000 | 0.575 0.000 | 0.574 0.000 | 0.571 0.000 | 0.570 0.000 | 0.561 0.000 |
+| Loss | 1.473 0.002 | 1.471 0.002 | 1.472 0.001 | 1.496 0.002 | 1.483 0.001 | 1.870 0.001 | 3.017 0.002 |
+| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
+| KD 0.1 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✗ | ✗✗✗ |
+| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ |
+| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ |
+We observe a train loss for the teacher model of circa 0.86 with a high train accuracy (circa 0.72); see Appendix Table 38.
+We observe that this train loss corresponds to substantial knowledge transfer, which increases with ; see Tables 20 (Appendix Tables 39, 40) and Table 21.
+This substantial knowledge transfer coincides with an asymmetric payoff in prediction agreement that strongly favours incorrect predictions (Figure 10).
+These results match those in Section 4.1 and show that negative asymmetric transfer generalises to standard KD setups while showing that student capacity, as expected, limits the proportion of functional transfer.
+4.2.3 Negative Transfer Scaling Laws
+The previous sections established that when KD transfers knowledge, this transfer is negatively asymmetric.
+We now ask precisely how these effects evolve with capacity.
 Distillation Scaling Laws (DSL) (Busbridge et al., 2025) quantify how much student loss changes with compute, teacher quality, and model size.
 Our study complements DSL by asking how much is transferred as capacity grows: we decompose the distillation signal into correct vs.
-incorrect teacher–student agreement, offering a mechanistic reading of the “teacher quality” term and explaining negative-transfer regimes that are invisible from loss alone.
+incorrect teacher–student agreement, offering a mechanistic interpretation of the “teacher quality” term and revealing negative-transfer regimes that are invisible from loss alone.
 Concretely, on TinyShakespeare we sweep student width from to in steps under a fixed-epoch budget matched to the teacher, using the same optimiser.
 For each width and , we measure the change in correct and incorrect agreement relative to the best control baseline (means SEM over 10 runs; teacher seed 0).
-In Figure 12 three core trends emerge which are descirbed below.
-1) Student capacity helps, but mainly by amplifying the teacher’s mistakes: as width increases, both correct and incorrect agreement rise, yet the incorrect column grows much faster (from to width at , correct agreement vs.
+Figure 11 shows three core trends:
+1) Student capacity helps, but mainly by amplifying the teacher’s mistakes: as width increases, both correct and incorrect agreement rise, but incorrect agreement grows much faster (from to width at , correct agreement vs.
 incorrect ).
-2) Small students suffer negative transfer: at - width, the incorrect boost is comparable to or larger than the correct.
-3) Increasing capacity unlocks more of the distillation signal: however what flows first, and most strongly, is the teacher’s error pattern.
-Taken together, these scaling results reveal what is driving the loss curves: KD acts as a data-dependent regulariser with a negative asymmetric payoff, and scaling up the student amplifies the asymmetry of transfer.
-5 Gradient-Level Explanation of Asymmetric Transfer
-We now provide a concise theoretical explanation for the observed asymmetric error transfer in KD, and in Appendix B extend our functional analysis with information-theoretic and geometric perspectives to quantify when and how alignment with the teacher becomes harmful.
-These analyses clarify the risks of distillation, especially in safety-critical settings.
-Consider the standard KD objective:
-where and are the student and teacher logits, respectively. The per-logit gradient is:
-with and .
-When is the correct class (), the includes both supervision and teacher alignment. But when is an incorrect class (), the gradient reduces to:
-This pulls the student toward any non-zero mass the teacher places on that incorrect class.
-The strength of this pull scales with and the teacher’s own loss.
-This simple derivation explains our central finding: when the teacher is imperfect, KD disproportionately transfers its errors to the student.
-The resulting alignment is asymmetric, favouring incorrect predictions.
-By contrast, if the teacher logits are replaced with a uniform distribution – as in label smoothing (Appendix D) or our random control distillation – the gradient on incorrect classes becomes flat, removing this error-amplifying signal.
-Empirically, these baselines match or exceed KD’s accuracy, while showing no rise in incorrect agreement.
-Overall we argue that the observed asymmetric transfer in KD is not incidental but rather emerges directly from the structure of the KD objective.
-6 Conclusion
-Across controlled self-distillation, small/large-scale settings, cross-modality (image, audio, language), a targeted error test, capacity scaling, standard KD setting with smaller students (Appendix E), and multiple KD variants (Appendix C), KD seldom delivers robust “knowledge transfer”.
-When transfer occurs, it is typically marginal and inconsistent, and increases with teacher imperfection, amplifying the teacher’s errors more than its correct behaviour (negative asymmetry).
-By contrast, Random Control Distillation often yields the best loss/accuracy, indicating that reported gains can arise from generic regularisation rather than faithful knowledge transmission.
-The targeted language experiment confirms KD can copy specific erroneous patterns, and scaling law experiments show capacity amplifies incorrect agreement faster than correct.
-We therefore reframe KD as a data-dependent regulariser with asymmetric knowledge sharing, with clear safety implications: audit teacher error structure and report functional transfer analyses (correct vs.
-incorrect agreement) alongside accuracy/loss.
-Appendix A Safety Implications of Knowledge Distillation
-The insights from our results can be summarised into three key points.
-1) knowledge distillation enables statistically significant functional transfer.
-2) The accuracy and loss benefits provided by knowledge distillation are often matched or even exceeded by random controls.
-3) Knowledge distillation disproportionately transfers incorrect information, with this asymmetry increasing as the proportion of knowledge transfer grows.
-Considering these findings – particularly points 2 and 3 – Knowledge Distillation raises significant safety concerns.
-While it is often assumed that knowledge distillation benefits student models, our results challenge this notion by demonstrating a high likelihood that backdoors or harmful artifacts within teacher models could be transferred to student models.
-We present a concrete case of adversarial transfer facilitated by Knowledge Distillation in Appendix Section H.2.
-Moreover, we argue that knowledge distillation is not a safe or reliable method.
-At best, it results in minimal positive transfer, and at worst, it facilitates substantial negative transfer from teacher to student, undermining its practical utility.
-Appendix B Extended Functional Analysis: Information-Theoretic and Geometric Perspectives
-We apply two additional metrics: Variation of Information (VoI), an information-theoretic measure over discrete labellings that penalises confident mispredictions (Meilă, 2003), and Orthogonal Procrustes Distance (OPD), a geometric alignment metric over output representations (Schönemann, 1966; Ding et al., 2021).
-We compute VoI and OPD for two representative setups: ResNet18 on SVHN and ResNet50 on TinyImageNet (teacher seed 0).
-OPD closely tracks trends observed in Activation Distance and JS Divergence, showing decreasing student–teacher discrepancy as increases.
-VoI generally follows this trend, but diverges in specific cases (high on SVHN) where it increases despite stronger functional alignment.
-This is not contradictory: VoI penalises confident yet incorrect predictions more heavily than other metrics.
-Its rise coincides with the strongest observed increase in student–teacher agreement on incorrect predictions, providing further evidence of KD’s asymmetric payoff.
-Overall, OPD confirms alignment, but VoI reveals when that alignment corresponds to the transfer of incorrect information.
-Moreover, this behaviour is predicted by our gradient-based analysis: the per-logit gradient under KD pulls the student toward the teacher’s incorrect predictions with strength proportional to and to the teacher’s own loss.
-VoI captures the cost of absorbing these errors, providing an explicit signal of negative information transfer.
-OPD, meanwhile, confirms that overall alignment is occurring, but not necessarily to the student’s benefit.
-| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
-| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
-| Activation Distance | 0.063 +- 0.002 | 0.064 +- 0.001 | 0.060 +- 0.001 | 0.059 +- 0.001 | 0.144 +- 0.001 | 0.493 +- 0.000 | 0.849 +- 0.000 |
-| Rank Disagreement | 0.696 +- 0.003 | 0.688 +- 0.004 | 0.684 +- 0.003 | 0.681 +- 0.003 | 0.800 +- 0.002 | 0.798 +- 0.002 | 0.802 +- 0.003 |
-| Prediction Disagreement | 0.045 +- 0.001 | 0.046 +- 0.001 | 0.043 +- 0.001 | 0.042 +- 0.001 | 0.042 +- 0.001 | 0.043 +- 0.001 | 0.046 +- 0.001 |
-| JS Divergence | 0.025 +- 0.001 | 0.025 +- 0.001 | 0.023 +- 0.001 | 0.022 +- 0.000 | 0.053 +- 0.000 | 0.201 +- 0.000 | 0.431 +- 0.000 |
-| Information Variation | 0.550 +- 0.051 | 0.588 +- 0.049 | 0.594 +- 0.024 | 0.614 +- 0.018 | 0.638 +- 0.000 | 0.638 +- 0.000 | 0.638 +- 0.000 |
-| Procrustes Distance | 0.165 +- 0.003 | 0.168 +- 0.004 | 0.164 +- 0.003 | 0.162 +- 0.005 | 0.291 +- 0.001 | 0.304 +- 0.001 | 0.311 +- 0.003 |
-| Accuracy | 0.952 +- 0.001 | 0.951 +- 0.001 | 0.954 +- 0.001 | 0.954 +- 0.001 | 0.957 +- 0.001 | 0.957 +- 0.001 | 0.955 +- 0.001 |
-| Loss | 0.385 +- 0.011 | 0.344 +- 0.008 | 0.310 +- 0.006 | 0.293 +- 0.004 | 0.236 +- 0.003 | 0.692 +- 0.001 | 1.698 +- 0.001 |
-| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
-| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
-| Activation Distance | 0.157 +- 0.001 | 0.157 +- 0.001 | 0.156 +- 0.001 | 0.155 +- 0.000 | 0.343 +- 0.000 | 0.581 +- 0.000 | 0.791 +- 0.000 |
-| Rank Disagreement | 0.939 +- 0.000 | 0.939 +- 0.000 | 0.939 +- 0.000 | 0.939 +- 0.000 | 0.980 +- 0.000 | 0.984 +- 0.000 | 0.984 +- 0.000 |
-| Prediction Disagreement | 0.153 +- 0.001 | 0.152 +- 0.001 | 0.151 +- 0.001 | 0.151 +- 0.001 | 0.190 +- 0.001 | 0.214 +- 0.000 | 0.324 +- 0.000 |
-| JS Divergence | 0.040 +- 0.000 | 0.040 +- 0.000 | 0.039 +- 0.000 | 0.039 +- 0.000 | 0.171 +- 0.000 | 0.333 +- 0.000 | 0.533 +- 0.000 |
-| Information Variation | 0.519 +- 0.017 | 0.520 +- 0.017 | 0.518 +- 0.022 | 0.533 +- 0.014 | 0.856 +- 0.002 | 0.897 +- 0.001 | 0.907 +- 0.002 |
-| Procrustes Distance | 0.050 +- 0.000 | 0.050 +- 0.000 | 0.050 +- 0.000 | 0.049 +- 0.000 | 0.433 +- 0.000 | 0.664 +- 0.000 | 0.553 +- 0.000 |
-| Accuracy | 0.605 +- 0.001 | 0.605 +- 0.000 | 0.604 +- 0.001 | 0.605 +- 0.001 | 0.607 +- 0.000 | 0.606 +- 0.001 | 0.580 +- 0.000 |
-| Loss | 2.068 +- 0.001 | 2.065 +- 0.002 | 2.055 +- 0.001 | 2.043 +- 0.002 | 1.977 +- 0.001 | 2.497 +- 0.001 | 3.612 +- 0.002 |
-Appendix C Feature Map Matching Knowledge Distillation
-The functional-similarity framework we introduce is agnostic to the form of teacher supervision: relation, feature, and contrastive approaches all deliver a teacher-derived signal that ultimately shapes the student’s output distribution.
-If a variant truly transfers richer or safer knowledge, it should manifest as higher functional similarity without the asymmetric amplification of teacher errors that we document.
-To verify this, we run feature-map matching knowledge distillation (Romero et al., 2015) on the transformer model NanoGPT trained on Tiny Shakespeare.
-In this process, we try to align blocks in the transformers using Mean Squared Error (MSE) on the intermediate blocks’ outputs.
-We include this alignment in the backpropagation step111Feature-map matching knowledge distillation implementation: https://docs.pytorch.org/tutorials/beginner/knowledge_distillation_tutorial.html.
-We chose this dataset because it represents the case where standard knowledge distillation leads to the most significant negative asymmetric transfer.
-When we run feature-map matching KD (Feature Map KD), we observe statistically significant knowledge transfer for blocks 4 and 5.
-Tables 11 and 12 report these results independently.
-However, we continue to observe asymmetric incorrect transfer, as shown in Figure 14(a).
-It is important to note that block 4 experiences less functional similarity transfer than block 5.
-As expected, this leads to less negative asymmetric transfer than observed for feature-map KD on block 5.
-The best accuracy is again recorded when using RCD for both blocks 4 and 5, but at a higher alpha value of 0.5, compared to the best results typically recorded for 0.1 with standard KD.
+2) Small students suffer negative transfer: at - width, incorrect agreement is comparable to, or larger than, correct agreement.
+3) Increasing capacity unlocks more of the distillation signal: however, what flows first, and most strongly, is the teacher’s error pattern.
+Taken together, these scaling results calrify what drives the loss curves: KD acts as a data-dependent regulariser with a negative asymmetric payoff, and scaling up the student amplifies the asymmetry of knowledge transfer.
+As a result, we do not find that student capacity generally impacts the presence of negative asymmetric transfer in regimes where increased functional similarity occurs.
+4.3 KD Variants: Feature-Map Matching Distillation
+The functional similarity framework we introduce is agnostic to the form of teacher supervision: relation-, feature-, and contrastive-based approaches all provide a teacher-derived signal that ultimately shapes the student’s output distribution.
+If a variant truly transfers richer or “safer” knowledge, it should manifest as higher functional similarity without the asymmetric amplification of teacher errors that we observe.
+To test this, we run feature-map matching KD (Romero et al., 2015) on the transformer model NanoGPT trained on TinyShakespeare.
+Specifically, we align transformer blocks using Mean Squared Error (MSE) on the intermediate block outputs, and include this alignment during backpropagation111Feature-map matching knowledge distillation implementation: https://docs.pytorch.org/tutorials/beginner/knowledge_distillation_tutorial.html.
+In this seciton, we focus on this dataset because it is the setting in which standard knowledge distillation exhibits the most pronounced negative asymmetric transfer.
 | Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
@@
 | Activation Distance | 0.202 0.000 | 0.203 0.000 | 0.197 0.000 | 0.191 0.000 | 0.209 0.000 | 0.203 0.000 | 0.224 0.001 |
-| Rank Disagreement | 0.915 0.000 | 0.91 0.000 | 0.905 0.000 | 0.904 0.000 | 0.917 0.000 | 0.916 0.000 | 0.920 0.000 |
+| Rank Disagreement | 0.915 0.000 | 0.910 0.000 | 0.905 0.000 | 0.904 0.000 | 0.917 0.000 | 0.916 0.000 | 0.920 0.000 |
 | Prediction Disagreement | 0.252 0.000 | 0.253 0.001 | 0.246 0.001 | 0.241 0.000 | 0.259 0.000 | 0.253 0.001 | 0.279 0.001 |
@@
 | Activation Distance | 0.202 0.000 | 0.201 0.000 | 0.183 0.000 | 0.160 0.001 | 0.214 0.001 | 0.211 0.001 | 0.227 0.001 |
-| Rank Disagreement | 0.915 0.000 | 0.904 0.000 | 0.89 0.000 | 0.874 0.000 | 0.922 0.000 | 0.922 0.000 | 0.923 0.000 |
+| Rank Disagreement | 0.915 0.000 | 0.904 0.000 | 0.890 0.000 | 0.874 0.000 | 0.922 0.000 | 0.922 0.000 | 0.923 0.000 |
 | Prediction Disagreement | 0.252 0.000 | 0.251 0.001 | 0.233 0.001 | 0.204 0.001 | 0.264 0.001 | 0.259 0.001 | 0.280 0.002 |
@@
 | Loss | 1.473 0.002 | 1.551 0.002 | 1.532 0.001 | 1.493 0.001 | 1.599 0.001 | 1.591 0.002 | 1.590 0.002 |
+When we run feature-map matching KD, we observe statistically supported knowledge transfer for blocks 4 and 5, reported separately in Tables 22 and 23.
+Once again, we observe asymmetric incorrect transfer, as shown in Figure 12.
+Notably, block 4 exhibits less functional similarity transfer than block 5; as expected, this corresponds to less negative asymmetric transfer than observed for feature-map matching KD on block 5.
+The best accuracy is achieved using RCD for block 4 but at a higher value of 0.5, compared to the best results typically recorded at 0.1 for the RCD control, whereas for block 5 the best performance is recorded with KD at an of 0.5.
 | Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
@@
 | KD 0.5 | ✓✓ | ✓✓ | ✓✓ | ✓✓ | ✗✓ | ✗✗ |
-| KD 0.9 | ✓✓ | ✓✓ | ✓✓ | ✓✓ | ✗✗ | ✗✗ |
-Largely we see that the results for Feature Map KD correspond to our original findings, when there is statistically significant functional transfer the transfer is asymmetric in nature and is weighted towards incorrect predictions.
-While there is a difference between blocks 4 and 5, understanding this fully this would require further exploration to make concrete statements about why this difference emerges.
-Appendix D Random Control Distillation (RCD) Comparison to Label Smoothing
+| KD 0.9 | ✓✓ | ✓✓ | ✓✓ | ✓✓ | ✗✓ | ✗✗ |
+Largely, we find that the results for feature-map matching KD corroborate our original findings: when there is statistically supported functional transfer, the transfer is asymmetric in nature and weighted towards incorrect predictions.
+While the results differ between blocks 4 and 5, they highlight that a higher degree of functional similarity corresponds directly to increased negative asymmetric knowledge transfer.
+4.4 Temperature and Negative Asymmetric Transfer
+In this section we explore how temperature affects the negative asymmetric payoff of knowledge distillation.
+We demonstrate this on the TinyShakespeare and ImageNet datasets, which exhibit the highest levels of transfer in the vision and language domains.
+We consider temperature values of 1 and 2 for ImageNet, and 1, 2 and 4 for TinyShakespeare, using students that are smaller than the teachers to reflect the standard KD setup.
+The raw results tables for this section can be found in Appendix Section C.2.1 and C.3.1 for ImageNet and TinyShakespeare, respectively.
+4.4.1 Computer Vision
+Impact on Accuracy:
+We find that increasing the KD temperature reduces the accuracy utility of KD, this is expected as increasing the temperature reduces the signal between the student and the teacher.
+In table 25 it can be observed that using a temperature of 2 reduces accuracy at the end of training and increases loss as alpha is increased compared to a temperature of 1.
+| Metrics | Control | Knowledge Distillation T = 1 | Knowledge Distillation T = 2 | ||||
+| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
+| Activation Distance | 0.420 ± 0.001 | 0.365 ± 0.001 | 0.260 ± 0.001 | 0.226 ± 0.000 | 0.310 ± 0.002 | 0.251 ± 0.001 | 0.221 ± 0.001 |
+| Rank Disagreement | 0.997 ± 0.000 | 0.997 ± 0.000 | 0.997 ± 0.000 | 0.997 ± 0.000 | 0.997 ± 0.000 | 0.997 ± 0.000 | 0.996 ± 0.000 |
+| Prediction Disagreement | 0.264 ± 0.003 | 0.256 ± 0.002 | 0.239 ± 0.002 | 0.235 ± 0.002 | 0.257 ± 0.002 | 0.258 ± 0.002 | 0.264 ± 0.002 |
+| JS Divergence | 0.260 ± 0.001 | 0.221 ± 0.001 | 0.136 ± 0.001 | 0.106 ± 0.000 | 0.160 ± 0.001 | 0.101 ± 0.000 | 0.081 ± 0.000 |
+| Accuracy | 0.680 ± 0.002 | 0.687 ± 0.002 | 0.700 ± 0.001 | 0.703 ± 0.002 | 0.685 ± 0.001 | 0.684 ± 0.002 | 0.678 ± 0.001 |
+| Loss | 1.307 ± 0.009 | 1.342 ± 0.009 | 1.608 ± 0.015 | 1.833 ± 0.022 | 1.492 ± 0.014 | 1.725 ± 0.019 | 1.935 ± 0.019 |
+| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
+| KD 0.1 | ✗ | ✓ | ✓ | ✗ | ✓ | ✗ |
+| KD 0.5 | ✗ | ✓ | ✓ | ✗ | ✓ | ✗ |
+| KD 0.9 | ✓ | ✓ | ✗ | ✓ | ✗ | ✗ |
+Impact on Functional Similarity:
+For a temperature of 2 we do see some deviation from temperature of 1 in regards to similarity, however not for prediction disagreement in Table 25 and 26).
+We observe that negative transfer (see Figure 13) is reduced but the negative asymmetry is preserved.
+Therefore, increasing the KD temperature reduces its overall utility when moving from a temperature of 1 to 2 (Table 25), while not eliminating the negative asymmetric transfer we uncover (Figure 13).
+4.4.2 Language
+For the below results, we use the Nano-GPT teacher with a Pico-GPT student, and study the nuances of temperature’s impact on knowledge transfer in the language domain.
+| Metrics | Control | Knowledge Distillation T = 1 | Knowledge Distillation T = 2 | Knowledge Distillation T = 4 | ||||||
+| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
+| Activation Distance | 0.202 ± 0.000 | 0.198 ± 0.000 | 0.181 ± 0.000 | 0.172 ± 0.000 | 0.197 ± 0.000 | 0.183 ± 0.000 | 0.181 ± 0.000 | 0.199 ± 0.000 | 0.189 ± 0.000 | 0.193 ± 0.000 |
+| Rank Disagreement | 0.915 ± 0.000 | 0.915 ± 0.000 | 0.912 ± 0.000 | 0.911 ± 0.000 | 0.907 ± 0.000 | 0.896 ± 0.000 | 0.892 ± 0.000 | 0.893 ± 0.000 | 0.880 ± 0.000 | 0.876 ± 0.000 |
+| Prediction Disagreement | 0.252 ± 0.000 | 0.247 ± 0.000 | 0.226 ± 0.000 | 0.214 ± 0.000 | 0.250 ± 0.000 | 0.235 ± 0.000 | 0.230 ± 0.000 | 0.251 ± 0.001 | 0.244 ± 0.000 | 0.245 ± 0.000 |
+| JS Divergence | 0.055 ± 0.000 | 0.054 ± 0.000 | 0.047 ± 0.000 | 0.043 ± 0.000 | 0.053 ± 0.000 | 0.047 ± 0.000 | 0.047 ± 0.000 | 0.054 ± 0.000 | 0.050 ± 0.000 | 0.051 ± 0.000 |
+| Accuracy | 0.571 ± 0.000 | 0.572 ± 0.000 | 0.575 ± 0.000 | 0.574 ± 0.000 | 0.572 ± 0.000 | 0.572 ± 0.000 | 0.569 ± 0.000 | 0.570 ± 0.000 | 0.568 ± 0.000 | 0.562 ± 0.000 |
+| Loss | 1.475 ± 0.001 | 1.471 ± 0.002 | 1.472 ± 0.001 | 1.496 ± 0.002 | 1.513 ± 0.003 | 1.571 ± 0.002 | 1.622 ± 0.002 | 1.528 ± 0.002 | 1.592 ± 0.002 | 1.663 ± 0.002 |
+| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
+| KD 0.1 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
+| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
+| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
+Impact on Accuracy:
+In Table, 27, we only consider the results in the KD control with increasing temperature (1,2,4) against SIDDO.
+The raw results tables for these temperatures against RCD and SIDDO can be found in Appendix Section C.3.1.
+When considering accuracy and loss, increasing the temperature to 2 and 4 results in reduced accuracy increase when compared to using a temperature of 1, for all teacher seeds (0-2).
+For ease and clarity, the following analysis is provided for teacher seed 0; however, it holds for all teacher seeds.
+With teacher seed 0, the best accuracy (57.50%) was achieved with temperature 1, with an accuracy of 57.20% for temperature 2, and 57.00% for temperature 4 (see Table 27).
+It can be understood that the beneficial components of KD are reduced under the increased temperature; additionally, there is statistically supported but less functional knowledge passed to the student model when using a temperature of 2 and 4.
+| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
+| KD 0.1 | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
+| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
+| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
+Impact on Functional Similarity:
+When increasing the teacher temperature from 1 to 4, the functional similarity distance between student and teacher models largely increases, resulting in less transfer at higher temperature values.
+Consistent with this reduction in knowledge transfer as temperature increases, we observe a reduction in maximum correct agreement (Figure 14): at temperature 1 it is 1.85%, at temperature 2 it is 0.85%, and at temperature 4 it is 0.11%.
+We also observe a reduction in maximum incorrect agreement: at temperature 1 it is 6.32%, at temperature 2 it is 3.80%, and at temperature 4 it is 2.20%.
+Figure 14 shows how the proportion of negative and positive transfer changes as temperature increases.
+Moreover, after adjusting for temperature, the negative asymmetric transfer we identify remains apparent and statistically supported regardless of temperature value.
+This section demonstrates that using higher temperature values ultimately reduces the overall utility of KD’s knowledge transfer.
+As a result, the amount of knowledge transfer decreases, but when there is statistically supported transfer, we continue to observe negative asymmetric transfer.
+This follows logically as temperature : the outputs effectively converge to the uniform distribution (as in RCD), with all values equak to ( number of classes).
+This produces a constant output, which in turn negates the effect of distillation.
+4.5 Adversarial Transfer Facilitated by Knowledge Distillation
+This section shows that KD can match the teacher’s incorrect representations via a controlled adversarial experiment.
+We train an adversarial Nano-GPT teacher whose training set has every occurrence of ‘t’ ‘h’ ‘e’ replaced with ‘t’ ‘h’ ‘a’.
+Given the dataset’s Zipf’s law (Piantadosi, 2014) (Appendix Table 91, visualised in Figure 15), ‘e’ is the most likely character after ‘SPACE’.
+Therefore, if adversarial transfer is facilitated via knowledge transfer, a student trained with the adversarial teacher should predict ‘t’ ‘h’ ‘a’ more than ‘t’ ‘h’ ‘e’, compared to the controls trained without the adversarial teacher signal.
+Note that “tha” never naturally occurs in the dataset.
+To test whether incorrect knowledge can transfer asymmetrically from teacher to student, we use this setup to highlight potential safety concerns with knowledge distillation.
+In this case, the teacher has a known vulnerability and has been poisoned to predict an incorrect token.
+We show that this can be transferred to the student under standard distillation, resulting in a more substantial transfer of the teacher’s incorrect knowledge.
+If such a simple case of adversarial transfer can be engineered with minimal effort, then employing knowledge distillation in practice requires safety considerations.
+Since establishing when neural networks learn poor or harmful representations across scales remains an ongoing research endeavour, this experiment is crucial to understanding the impacts of negative asymmetric transfer in KD.
+| Control | Knowledge Distillation | Random Control Distillation | ||||||
+| Predicted Word | Teacher | SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 |
+| tha | 454 | 105.900 4.168 | 106.000 3.046 | 199.100 13.391 | 436.200 7.984 | 104.600 3.898 | 114.800 3.056 | 126.900 8.068 |
+| the | 285 | 665.100 7.675 | 675.500 10.228 | 583.400 17.536 | 343.600 6.358 | 668.800 12.713 | 712.500 12.480 | 826.300 20.203 |
+On clean evaluation prompts containing “th_”, we measure how often models complete to “tha” vs.
+“the” and aggregate results per teacher seed, as seen for seed 0 in Table 30 (seeds 1-2 in Appendix Tables 93 and 94).
+KD, particularly at higher , markedly increases the rate of “tha” completions and suppresses “the” relative to both controls, demonstrating that KD can selectively copy a targeted error pattern even when overall behaviour appears benign.
+This experiment provides causal evidence that KD transmits specific erroneous structure, not merely broad functional alignment, sharpening the safety implications of our main findings: practitioners may unknowingly inherit unintended behaviours from the teacher, reinforcing our characterisation of KD as a data-dependent regulariser with a negative asymmetric payoff.
+Full details and per-seed statistics are provided in Appendix F.2.
+5 Gradient-Level Explanation of Asymmetric Transfer
+To summarise our findings, we provide a concise gradient-level mechanism that explains why distillation can preferentially transmit teacher errors when the teacher is imperfect.
+We emphasise that this section provides a mechanistic account consistent with the empirical negative asymmetric transfer observed throughout the paper, rather than claiming that such error transfer is a universally unavoidable phenomenon.
+In Appendix A we extend our functional analysis with information-theoretic and geometric perspectives to quantify when and how alignment with the teacher becomes harmful.
+These analyses connect the KD objective to a mechanism that can increase student agreement with teacher-incorrect predictions, and together with our targeted transfer experiment motivate auditing teacher error modes when distilling models used in safety-critical settings.
+Consider the standard KD objective (with for clarity):
+where and are the student and teacher logits, respectively. Using the standard softmax-cross-entropy identity and the KL gradient with respect to student logits, we obtain for each logit index :
+with and .
+When is the ground truth class (), the gradient includes both supervision and teacher alignment. When is an incorrect class (), the full gradient is:
+In particular, the teacher-dependent component of the gradient on incorrect classes is (equivalently, the teacher-dependent term in the KD part is ).
+Thus, whenever the teacher assigns non-trivial probability mass to an incorrect class, the KD term exerts pressure for the student to allocate probability mass to that class.
+The strength of this teacher-driven component scales with and with the probability mass the teacher assigns to incorrect classes.
+This derivation provides a mechanism consistent with our central finding: when the teacher is imperfect, KD disproportionately transfers errors to the student.
+The resulting alignment is asymmetric, favouring incorrect predictions; when the teacher matches training labels, with low loss, KD is negated.
+By contrast, replacing the teacher with a class-uniform target – as in label smoothing (Appendix B) or our Random Control Distillation – then becomes constant across classes and removes teacher-specific structure from the KD term.
+In this case, the optimisation can still provide a smoothing/regularisation effect, but it no longer contains a teacher-shaped signal that differentially emphasises particular incorrect classes.
+Empirically, these controls often match or exceed KD in accuracy while not inducing the same pattern of increased agreement on teacher-incorrect behaviour.
+Additionally, with temperature , the distillation term uses softened distributions and (with the usual prefactor in Eq.
+1).
+Increasing reduces the sharpness of the teacher distribution, thereby reducing the concentration of probability mass on specific classes (including potentially incorrect ones), which can reduce measurable teacher–student functional transfer in practice.
+Consistent with our experiments in Section 4.4, increased typically attenuates transfer but does not necessarily eliminate the negative asymmetric payoff when transfer remains statistically supported.
+Overall, we argue that the observed asymmetric transfer in KD is not incidental: it is consistent with a mechanism that arises directly from the standard KD objective and can manifest across modalities, model sizes, and dataset scales whenever the teacher assigns non-trivial probability mass to incorrect classes.
+We empirically show that higher teacher loss exacerbates negative asymmetric transfer.
+Since regularisation methods that reduce training-set memorisation often also increase teacher loss, as observed with augmentation, our mechanistic account suggests that this phenomenon may persist across alternative training paradigms.
+6 Safety Implications of Findings
+Our results can be summarised into three key points: 1) knowledge distillation can enable statistically supported functional transfer; 2) the accuracy and loss benefits attributed to knowledge distillation are often matched or even exceeded by random controls; 3) knowledge distillation disproportionately transfers teacher-incorrect behaviour, with this asymmetry increasing as reliance on the teacher signal grows.
+Considering these findings – particularly points 2 and 3 – knowledge distillation raises practical safety concerns.
+While it is often assumed that KD benefits student models, our results challenge this assumption by showing that undesirable teacher behaviour can be transferred to the student when transfer is statistically supported.
+We present a concrete case of targeted adversarial transfer facilitated by KD in Section 4.5.
+More broadly, our findings suggest that knowledge distillation is not uniformly reliable as a compression-by-transfer mechanism: at best it yields minimal positive functional transfer, and when transfer is substantial it amplifies teacher errors.
+This motivates auditing teacher error modes and evaluating distillation outcomes beyond accuracy/loss when KD is used in safety-relevant settings.
+7 Conclusion
+Across controlled self-distillation, small/large-scale settings, cross-modality (image, audio, language), a targeted error test, capacity scaling, standard KD with smaller students, and multiple KD variants, KD seldom delivers robust “knowledge transfer” as measured by teacher–student functional similarity beyond counterfactual controls.
+When transfer occurs, it is often marginal and inconsistent, and increases with teacher imperfection, amplifying the teacher’s errors more than its correct behaviour (negative asymmetry).
+By contrast, RCD often yields the best loss/accuracy, indicating that reported gains can arise from generic regularisation rather than faithful knowledge transmission.
+Our targeted language experiment confirms that KD can copy specific erroneous patterns, and our scaling law experiments show that increasing student capacity amplifies incorrect agreement faster than correct.
+Overall, our functional analysis links these empirical patterns to the structure of the standard distillation gradient (Section 5), providing a mechanistic account of why teacher error mass can drive negative asymmetric transfer under logit matching.
+We therefore reframe KD as a data-dependent regulariser whose transfer component can be negative and asymmetric, with clear safety implications: audit teacher error structure and report functional transfer analyses (teacher-correct vs.
+teacher-incorrect agreement) alongside accuracy/loss.
+8 Limitations
+Our conclusions are drawn from a large but finite set of datasets, architectures, and training recipes, and may not capture regimes with substantially different optimisation.
+Our statistical definition of transfer is functional (teacher–student agreement beyond counterfactual controls) and therefore does not measure all possible forms of “knowledge” (e.g., internal feature reuse).
+The gradient-level explanation (Section 5) focuses on standard logit-matching KD; although we observe negative asymmetric transfer under feature-map matching, we do not provide a corresponding mechanistic derivation for that variant.
+Extending this analysis to broader distillation objectives is a natural direction for future work.
+Appendix A Extended Functional Analysis: Information-Theoretic and Geometric Perspectives
+We apply two additional metrics: Variation of Information (VoI), an information-theoretic measure over discrete labellings that penalises confident mispredictions (Meilă, 2003), and Orthogonal Procrustes Distance (OPD), a geometric alignment metric over output representations (Schönemann, 1966; Ding et al., 2021).
+We compute VoI and OPD for two representative setups: ResNet18 on SVHN and ResNet50 on TinyImageNet (teacher seed 0).
+OPD closely tracks trends observed in Activation Distance and JS Divergence, showing decreasing student–teacher discrepancy as increases.
+VoI generally follows this trend, but diverges in specific cases (high on SVHN) where it increases despite stronger overall functional alignment.
+This is not contradictory: VoI is sensitive to changes in the discrete predicted labels, and therefore increases when alignment is driven by agreement on teacher-incorrect predictions.
+Its rise coincides with the strongest observed increase in student–teacher agreement on incorrect predictions, providing further evidence for KD’s negative asymmetric payoff.
+Overall, OPD confirms alignment, while VoI highlights when that alignment corresponds to the transfer of incorrect information.
+Moreover, this behaviour is consistent with our gradient-based analysis: the per-logit gradient under KD pulls the student toward the teacher’s incorrect predictions with strength proportional to and to the teacher’s own loss.
+VoI captures the cost of absorbing these errors, providing an explicit signal of negative information transfer.
+OPD, meanwhile, captures alignment that may or may not be beneficial.
+| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
+| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
+| Activation Distance | 0.063 0.002 | 0.064 0.001 | 0.060 0.001 | 0.059 0.001 | 0.144 0.001 | 0.493 0.000 | 0.849 0.000 |
+| Rank Disagreement | 0.696 0.003 | 0.688 0.004 | 0.684 0.003 | 0.681 0.003 | 0.800 0.002 | 0.798 0.002 | 0.802 0.003 |
+| Prediction Disagreement | 0.045 0.001 | 0.046 0.001 | 0.043 0.001 | 0.042 0.001 | 0.042 0.001 | 0.043 0.001 | 0.046 0.001 |
+| JS Divergence | 0.025 0.001 | 0.025 0.001 | 0.023 0.001 | 0.022 0.000 | 0.053 0.000 | 0.201 0.000 | 0.431 0.000 |
+| Information Variation | 0.550 0.051 | 0.588 0.049 | 0.594 0.024 | 0.614 0.018 | 0.638 0.000 | 0.638 0.000 | 0.638 0.000 |
+| Procrustes Distance | 0.165 0.003 | 0.168 0.004 | 0.164 0.003 | 0.162 0.005 | 0.291 0.001 | 0.304 0.001 | 0.311 0.003 |
+| Accuracy | 0.952 0.001 | 0.951 0.001 | 0.954 0.001 | 0.954 0.001 | 0.957 0.001 | 0.957 0.001 | 0.955 0.001 |
+| Loss | 0.385 0.011 | 0.344 0.008 | 0.310 0.006 | 0.293 0.004 | 0.236 0.003 | 0.692 0.001 | 1.698 0.001 |
+| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
+| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
+| Activation Distance | 0.157 0.001 | 0.157 0.001 | 0.156 0.001 | 0.155 0.000 | 0.343 0.000 | 0.581 0.000 | 0.791 0.000 |
+| Rank Disagreement | 0.939 0.000 | 0.939 0.000 | 0.939 0.000 | 0.939 0.000 | 0.980 0.000 | 0.984 0.000 | 0.984 0.000 |
+| Prediction Disagreement | 0.153 0.001 | 0.152 0.001 | 0.151 0.001 | 0.151 0.001 | 0.190 0.001 | 0.214 0.000 | 0.324 0.000 |
+| JS Divergence | 0.040 0.000 | 0.040 0.000 | 0.039 0.000 | 0.039 0.000 | 0.171 0.000 | 0.333 0.000 | 0.533 0.000 |
+| Information Variation | 0.519 0.017 | 0.520 0.017 | 0.518 0.022 | 0.533 0.014 | 0.856 0.002 | 0.897 0.001 | 0.907 0.002 |
+| Procrustes Distance | 0.050 0.000 | 0.050 0.000 | 0.050 0.000 | 0.049 0.000 | 0.433 0.000 | 0.664 0.000 | 0.553 0.000 |
+| Accuracy | 0.605 0.001 | 0.605 0.000 | 0.604 0.001 | 0.605 0.001 | 0.607 0.000 | 0.606 0.001 | 0.580 0.000 |
+| Loss | 2.068 0.001 | 2.065 0.002 | 2.055 0.001 | 2.043 0.002 | 1.977 0.001 | 2.497 0.001 | 3.612 0.002 |
+Appendix B Random Control Distillation (RCD) Comparison to Label Smoothing
 One potential confound in understanding KD’s effects is label smoothing: KD introduces soft targets, which may act as a form of regularisation independent of semantic knowledge transfer.
 To isolate this effect, we evaluate a baseline trained with classic label smoothing (LS), using the same loss structure but no teacher.
-We also rely on RCD, which retains soft targets but replaces the teacher’s logits with uniform noise.
-RCD preserves any label-smoothing benefit while removing semantic content.
-Across all metrics, we find that LS and RCD match or exceed KD in accuracy, yet exhibit no increase in functional similarity with the teacher, particularly on incorrect predictions.
-This confirms that KD’s asymmetric error transfer arises from the specific structure of the teacher’s logits, not from softening per se.
+We also rely on RCD, which retains the distillation form but replaces the teacher distribution with a class-uniform target ().
+RCD therefore preserves the label-smoothing effect while removing teacher-specific semantic content.
+Across all metrics, we find that LS and RCD match or exceed KD in accuracy, yet exhibit no increase in functional similarity with the teacher, particularly on incorrect predictions when accuracy using RCD is preserved.
+This supports the interpretation that KD’s asymmetric error transfer is driven by the specific structure of the teacher logits, rather than softening per se.
 | Metrics | Control | Knowledge Distillation | Random Control Distillation | Label Smoothing | ||||||
@@
 | Loss | 2.068 0.001 | 2.065 0.002 | 2.055 0.001 | 2.043 0.002 | 1.977 0.001 | 2.497 0.001 | 3.612 0.002 | 1.976 0.001 | 2.498 0.001 | 3.612 0.002 |
-Appendix E Knowledge Distillation to Smaller Student
-Justification:
-This setup allows for an analysis of Knowledge Distillation where the student model is smaller than the teacher model, as expected in practice.
-Caveat:
-Although this moves away from our traditional experiential setup where the student can perfectly match the teacher, we use this example to show how transfer works between a larger teacher to a smaller student.
-It is important to note that using a smaller student introduces uncertainty on if the student capacity is a bottleneck to knowledge transfer.
-However, given that in practice Knowledge Distillation is used in this setting we show how our fundamental insights from the self distillation case transfer to other cases of dilatation.
-Our study of using a smaller students is not exhaustive but demonstrative and verifies the findings presented in the main body of the paper, and the utility of our initial experimental setup.
-Other than the architecture’s implicit bias towards the problem, which affects its performance (loss and accuracy), there are no confounding factors that could influence Knowledge Distillation.
-E.1 TinyImageNet ResNet50 Teacher to ResNet18 Student
+Appendix C Knowledge Distillation to Smaller Student
+In this section we present the raw results for the other teacher seeds explored in this paper (1-2) to show that our findings are not the effect of a single seed but are general across teacher models and training seeds.
+C.1 TinyImageNet ResNet50 Teacher to ResNet18 Student
 Training Settings:
@@
 The ResNet18 student model was trained under the same conditions.
-Findings:
-We observe a low train loss for the teacher model circa 0.0014 with a high train accuracy circa 0.9998; see Table 15.
-This low train loss corresponds as expected, with no significant knowledge transfer across alpha values; see Tables 16, 17, 18 and 19.
-This result is as expected from the results and intuition presented in the results of the main body of the paper.
-It highlights how this finding generalises to the practical KD environment.
 | Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
@@
 | 2 | 0.001436 | 0.999800 | 2.051024 | 0.610600 |
-| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
-| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
-| Activation Distance | 0.548 0.000 | 0.548 0.000 | 0.547 0.000 | 0.547 0.000 | 0.565 0.000 | 0.651 0.000 | 0.828 0.000 |
-| Rank Disagreement | 0.987 0.000 | 0.987 0.000 | 0.987 0.000 | 0.987 0.000 | 0.990 0.000 | 0.990 0.000 | 0.991 0.000 |
-| Prediction Disagreement | 0.498 0.001 | 0.497 0.000 | 0.497 0.001 | 0.497 0.000 | 0.512 0.001 | 0.493 0.001 | 0.754 0.000 |
-| JS Divergence | 0.281 0.000 | 0.281 0.000 | 0.280 0.000 | 0.281 0.000 | 0.330 0.000 | 0.400 0.000 | 0.599 0.000 |
-| Accuracy | 0.503 0.001 | 0.504 0.001 | 0.504 0.000 | 0.503 0.000 | 0.493 0.000 | 0.512 0.000 | 0.236 0.000 |
-| Loss | 2.604 0.001 | 2.602 0.002 | 2.594 0.001 | 2.589 0.001 | 2.434 0.001 | 2.641 0.001 | 4.684 0.002 |
 | Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
@@
 | Loss | 2.604 0.001 | 2.602 0.001 | 2.594 0.001 | 2.587 0.001 | 2.434 0.001 | 2.641 0.001 | 4.684 0.002 |
-| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
-| KD 0.1 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
-| KD 0.5 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
-| KD 0.9 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
-E.2 Tiny Shakespeare Nano-GPT Teacher to Pico-GPT Student
+C.2 ImageNet ResNet50 Teacher to ResNet18 Student
 Training Settings:
-The Nano-GPT Teacher is a GPT2-style transformer with an embedding dimension of 384, a vocabulary size of 65, six attention heads, six transformer blocks, a dropout of 0.200, and a block size of 256.
-The Pico-GPT student has an embedding dimension of 192, halving the internal width of the model; all other model settings are the same as the teacher.
-The teacher and student are trained on the Tiny Shakespeare dataset, with the first 90% used for training and the last 10% used for testing.
-The dataset was tokenised via a character tokeniser, and the model was trained auto-regressively to predict the next character token.
-The teacher and student are trained with the Adam optimiser with a learning rate of 3e-4 with a batch size of 64 for 5000 iterations.
-The student models are trained with the same seeds and data orders from seeds 10 to 19 for the 10 models used for averaging.
-This is repeated for the three teachers trained on seeds 0 to 2.
-Justification:
-This setup allows for an analysis of Knowledge Distillation where the student model is smaller than the teacher model, as expected in practice.
-It is not exhaustive but demonstrative that the findings we present in the main body of the paper generalise to this case.
-Other than the architecture’s implicit bias towards the problem, which affects its performance (loss and accuracy), no confounding factors could influence Knowledge Distillation.
-Findings:
-We observe a high train loss for the teacher model circa 0.86 with a high train accuracy circa 0.72; see Table 20.
-This high train loss corresponds as expected with a substantial knowledge transfer which increases as alpha increases, see Tables 96, 97, 98 and 99.
-This substantial knowledge transfer coincides with an asymmetric payoff in prediction agreement, strongly favouring incorrect predictions, see Figure 65(a).
-This result is as expected from the results and intuition presented in the results of the main body of the paper and highlights how this finding generalises.
+A pre-trained ResNet50 model taken from PyTorch with a top-1-accuracy of 80.858 and a top-5-accuracy of 95.434222https://docs.pytorch.org/vision/main/models/generated/torchvision.models.resnet50.html#torchvision.models.ResNet50_Weights is used as the teacher.
+As Pytorch only provides one set of pre-trained model weights there is only one teacher seed for this experiment.
+The ResNet18 student was trained on ImageNet (Deng et al., 2009) using the FFCV setup (Leclerc et al., 2023), where 100% of the training images were compressed to a JPEG with 90% quality.
+The data was normalized with a mean of (0.485, 0.456, 0.406) and a standard deviation of (0.229, 0.224, 0.225).
+The model utilised BlurPools (Zhang, 2019) within the convolutional layers, and was trained for 56 epochs, with a batch size of 1024 using SGD, momentum of 0.9, weight decay of 5e-5, a learning rate of 0.5 using a cyclic scheduler with a learning rate step ratio of 0,1 and step length of 30.
+The learning rate peak was at epoch 2.
+The input resolution started at 160 by 160, and started to ramp up to 192 by 192 at epoch 41 and ended at 192 by 192 at epoch 48.
+C.2.1 The Effect Of Temperature
+This section presents the raw results for temperature of 2 on the ImageNet dataset on seed 0, due to computational constraints this test was only conducted on seed 0.
+| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
+| SIDDO | KD 0.1 | KD 0.5 | KD 0.9 | Rand KD 0.1 | Rand KD 0.5 | Rand KD 0.9 | |
+| Activation Distance | 0.420 0.001 | 0.310 0.001 | 0.251 0.001 | 0.221 0.001 | 0.304 0.001 | 0.247 0.001 | 0.282 0.001 |
+| Rank Disagreement | 0.997 0.000 | 0.997 0.000 | 0.997 0.000 | 0.996 0.000 | 0.997 0.000 | 0.997 0.000 | 0.997 0.000 |
+| Prediction Disagreement | 0.263 0.002 | 0.257 0.001 | 0.256 0.001 | 0.264 0.001 | 0.259 0.001 | 0.273 0.001 | 0.310 0.001 |
+| JS Divergence | 0.259 0.000 | 0.160 0.000 | 0.101 0.000 | 0.081 0.000 | 0.151 0.001 | 0.096 0.000 | 0.116 0.001 |
+| Accuracy | 0.681 0.001 | 0.685 0.001 | 0.685 0.001 | 0.679 0.001 | 0.684 0.001 | 0.671 0.001 | 0.641 0.001 |
+| Loss | 1.303 0.006 | 1.492 0.008 | 1.718 0.012 | 1.922 0.013 | 1.532 0.009 | 1.919 0.015 | 3.022 0.014 |
+C.3 TinyShakespeare Nano-GPT Teacher to Pico-GPT Student
 | Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
@@
 | 2 | 0.861098 | 0.721140 | 1.562137 | 0.573033 |
-| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
-| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
-| Activation Distance | 0.202 0.000 | 0.198 0.000 | 0.181 0.000 | 0.172 0.000 | 0.221 0.000 | 0.399 0.000 | 0.663 0.000 |
-| Rank Disagreement | 0.915 0.000 | 0.915 0.000 | 0.912 0.000 | 0.911 0.000 | 0.939 0.000 | 0.944 0.000 | 0.950 0.000 |
-| Prediction Disagreement | 0.252 0.000 | 0.247 0.000 | 0.226 0.000 | 0.214 0.000 | 0.252 0.000 | 0.253 0.001 | 0.272 0.001 |
-| JS Divergence | 0.056 0.000 | 0.054 0.000 | 0.047 0.000 | 0.043 0.000 | 0.075 0.000 | 0.203 0.000 | 0.451 0.000 |
-| Accuracy | 0.571 0.000 | 0.572 0.000 | 0.575 0.000 | 0.574 0.000 | 0.571 0.000 | 0.570 0.000 | 0.561 0.000 |
-| Loss | 1.473 0.002 | 1.471 0.002 | 1.472 0.001 | 1.496 0.002 | 1.483 0.001 | 1.870 0.001 | 3.017 0.002 |
 | Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
@@
 | Loss | 1.475 0.001 | 1.470 0.001 | 1.471 0.002 | 1.491 0.002 | 1.482 0.001 | 1.865 0.002 | 3.017 0.001 |
-| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
-| KD 0.1 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✗ | ✗✗✓ |
-| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ |
-| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ |
-Appendix F Vision Results
-F.1 TinyImageNet
+C.3.1 The Effect Of Temperature
+This section provides the raw results for temperature 2 and 4 on the TinyShakespeare dataset.
+| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
+| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
+| Activation Distance | 0.202 0.000 | 0.197 0.000 | 0.183 0.000 | 0.181 0.000 | 0.213 0.000 | 0.305 0.001 | 0.617 0.000 |
+| Rank Disagreement | 0.915 0.000 | 0.907 0.000 | 0.896 0.000 | 0.892 0.000 | 0.94 0.000 | 0.945 0.000 | 0.95 0.000 |
+| Prediction Disagreement | 0.252 0.000 | 0.25 0.000 | 0.235 0.000 | 0.23 0.000 | 0.252 0.000 | 0.253 0.000 | 0.27 0.000 |
+| JS Divergence | 0.056 0.000 | 0.053 0.000 | 0.047 0.000 | 0.047 0.000 | 0.072 0.000 | 0.152 0.000 | 0.403 0.000 |
+| Accuracy | 0.571 0.000 | 0.572 0.000 | 0.572 0.000 | 0.569 0.000 | 0.571 0.000 | 0.571 0.000 | 0.562 0.000 |
+| Loss | 1.473 0.002 | 1.513 0.003 | 1.571 0.002 | 1.622 0.002 | 1.493 0.001 | 1.736 0.001 | 2.732 0.001 |
+| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
+| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
+| Activation Distance | 0.201 0.000 | 0.195 0.000 | 0.181 0.000 | 0.179 0.000 | 0.209 0.000 | 0.298 0.000 | 0.609 0.000 |
+| Rank Disagreement | 0.916 0.000 | 0.907 0.000 | 0.896 0.000 | 0.892 0.000 | 0.94 0.000 | 0.945 0.000 | 0.95 0.000 |
+| Prediction Disagreement | 0.258 0.001 | 0.254 0.000 | 0.24 0.000 | 0.236 0.000 | 0.256 0.000 | 0.258 0.000 | 0.279 0.000 |
+| JS Divergence | 0.055 0.000 | 0.052 0.000 | 0.047 0.000 | 0.046 0.000 | 0.071 0.000 | 0.15 0.000 | 0.401 0.000 |
+| Accuracy | 0.571 0.000 | 0.571 0.000 | 0.572 0.000 | 0.569 0.000 | 0.572 0.000 | 0.571 0.000 | 0.56 0.000 |
+| Loss | 1.474 0.002 | 1.512 0.003 | 1.569 0.002 | 1.613 0.003 | 1.489 0.001 | 1.732 0.001 | 2.739 0.001 |
+| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
+| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
+| Activation Distance | 0.201 0.000 | 0.195 0.000 | 0.181 0.000 | 0.18 0.000 | 0.21 0.000 | 0.301 0.000 | 0.615 0.000 |
+| Rank Disagreement | 0.915 0.000 | 0.906 0.000 | 0.896 0.000 | 0.892 0.000 | 0.94 0.000 | 0.945 0.000 | 0.95 0.000 |
+| Prediction Disagreement | 0.251 0.001 | 0.247 0.000 | 0.235 0.000 | 0.23 0.000 | 0.249 0.000 | 0.252 0.000 | 0.274 0.000 |
+| JS Divergence | 0.055 0.000 | 0.052 0.000 | 0.046 0.000 | 0.046 0.000 | 0.071 0.000 | 0.15 0.000 | 0.403 0.000 |
+| Accuracy | 0.571 0.000 | 0.571 0.000 | 0.571 0.000 | 0.569 0.000 | 0.572 0.000 | 0.571 0.000 | 0.56 0.000 |
+| Loss | 1.474 0.002 | 1.513 0.001 | 1.576 0.001 | 1.619 0.003 | 1.489 0.001 | 1.732 0.001 | 2.739 0.001 |
+| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
+| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
+| Activation Distance | 0.202 0.000 | 0.199 0.000 | 0.189 0.000 | 0.193 0.000 | 0.206 0.000 | 0.262 0.000 | 0.568 0.001 |
+| Rank Disagreement | 0.915 0.000 | 0.893 0.000 | 0.88 0.000 | 0.876 0.000 | 0.94 0.000 | 0.945 0.000 | 0.951 0.000 |
+| Prediction Disagreement | 0.252 0.000 | 0.251 0.001 | 0.244 0.000 | 0.245 0.000 | 0.253 0.000 | 0.253 0.000 | 0.27 0.000 |
+| JS Divergence | 0.056 0.000 | 0.054 0.000 | 0.05 0.000 | 0.051 0.000 | 0.067 0.000 | 0.127 0.000 | 0.362 0.000 |
+| Accuracy | 0.571 0.000 | 0.57 0.000 | 0.568 0.000 | 0.562 0.000 | 0.572 0.000 | 0.571 0.000 | 0.562 0.000 |
+| Loss | 1.473 0.002 | 1.528 0.002 | 1.592 0.002 | 1.663 0.002 | 1.491 0.002 | 1.68 0.000 | 2.544 0.002 |
+| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
+| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
+| Activation Distance | 0.201 0.000 | 0.196 0.000 | 0.188 0.000 | 0.191 0.000 | 0.203 0.000 | 0.256 0.000 | 0.562 0.000 |
+| Rank Disagreement | 0.916 0.000 | 0.893 0.000 | 0.88 0.000 | 0.876 0.000 | 0.94 0.000 | 0.945 0.000 | 0.951 0.000 |
+| Prediction Disagreement | 0.258 0.001 | 0.256 0.000 | 0.25 0.000 | 0.249 0.000 | 0.256 0.000 | 0.258 0.000 | 0.278 0.000 |
+| JS Divergence | 0.055 0.000 | 0.052 0.000 | 0.049 0.000 | 0.05 0.000 | 0.066 0.000 | 0.126 0.000 | 0.361 0.000 |
+| Accuracy | 0.571 0.000 | 0.57 0.000 | 0.568 0.000 | 0.563 0.000 | 0.571 0.000 | 0.571 0.000 | 0.561 0.000 |
+| Loss | 1.474 0.002 | 1.528 0.002 | 1.59 0.002 | 1.653 0.003 | 1.489 0.001 | 1.677 0.001 | 2.55 0.002 |
+| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
+| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
+| Activation Distance | 0.201 0.000 | 0.197 0.000 | 0.189 0.000 | 0.192 0.000 | 0.204 0.000 | 0.259 0.000 | 0.567 0.000 |
+| Rank Disagreement | 0.915 0.000 | 0.893 0.000 | 0.879 0.000 | 0.876 0.000 | 0.94 0.000 | 0.945 0.000 | 0.951 0.000 |
+| Prediction Disagreement | 0.251 0.001 | 0.25 0.001 | 0.245 0.001 | 0.245 0.000 | 0.25 0.001 | 0.253 0.000 | 0.275 0.000 |
+| JS Divergence | 0.055 0.000 | 0.053 0.000 | 0.049 0.000 | 0.05 0.000 | 0.066 0.000 | 0.127 0.000 | 0.363 0.000 |
+| Accuracy | 0.571 0.000 | 0.57 0.000 | 0.568 0.000 | 0.562 0.000 | 0.571 0.000 | 0.571 0.000 | 0.561 0.000 |
+| Loss | 1.474 0.002 | 1.53 0.001 | 1.594 0.002 | 1.658 0.002 | 1.489 0.001 | 1.677 0.001 | 2.55 0.002 |
+Appendix D Vision Results
+D.1 TinyImageNet
 Training Settings:
@@
 The VGG19 and VGG19 with RandAugment has the same setup as the ResNet50 and ResNet50 with RandAugment respectively however it was trained with momentum of 0.9.
-F.1.1 ResNet50
-Findings:
-For the ResNet50 on TinyImageNet, we observe that the teacher seeds, Table 25, obtain a low train loss of 0.001 and a train accuracy of 0.99.
-This train performance coincides with a test accuracy of circa 0.60, resulting in a generalisation gap of circa 0.39.
-For an alpha of 0.1, Table 29, we observe no significant knowledge transfer across all metrics except for Rank Disagreement with teacher seed 0.
-It has statistically significant transfer, but the increased similarity is extremely marginal, as observed with SIDDO and KD 0.1 having the same value to 3 significant figures, see Table 26.
-With this, we see a marginal prediction agreement of less than 0.5% for correct and incorrect predictions across teacher seeds, Figure 23(a).
-For alpha 0.5 and 0.9, we observe significant knowledge transfer for all bar Prediction Disagreement with alpha of 0.5 and 0.9 for teacher seed 2.
-However, this transfer is marginal, Tables 26, 27 and 28, and we observe a prediction agreement of less than 0.5% for correct and incorrect predictions across teacher seeds, Figure 17(a).
-| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
-| 0 | 0.001426 | 0.999800 | 2.070590 | 0.605300 |
-| 1 | 0.001393 | 0.999800 | 2.051494 | 0.607900 |
-| 2 | 0.001436 | 0.999800 | 2.051024 | 0.610600 |
-| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
-| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
-| Activation Distance | 0.157 0.001 | 0.157 0.001 | 0.156 0.001 | 0.155 0.000 | 0.343 0.000 | 0.581 0.000 | 0.791 0.000 |
-| Rank Disagreement | 0.939 0.000 | 0.939 0.000 | 0.939 0.000 | 0.939 0.000 | 0.980 0.000 | 0.984 0.000 | 0.984 0.000 |
-| Prediction Disagreement | 0.153 0.001 | 0.152 0.001 | 0.151 0.001 | 0.151 0.001 | 0.190 0.001 | 0.214 0.000 | 0.324 0.000 |
-| JS Divergence | 0.040 0.000 | 0.040 0.000 | 0.039 0.000 | 0.039 0.000 | 0.171 0.000 | 0.333 0.000 | 0.533 0.000 |
-| Accuracy | 0.605 0.001 | 0.605 0.000 | 0.604 0.001 | 0.605 0.001 | 0.607 0.000 | 0.606 0.001 | 0.580 0.000 |
-| Loss | 2.068 0.001 | 2.065 0.002 | 2.055 0.001 | 2.043 0.002 | 1.977 0.001 | 2.497 0.001 | 3.612 0.002 |
+D.1.1 ResNet50
+Here we present the results of the ResNet50 on TinyImageNet in the Self Distillation regime for teacher seeds 1 and 2 - the corresponding analysis can be found in the main body in Section 4.1.
 | Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
@@
 | Loss | 2.054 0.002 | 2.050 0.002 | 2.040 0.003 | 2.025 0.002 | 1.967 0.001 | 2.494 0.001 | 3.602 0.002 |
-| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
-| KD 0.1 | ✗✗✗ | ✓✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
-| KD 0.5 | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
-| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✗ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
-F.1.2 ResNet50 with RandAugment
-Findings:
-For the ResNet50 on TinyImageNet with RandAugment, we observe that the teacher seeds, Table 25, obtain a high train loss and a train accuracy of circa 0.84.
-This train performance coincides with a test accuracy of circa 0.64, resulting in a generalisation gap of circa 0.2.
-We observe significant knowledge transfer for all alpha values with a strong asymmetric transfer of knowledge favouring incorrect predictions as shown in Table 34 and Figure 26(a), respectively.
-However, it is important to note that despite significant and substantial knowledge transfer, we do not see any improvement in test accuracy over the control and random controls.
-| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
-| 0 | 0.672748 | 0.840410 | 1.620552 | 0.638800 |
-| 1 | 0.678245 | 0.839200 | 1.629393 | 0.641800 |
-| 2 | 0.667570 | 0.840750 | 1.624969 | 0.641100 |
-| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
-| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
-| Activation Distance | 0.193 0.000 | 0.183 0.000 | 0.150 0.000 | 0.131 0.000 | 0.245 0.001 | 0.501 0.001 | 0.781 0.000 |
-| Rank Disagreement | 0.959 0.000 | 0.957 0.000 | 0.948 0.000 | 0.943 0.000 | 0.975 0.000 | 0.981 0.000 | 0.987 0.000 |
-| Prediction Disagreement | 0.196 0.001 | 0.188 0.001 | 0.154 0.001 | 0.136 0.001 | 0.195 0.001 | 0.240 0.001 | 0.572 0.001 |
-| JS Divergence | 0.058 0.000 | 0.052 0.000 | 0.036 0.000 | 0.028 0.000 | 0.094 0.000 | 0.266 0.000 | 0.563 0.000 |
-| Accuracy | 0.640 0.000 | 0.643 0.001 | 0.644 0.000 | 0.642 0.000 | 0.646 0.001 | 0.657 0.001 | 0.400 0.001 |
-| Loss | 1.619 0.003 | 1.600 0.001 | 1.578 0.001 | 1.577 0.001 | 1.551 0.001 | 1.984 0.002 | 4.211 0.001 |
+D.1.2 ResNet50 with RandAugment
+Here we present the results of the ResNet50 on TinyImageNet using RandAugment for student and teacher in the Self Distillation regime for teacher seeds 1 and 2 - the corresponding analysis can be found in the main body in Section 4.1.
 | Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
@@
 | Loss | 1.621 0.002 | 1.606 0.001 | 1.581 0.001 | 1.582 0.001 | 1.552 0.001 | 1.982 0.002 | 4.180 0.002 |
-| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
-| KD 0.1 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
-| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
-| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
-F.1.3 VGG19
+D.1.3 VGG19
 Findings:
 For the VGG19 on the TinyImageNet, we observe a low train loss of circa and a train accuracy of 0.9998.
-As expected, given our results and discussion in the main body of the paper on the ResNet50, we see no significant transfer until an alpha of 0.9.
-With teacher seed 0 and 2 with an alpha of 0.9, we record significant transfer for Activation Distance and for teacher seed 0 on JS Divergence, as seen in Table 39.
-When we observe knowledge transfer with an alpha of 0.9, we observe a slight preference for positive agreement of test prediction; however, the results have a large SEM, and the amount of agreement is less than 0.5%, making the results less reliable and insignificant in either transfer direction.
+As expected, given our results and discussion in the main body of the paper on the ResNet50, we see no statistically supported until an of 0.9.
+With teacher seed 0 and 2 with an of 0.9, we record statistically supported transfer for Activation Distance and for teacher seed 0 on JS Divergence, as seen in Table 55.
+When we observe knowledge transfer with an of 0.9, we observe a slight preference for positive agreement of test prediction; however, the results have a large SEM, and the amount of agreement is less than 0.5%, making the results less reliable and insignificant in either transfer direction.
 | Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
@@
 | KD 0.5 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
-| KD 0.9 | ✓✗✓ | ✗✗✗ | ✗✗✗ | ✓✗✗ | ✗✗✗ | ✗✗✗ |
-F.1.4 VGG19 with RandAugment
+| KD 0.9 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
+D.1.4 VGG19 with RandAugment
 Findings:
 For the VGG19 on the TinyImageNet with RandAugment, we observe a high train loss of circa and a train accuracy of circa 0.93.
-As expected, given the results on the RandAugment ResNet50 that we present in the main body of the paper, we see substantial transfer across all alpha values; see Tables 41, 42, 43 and 44.
-This substantial and significant transfer of knowledge, as expected, coincides with a strong asymmetric transfer of knowledge favouring incorrect predictions, as shown in Figure 32(a).
+As expected, given the results on the RandAugment ResNet50 that we present in the main body of the paper, we see statistically supported transfer across all values; see Tables 57, 58, 59 and 60.
+This substantial and statistically supported transfer of knowledge, as expected, coincides with a strong asymmetric transfer of knowledge favouring incorrect predictions, as shown in Figure 19.
 | Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
@@
 | KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ |
-F.2 CIFAR10
+D.2 SVHN Dataset
 Training Settings:
-All CIFAR10 architectures are trained with Adam optimiser with a learning rate of 0.001 and a batch size of 256 for 100 epochs.
-All data is normalised with a mean of 0.5 and a standard deviation of 0.5.
-The student vision architectures are trained with the same seeds and data orders from seeds 10-19 for the 10 models used for averaging.
-As aligned with all experiments we conduct, this is repeated for the three teachers trained on seeds 0-2.
-Justification:
-This setup allows for a fair analysis of Knowledge Distillation as its role is isolated in the training process.
-Other than the architecture’s implicit bias towards the problem, which affects its performance (loss and accuracy), there are no confounding factors that could influence Knowledge Distillation.
-Findings:
-We find that the teacher models often significantly transfer knowledge to the student model, and this coincides with the teacher’s high loss on the training dataset.
-The ResNet has the lowest loss and no transfer, the VGG has a higher loss and some transfer, and the ViT has the highest loss and the most transfer.
-However, when knowledge is transferred, it often has a negative asymmetric payoff towards agreement between the teacher and the student on incorrect predictions.
-F.2.1 ResNet18
-Findings:
-For the ResNet18 on CIFAR10, we observe that the teacher seeds, Table 45, obtain a very low train loss of and a train accuracy of 1.
-This train performance coincides with a high test accuracy of circa 0.86, resulting in a generalisation gap of circa 0.14.
-Table 49 shows no significant knowledge transfer across teacher seeds.
-Due to the low train loss on the teacher seed, the teacher model is a nearly identical representation of the training labels, meaning there is low utility in the teacher model.
-As we observe, the controls of the models trained in the SIDDO condition is functionally different from the teacher, Tables 46, 47 and 48; despite having the same initialisation and only changing the data order, it is not a surprise that Knowledge Distillation in the setup does not add anything as the teacher is essentially the label, and thus creates a similar setup to the SIDDO condition.
-| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
-| 0 | 0.000010 | 1.000000 | 0.869184 | 0.862100 |
-| 1 | 0.000006 | 1.000000 | 0.833735 | 0.867200 |
-| 2 | 0.000030 | 1.000000 | 0.739927 | 0.867000 |
-| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
-| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
-| Activation Distance () | 0.174 | 0.175 | 0.172 | 0.174 | 0.244 | 0.538 | 0.843 |
-| Rank Disagreement () | 0.659 | 0.659 | 0.656 | 0.655 | 0.795 | 0.802 | 0.807 |
-| Prediction Disagreement () | 0.128 | 0.129 | 0.127 | 0.128 | 0.131 | 0.143 | 0.150 |
-| JS Divergence () | 0.070 | 0.070 | 0.069 | 0.068 | 0.097 | 0.229 | 0.432 |
-| Accuracy () | 0.861 | 0.862 | 0.862 | 0.862 | 0.865 | 0.856 | 0.854 |
-| Loss () | 0.961 | 0.903 | 0.895 | 0.827 | 0.539 | 0.902 | 1.772 |
-| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
-| SIDDO | 0.1 | 0.5 | 0.900 | 0.1 | 0.5 | 0.900 | |
-| Activation Distance () | 0.167 | 0.164 | 0.165 | 0.165 | 0.240 | 0.533 | 0.841 |
-| Rank Disagreement () | 0.653 | 0.649 | 0.650 | 0.650 | 0.796 | 0.803 | 0.807 |
-| Prediction Disagreement () | 0.122 | 0.120 | 0.121 | 0.120 | 0.126 | 0.134 | 0.139 |
-| JS Divergence () | 0.066 | 0.065 | 0.065 | 0.064 | 0.095 | 0.226 | 0.430 |
-| Accuracy () | 0.865 | 0.867 | 0.866 | 0.867 | 0.866 | 0.860 | 0.859 |
-| Loss () | 0.858 | 0.877 | 0.824 | 0.816 | 0.533 | 0.896 | 1.767 |
-| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
-| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
-| Activation Distance () | 0.166 | 0.169 | 0.167 | 0.172 | 0.242 | 0.533 | 0.839 |
-| Rank Disagreement () | 0.646 | 0.647 | 0.638 | 0.646 | 0.799 | 0.803 | 0.805 |
-| Prediction Disagreement () | 0.122 | 0.124 | 0.124 | 0.127 | 0.132 | 0.140 | 0.142 |
-| JS Divergence () | 0.065 | 0.066 | 0.064 | 0.067 | 0.096 | 0.226 | 0.429 |
-| Accuracy () | 0.865 | 0.864 | 0.864 | 0.861 | 0.862 | 0.857 | 0.857 |
-| Loss () | 0.892 | 0.887 | 0.803 | 0.798 | 0.549 | 0.900 | 1.769 |
-| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
-| KD 0.1 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
-| KD 0.5 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
-| KD 0.9 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
-F.2.2 VGG19
-Findings:
-For the VGG19 on CIFAR10, we observe that the teacher seeds, Table 50, obtain a low train loss of circa 0.01 and a train accuracy of approximately 0.996.
-This train performance coincides with a high test accuracy of circa 0.86, resulting in a generalisation gap of circa 0.14.
-Table 54 shows a significant knowledge transfer with regard to Rank Disagreement for all teacher seeds when alpha is at 0.9.
-At alpha 0.9 for teacher seed 0 and 2, there is an increase in agreement between the student and teacher on incorrect predictions over the correct predictions, Figure 38(a), which corresponds with the knowledge transfer.
-This result coincides with teachers seed 0 and 2 having a higher train loss than teacher seed 1, indicating that the teacher train loss plays an important role in knowledge transfer.
-For teacher seed 1, Figure 38(a), there is no significant increase in correct or incorrect prediction agreement between the student model and the teacher due to the deviation in the SEM.
-| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
-| 0 | 0.011608 | 0.996760 | 0.858675 | 0.863900 |
-| 1 | 0.009228 | 0.997080 | 0.798530 | 0.860800 |
-| 2 | 0.012352 | 0.996420 | 0.801562 | 0.867100 |
-| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
-| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
-| Activation Distance () | 0.206 | 0.199 | 0.203 | 0.197 | 0.264 | 0.541 | 0.842 |
-| Rank Disagreement () | 0.701 | 0.705 | 0.658 | 0.640 | 0.811 | 0.819 | 0.819 |
-| Prediction Disagreement () | 0.152 | 0.147 | 0.151 | 0.146 | 0.148 | 0.146 | 0.150 |
-| JS Divergence () | 0.090 | 0.085 | 0.086 | 0.083 | 0.109 | 0.230 | 0.429 |
-| Accuracy () | 0.864 | 0.869 | 0.867 | 0.869 | 0.870 | 0.871 | 0.868 |
-| Loss () | 0.849 | 0.725 | 0.676 | 0.649 | 0.562 | 0.880 | 1.762 |
-| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
-| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
-| Activation Distance () | 0.199 | 0.202 | 0.202 | 0.201 | 0.263 | 0.543 | 0.842 |
-| Rank Disagreement () | 0.726 | 0.684 | 0.662 | 0.639 | 0.803 | 0.801 | 0.810 |
-| Prediction Disagreement () | 0.147 | 0.150 | 0.150 | 0.149 | 0.148 | 0.149 | 0.153 |
-| JS Divergence () | 0.086 | 0.087 | 0.086 | 0.085 | 0.107 | 0.230 | 0.428 |
-| Accuracy () | 0.868 | 0.866 | 0.865 | 0.866 | 0.870 | 0.869 | 0.866 |
-| Loss () | 0.799 | 0.735 | 0.680 | 0.666 | 0.562 | 0.887 | 1.762 |
-| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
-| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
-| Activation Distance () | 0.196 | 0.199 | 0.196 | 0.193 | 0.258 | 0.541 | 0.844 |
-| Rank Disagreement () | 0.672 | 0.649 | 0.633 | 0.602 | 0.809 | 0.817 | 0.816 |
-| Prediction Disagreement () | 0.142 | 0.146 | 0.143 | 0.141 | 0.142 | 0.143 | 0.149 |
-| JS Divergence () | 0.084 | 0.086 | 0.083 | 0.081 | 0.106 | 0.229 | 0.429 |
-| Accuracy () | 0.870 | 0.864 | 0.868 | 0.867 | 0.871 | 0.871 | 0.867 |
-| Loss () | 0.801 | 0.734 | 0.665 | 0.639 | 0.560 | 0.884 | 1.762 |
-| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
-| KD 0.1 | ✗✗✗ | ✗✓✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
-| KD 0.5 | ✗✗✗ | ✓✓✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
-| KD 0.9 | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✓✗✗ | ✗✗✗ | ✗✗✗ |
-F.2.3 ViT
-Findings:
-For the ViT on CIFAR10, we observe that the teacher seeds, Table 55, obtain a high train loss of and a train accuracy of approximately 0.98.
-This train performance coincides with a test accuracy of circa 0.63, resulting in a generalisation gap of circa 0.35.
-Table 59 shows a significant knowledge transfer on all teacher seeds when alpha is 0.5 and 0.9.
-For teacher seed 0 and 1 using alpha at 0.9, where there is sizeable knowledge transfer, we observe an asymmetric knowledge transfer favouring negative transfer in Figure 41(a).
-| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
-| 0 | 0.043291 | 0.988260 | 1.864339 | 0.626900 |
-| 1 | 0.056539 | 0.983160 | 1.772490 | 0.634200 |
-| 2 | 0.046902 | 0.987100 | 1.714442 | 0.649600 |
-| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
-| SIDDO | 0.1 | 0.5 | 0.900 | 0.1 | 0.5 | 0.900 | |
-| Activation Distance () | 0.491 | 0.487 | 0.473 | 0.470 | 0.496 | 0.611 | 0.793 |
-| Rank Disagreement () | 0.734 | 0.730 | 0.724 | 0.722 | 0.808 | 0.812 | 0.817 |
-| Prediction Disagreement () | 0.385 | 0.383 | 0.374 | 0.373 | 0.383 | 0.380 | 0.386 |
-| JS Divergence () | 0.201 | 0.198 | 0.189 | 0.186 | 0.206 | 0.277 | 0.411 |
-| Accuracy () | 0.634 | 0.634 | 0.641 | 0.637 | 0.640 | 0.641 | 0.627 |
-| Loss () | 1.773 | 1.695 | 1.52 | 1.451 | 1.258 | 1.351 | 1.943 |
-| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
-| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
-| Activation Distance () | 0.485 | 0.477 | 0.461 | 0.455 | 0.489 | 0.609 | 0.791 |
-| Rank Disagreement () | 0.733 | 0.728 | 0.717 | 0.714 | 0.806 | 0.808 | 0.816 |
-| Prediction Disagreement () | 0.382 | 0.375 | 0.367 | 0.363 | 0.379 | 0.380 | 0.382 |
-| JS Divergence () | 0.198 | 0.193 | 0.182 | 0.178 | 0.202 | 0.275 | 0.410 |
-| Accuracy () | 0.637 | 0.643 | 0.644 | 0.648 | 0.643 | 0.636 | 0.630 |
-| Loss () | 1.781 | 1.668 | 1.466 | 1.366 | 1.253 | 1.359 | 1.942 |
-| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
-| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
-| Activation Distance () | 0.476 | 0.468 | 0.459 | 0.456 | 0.486 | 0.612 | 0.797 |
-| Rank Disagreement () | 0.730 | 0.725 | 0.720 | 0.718 | 0.806 | 0.811 | 0.817 |
-| Prediction Disagreement () | 0.372 | 0.366 | 0.363 | 0.360 | 0.371 | 0.374 | 0.375 |
-| JS Divergence () | 0.195 | 0.189 | 0.183 | 0.180 | 0.201 | 0.277 | 0.413 |
-| Accuracy () | 0.636 | 0.641 | 0.644 | 0.639 | 0.637 | 0.635 | 0.631 |
-| Loss () | 1.788 | 1.673 | 1.498 | 1.458 | 1.282 | 1.361 | 1.942 |
-| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
-| KD 0.1 | ✗✓✓ | ✓✓✓ | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
-| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✓ | ✗✗✗ |
-| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
-F.3 SVHN Dataset
-Training Settings:
-All SVHN architectures are trained with Adam optimiser with a learning rate of 0.001 and a batch size of 256 for 100 epochs.
+All SVHN architectures are trained with Adam with a learning rate of 0.001 and a batch size of 256 for 100 epochs.
 All data is normalised with a mean of 0.5 and a standard deviation of 0.5.
@@
 We repeated this, in line with our other experiments for the three teachers trained on seeds 0-2.
-Justification:
-This setup allows for a fair analysis of Knowledge Distillation as its role is isolated in the training process.
-Other than the architecture’s implicit bias towards the problem, which affects its performance (loss and accuracy), there are no confounding factors that could influence Knowledge Distillation.
-Findings:
-The teacher models often significantly transfer knowledge to the student model. However, the knowledge transfer is often inconsistent, and when transferred, it often has an asymmetric negative payoff.
-F.3.1 ResNet18
-Findings:
-For the ResNet on SVHN, we observe that the teacher seeds, Table 60, obtain a range of train loss values of 0.000646, 0.000061 and 0.004657 for teacher seeds 0, 1, and 2, respectively.
-The train accuracies are approximately 0.99.
-This train performance coincides with a test accuracy of circa 0.95, resulting in a generalisation gap of circa 0.04.
-The teacher model with a higher training loss (seed 2) has significant knowledge transfer, see Table 64, for all functional similarity metrics across alpha values 0.1, 0.5 and 0.9, except for Prediction Disagreement when alpha was 0.1.
-In this case, we also observe a large asymmetric payoff in prediction agreement, significantly favouring incorrect predictions, Figure 44(a).
-Whereas teacher seed 0 has a train loss of 0.000061 and has no significant transfer with alpha values of 0.1 and 0.5.
-However, with an alpha of 0.9, it does have a significant transfer across metrics except for Prediction Disagreement, see Table 64.
-When alpha is 0.9, we observe an asymmetric payoff in prediction agreement, significantly favouring incorrect predictions.
-For teacher seed 0, which has a train loss of 0.000646, we observe significant knowledge transfer when alpha is 0.5 and 0.9, coinciding with an asymmetric payoff in prediction agreement, favouring incorrect predictions.
-| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
-| 0 | 0.000646 | 0.999850 | 0.381410 | 0.951829 |
-| 1 | 0.000061 | 0.999973 | 0.331054 | 0.952251 |
-| 2 | 0.004657 | 0.998580 | 0.309702 | 0.947104 |
-| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
-| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
-| Activation Distance () | 0.063 | 0.064 | 0.060 | 0.059 | 0.144 | 0.493 | 0.849 |
-| Rank Disagreement () | 0.696 | 0.688 | 0.684 | 0.681 | 0.800 | 0.798 | 0.802 |
-| Prediction Disagreement () | 0.045 | 0.046 | 0.043 | 0.042 | 0.042 | 0.043 | 0.046 |
-| JS Divergence () | 0.025 | 0.025 | 0.023 | 0.022 | 0.053 | 0.201 | 0.431 |
-| Accuracy () | 0.952 | 0.951 | 0.954 | 0.954 | 0.957 | 0.957 | 0.955 |
-| Loss () | 0.385 | 0.344 | 0.310 | 0.293 | 0.236 | 0.692 | 1.698 |
+D.2.1 ResNet18
 | Metrics | Control | Knowledge Distillation | Random Knowledge Distillation | ||||
@@
 | Loss () | 0.370 | 0.256 | 0.226 | 0.216 | 0.239 | 0.692 | 1.700 |
-| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
-| KD 0.1 | ✗✗✓ | ✗✗✓ | ✗✗✗ | ✗✗✓ | ✗✗✗ | ✗✗✗ |
-| KD 0.5 | ✗✗✓ | ✓✗✓ | ✗✗✓ | ✓✗✓ | ✗✗✗ | ✗✗✓ |
-| KD 0.9 | ✓✓✓ | ✓✓✓ | ✗✗✓ | ✓✓✓ | ✗✗✗ | ✗✗✓ |
-F.3.2 VGG19
+D.2.2 VGG19
 Findings:
-For the VGG19 on SVHN, we record a low train loss from we observe that the teacher seeds, Table 65, obtain a range of train loss values of 0.004511, 0.002757 and 0.00374 for teacher seeds 0, 1, and 2, respectively.
-The train accuracies are approximately 0.99.
-This train performance coincides with a test accuracy of circa 0.95, resulting in a generalisation gap of circa 0.04.
-The teacher model with a higher training loss (seed 2) has significant knowledge transfer, see Table 69, for only Rank Disagreement, across alpha values 0.1, 0.5 and 0.9.
-Due to limited statically significant functional transfer across metrics for this seed, we observe a small but inconsistent asymmetric payoff in prediction agreement, slightly favouring incorrect predictions, Figure 47(a).
-The story is very similar across the other teacher seeds; we see marginal functional transfer, and where a transfer is higher, we see negative transfer, but where it is marginal or largely insignificant, we see no preference for knowledge transfer, showing that in this case knowledge sharing can not be attributed to improved performance.
+For the VGG19 on SVHN, we record a low train loss from we observe that the teacher seeds, Table 63.
+The teacher model with a higher training loss (seed 2) has statistically supported knowledge transfer, see Table 67, for only Rank Disagreement, across values 0.1, 0.5 and 0.9.
+Due to the marginal statistically supported transfer across metrics for this seed, we observe a small but inconsistent asymmetric payoff in prediction agreement, slightly favouring incorrect predictions, Figure 20.
+Across the other teacher seeds; we see marginal statistically supported functional transfer, and where transfer is higher, we see negative transfer, but where it is marginal or largely insignificant, we see no preference for knowledge transfer, showing that in this case knowledge sharing can not be attributed to improved performance.
 | Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
@@
 | Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
-| KD 0.1 | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✗✓✗ | ✗✗✗ | ✗✗✗ |
-| KD 0.5 | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✗✓✗ | ✗✗✗ | ✗✗✗ |
-| KD 0.9 | ✗✓✗ | ✓✓✓ | ✗✗✗ | ✗✓✗ | ✗✗✗ | ✗✗✗ |
-F.3.3 ViT
+| KD 0.1 | ✗✗✗ | ✗✓✓ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
+| KD 0.5 | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
+| KD 0.9 | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✗✓✗ | ✗✗✗ | ✗✗✗ |
+D.2.3 ViT
 Findings:
-For the ViT on SVHN, we record a train loss from we observe that the teacher seeds, Table 70, obtain a range of train loss values of 0.018473, 0.019402 and 0.018580 for teacher seeds 0, 1, and 2, respectively.
-The train accuracies are approximately 0.99.
-This train performance coincides with a test accuracy of circa 0.85, resulting in a generalisation gap of circa 0.14.
-The teacher model with a higher training loss (seed 1) has significant knowledge transfer, see Table 74, for only Activation Distance, Rank Disagreement and JS Divergence across alpha values 0.5 and 0.9.
-In this case, we observe a small but inconsistent asymmetric payoff in prediction agreement, slightly favouring incorrect predictions, Figure 50(a).
-The story is very similar across the other teacher seeds; we see marginal functional transfer, and where a transfer is higher, we see negative transfer, but where it is marginal or largely insignificant, we see no real preference for knowledge transfer, showing that in this case knowledge sharing can not be attributed to improved performance.
+For ViT on SVHN, obtain a low train loss values across teacher seeds, Table 68, The teacher model with a higher training loss (seed 1) has statisically supported knowledge transfer, see Table 72, for only Activation Distance, Rank Disagreement and JS Divergence across values 0.5 and 0.9.
+In this case, we observe a small but inconsistent asymmetric payoff in prediction agreement, slightly favouring incorrect predictions, Figure 21.
 | Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
@@
 | Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
-| KD 0.1 | ✗✓✓ | ✗✗✓ | ✗✗✗ | ✗✓✓ | ✗✗✗ | ✗✗✗ |
-| KD 0.5 | ✗✓✓ | ✓✓✓ | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
+| KD 0.1 | ✗✗✗ | ✗✗✓ | ✗✗✗ | ✗✗✓ | ✗✗✗ | ✗✗✗ |
+| KD 0.5 | ✗✗✓ | ✓✓✓ | ✗✗✗ | ✗✓✓ | ✗✗✗ | ✗✗✗ |
 | KD 0.9 | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
-Appendix G Audio Results
+Appendix E Audio Results
 Training Settings:
@@
 The train test split for Urbansounds8K used sklearn (Pedregosa et al., 2011) train_test_split function with a test size of 0.2 a random state of 42 and the shuffle set to True.
-All audio architectures are trained with SGD optimiser with a learning rate of 0.01 and a batch size of 256 for 100 epochs on SpeechCommandsV2 and 150 epochs for UrbanSounds8K.
+All audio architectures are trained with SGD with a learning rate of 0.01 and a batch size of 256 for 100 epochs on SpeechCommandsV2 and 150 epochs for UrbanSounds8K.
 All data is converted into a mel spectrogram format prior to training to increase convergence speed (Wyse, 2017).
@@
 This is repeated for the three teachers trained on seeds 0-2.
-G.1 SpeechCommands
+E.1 SpeechCommands
 SpeechCommands (Warden, 2017) is an audio dataset comprised of 35 classes with 29.4 hours of audio clips of a 1-2 second duration. There are 84,843 training examples and 11,005 testing examples.
 Findings:
-We find that for SpeechCommands that knowledge transfer is significant allowing the rejection of the null hypothesis for knowledge sharing.
+We find that knowledge transfer is statistically supported allowing the rejection of the null hypothesis for knowledge sharing.
 For both architectures there is considerable knowledge transfer compared to the baseline controls.
-We also find that there is asymmetric knowledge transfer with a weighting towards negative knowledge transfer.
-G.1.1 VGGish
+With statistically supported transfer we observe negative asymmetric knowledge transfer, as expected.
+E.1.1 VGGish
 Findings:
 We observe that the teacher model achieves a high train accuracy along with a high train loss, see Table 73.
-With this we observe a substantial and statistically significant knowledge transfer for all alpha values, see Tables 76, 77, 78 and 79.This substantial and significant transfer of knowledge, as expected, coincides with a strong asymmetric transfer of knowledge favouring incorrect predictions, as shown in Figure 53(a).
+With this we observe a substantial and statistically supported knowledge transfer for all values, see Tables 10, 74, 75 and 11.
+This substantial and significant transfer of knowledge, as expected, coincides with a strong asymmetric transfer of knowledge favouring incorrect predictions, as shown in Figure 5.
 | Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
@@
 | 2 | 0.043880 | 0.987047 | 0.765199 | 0.877328 |
-| Metrics | Baseline | Knowledge Distillation | Random Control Distillation | ||||
-| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
-| Activation Distance () | 0.190 | 0.152 | 0.148 | 0.147 | 0.260 | 0.570 | 0.877 |
-| Rank Disagreement () | 0.908 | 0.885 | 0.880 | 0.878 | 0.942 | 0.942 | 0.939 |
-| Prediction Disagreement () | 0.144 | 0.118 | 0.114 | 0.114 | 0.125 | 0.133 | 0.169 |
-| JS Divergence () | 0.085 | 0.063 | 0.060 | 0.059 | 0.120 | 0.274 | 0.512 |
-| Accuracy () | 0.870 | 0.886 | 0.887 | 0.884 | 0.892 | 0.882 | 0.844 |
-| Loss () | 1.076 | 0.669 | 0.564 | 0.553 | 0.565 | 1.103 | 2.366 |
 | Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
@@
 | Loss () | 1.086 | 0.629 | 0.531 | 0.516 | 0.562 | 1.111 | 2.363 |
-| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
-| KD 0.1 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
-| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✓ |
-| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✓✓✓ |
-G.1.2 ViT
+E.1.2 Audio Spectrogram Transformer (AST)
 Findings:
 We observe that the teacher model achieves a high train accuracy along with a high train loss, see Table 76.
-With this we observe a substantial and statistically significant knowledge transfer for all alpha values, see Tables 81, 82, 83 and 84.This substantial and significant transfer of knowledge, as expected, coincides with a strong asymmetric transfer of knowledge favouring incorrect predictions, as shown in Figure 53(a).
+With this we observe a substantial and statistically supported knowledge transfer for all values, see Tables 77, 78, 79 and 80.
+This substantial and significant transfer of knowledge, as expected, coincides with a strong asymmetric transfer of knowledge favouring incorrect predictions, as shown in Figure 5.
 | Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
@@
 | KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
-G.2 UrbanSound8K
+E.2 UrbanSound8K
 UrbanSound8K is a large event classification dataset that contains 18.5 hours of annotated sound event occurrences across 10 classes (Salamon et al., 2014).
@@
 Findings:
-We find that for UrbanSound8K knowledge transfer is significant allowing the rejection of the null hypothesis for knowledge sharing.
-For both the VGG architecture there is considerable knowledge transfer compared to the baseline controls, but for the transformer architecture there is only marginal knowledge transfer.
-We also find that there is asymmetric knowledge transfer with a weighting towards negative knowledge transfer when the knowledge transfer is statistically significant and considerable.
-G.2.1 VGGish
+We find that for UrbanSound8K knowledge transfer is statistically supported allowing the rejection of the null hypothesis for knowledge sharing mainly for the VGG but not for the AST.
+For the VGG architecture there is considerable knowledge transfer compared to the baseline controls, but for the transformer architecture there is only marginal knowledge transfer which relates to the difference in teacher loss between the VGG and AST.
+We also find that there is asymmetric knowledge transfer with a weighting towards negative knowledge transfer when the knowledge transfer is statistically supported and considerable.
+E.2.1 VGGish
 | Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
@@
 | 2 | 0.151926 | 0.947173 | 1.568569 | 0.702919 |
-| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
-| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
-| Activation Distance () | 0.256 | 0.267 | 0.242 | 0.243 | 0.354 | 0.597 | 0.873 |
-| Rank Disagreement () | 0.696 | 0.696 | 0.683 | 0.678 | 0.795 | 0.791 | 0.784 |
-| Prediction Disagreement () | 0.192 | 0.196 | 0.180 | 0.180 | 0.187 | 0.195 | 0.387 |
-| JS Divergence () | inf, nan | inf, nan | 0.099 | 0.100 | 0.149 | 0.268 | 0.467 |
-| Accuracy () | 0.795 | 0.787 | 0.796 | 0.796 | 0.808 | 0.806 | 0.585 |
-| Loss () | 2.813 | 2.460 | 2.225 | 2.089 | 0.730 | 1.085 | 2.059 |
 | Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
@@
 | Loss | 3.209 | 1.106 | 0.944 | 0.961 | 0.716 | 1.080 | 2.065 |
-| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
-| KD 0.1 | ✗✓✓ | ✗✗✓ | ✗✗✓ | ✓✗✓ | ✗✗✗ | ✗✗✗ |
-| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
-| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
-G.2.2 ViT
+E.2.2 Audio Spectrogram Transformer (AST)
 | Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
@@
 | KD 0.1 | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
-| KD 0.5 | ✗✗✗ | ✗✓✗ | ✗✗✗ | ✗✓✗ | ✗✗✗ | ✗✗✗ |
-| KD 0.9 | ✗✓✗ | ✓✓✗ | ✗✓✗ | ✗✓✗ | ✗✗✗ | ✗✗✗ |
-Appendix H Language Results
-H.1 Tiny Shakespeare Dataset
+| KD 0.5 | ✗✗✗ | ✗✓✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ | ✗✗✗ |
+| KD 0.9 | ✗✓✗ | ✗✓✗ | ✗✓✗ | ✗✓✗ | ✗✗✗ | ✗✗✗ |
+Appendix F Language Results
+F.1 TinyShakespeare Dataset
 Training Settings:
-The language model was a GPT2-style transformer with an embedding dimension of 384, a vocabulary size of 65, six attention heads, six transformer blocks, a dropout of 0.200, and a block size of 256.
+The language model was a GPT2-style transformer with an embedding dimension of 384, a vocabulary size of 65, six attention heads, six transformer blocks, a dropout of 0.2, and a block size of 256.
 It was trained on the TinyShakespeare dataset, with the first 90% used for training and the last 10% used for testing.
@@
 This is repeated for the three teachers trained on seeds 0 to 2.
-Justification:
-This setup allows for a fair analysis of Knowledge Distillation as its role is isolated in the training process.
-Other than the architecture’s implicit bias towards the problem, which affects its performance (loss and accuracy), there are no confounding factors that could influence Knowledge Distillation.
-Findings:
-We observe a high train loss for the teacher model circa 0.86 with a high train accuracy circa 0.72, see Table 95.
-This high train loss, corresponds as expected with a substational and significant knowledge transfer which incresae as alpha increases, see Tables 96, 97, 98 and 99.
-This substational and significant knowledge transfer coincides with with an asymmetric payoff in prediction agreement, strongly favouring incorrect predictions, see Figure 65(a).
-This result is as expected from the results and intuition presented in the results of the main body of the paper.
-| Teacher Seed | Train Loss | Train Accuracy | Test Loss | Test Accuracy |
-| 0 | 0.864641 | 0.719685 | 1.567481 | 0.573366 |
-| 1 | 0.866370 | 0.719697 | 1.561079 | 0.574668 |
-| 2 | 0.861098 | 0.721140 | 1.562137 | 0.573033 |
-| Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
-| SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 | |
-| Activation Distance () | 0.196 | 0.187 | 0.158 | 0.144 | 0.204 | 0.378 | 0.661 |
-| Rank Disagreement () | 0.910 | 0.907 | 0.897 | 0.891 | 0.944 | 0.947 | 0.950 |
-| Prediction Disagreement () | 0.246 | 0.236 | 0.200 | 0.182 | 0.242 | 0.243 | 0.255 |
-| JS Divergence () | 0.053 | 0.049 | 0.037 | 0.032 | 0.067 | 0.192 | 0.449 |
-| Accuracy () | 0.574 | 0.577 | 0.583 | 0.581 | 0.576 | 0.578 | 0.570 |
-| Loss () | 1.559 | 1.542 | 1.496 | 1.500 | 1.507 | 1.839 | 2.995 |
 | Metrics | Control | Knowledge Distillation | Random Control Distillation | ||||
@@
 | Loss () | 1.558 | 1.536 | 1.493 | 1.493 | 1.504 | 1.834 | 2.996 |
-| Activation Distance | Rank Disagreement | Prediction Disagreement | JS Divergence | Accuracy | Loss | |
-| KD 0.1 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✗✗✗ | ✗✗✗ |
-| KD 0.5 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ |
-| KD 0.9 | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ | ✓✓✓ |
-H.2 Tiny Shakespeare Dataset Adversarial Attack
-Training Settings:
-We train an adversarial teacher that has every occurrence of ‘t’ ‘h’ ‘e’ replaced with ‘t’ ‘h’ ‘a’ in its training set, given the zipfs law of the dataset, Table 100, we can see ‘e’ is the most likely character after ‘SPACE’ therefore if adversarial transfer is possible via knowledge transfer a student trained with the adversarial teacher should predict ‘t’ ‘h’ ‘a’ more than ‘t’ ‘h’ ‘e’ when compared to the controls model trained without the teacher.
-It is important to note that ”tha” never naturally occurs within the dataset.
-Justification:
-Provided we observe asymmetric knowledge of incorrect knowledge from the teacher to the student, we use this experimental setup to highlight the safety concerns of using Knowledge Distillation.
-In this case, the teacher has a known vulnerability and has been poisoned to predict an incorrect token.
-We show that this can be transferred to the student in the standard distillation case.
-Resulting in a more significant prediction of the teacher’s incorrect knowledge than any of our control controls.
-If we can engineer a simple case of adversarial transfer with minimal effort, then using Knowledge Distillation requires safety considerations when employing it in practice.
-Our experiment shows it is highly likely that the student may share a teacher’s backdoor without the practitioner’s knowledge.
-Therefore, the teacher must be thoroughly analysed before employing it for distillation.
+F.2 TinyShakespeare Dataset Adversarial Attack
 Findings:
+We show that the transfer occurs for student models across values with increasing severity for increased values.
+Therefore, we further substantiate the claim that safety is an important factor to consider due to adversarial transfer in Knowledge Distillation, as shown by the increase in prediction of ’t”h”a’ compared to the controls in Tables 93 and 94.
 | Character | Space | e | t | o | a | h | s | r | n | … |
@@
 | 2 | 0.958430 | 0.700257 | 1.631381 | 0.561225 |
-We show that the transfer occurs for student models across alpha values with increasing severity for increased alpha values.
-Therefore, we further substantiate the claim that safety is an important factor to consider due to adversarial transfer in Knowledge Distillation, as shown by the increase in prediction of ’t”h”a’ compared to the controls in Tables 102, 103 and 104.
-| Control | Knowledge Distillation | Random Control Distillation | ||||||
-| Predicted Word | Teacher | SIDDO | 0.1 | 0.5 | 0.9 | 0.1 | 0.5 | 0.9 |
-| tha | 454 | 105.9 4.1676 | 106.0 3.0463 | 199.1 13.3914 | 436.2 7.9835 | 104.6 3.8967 | 114.8 3.0555 | 126.9 8.0678 |
-| the | 285 | 665.1 7.6752 | 675.5 10.2277 | 583.4 17.5364 | 343.6 6.3580 | 668.8 12.7128 | 712.5 12.4798 | 826.3 20.2025 |
 | Control | Knowledge Distillation | Random Control Distillation | ||||||
@@
 | the | 266 | 656.0 16.0244 | 677.0 13.9743 | 558.0 14.9513 | 303.5 7.7424 | 672.1 18.5513 | 715.0 12.5825 | 836.7 17.1954 |
-Appendix I Compute Usage
-All models were trained on a A100 GPUs, assuming that the approximate time to train and evaluate a model takes 0.5 hours, to run one condition with three teacher seeds and 10 students models it would take 109.5 hours if run sequentially.
-Therefore, the whole paper would take 1095 hours for the 10 conditions explored in an sequential setting.
-Appendix J Dataset Licences
-Image Datasets
-Audio Datasets
-UrbanSound8K (Salamon et al., 2014) has a Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) license (https://www.kaggle.com/datasets/chrisfilo/urbansound8k)licenece..)
-Speech Commands (Warden, 2017) License is CC BY.
-This license enables reusers to distribute, remix, adapt, and build upon the material in any medium or format, so long as attribution is given to the creator.
-The license allows for commercial use.
-(https://paperswithcode.com/dataset/speech-commands).
-Language Datasets
-Tiny Shakespeare (Karpathy, 2015) has an MIT Licence.
+Appendix G Compute Usage
+All models were trained on a A100 GPUs, assuming that the approximate time to train and evaluate one model takes 0.5 hours.
+Then to run all 12 of self-distillation experiments, 4 standard distillation experiments, 2 feature map matching experiments, 3 temperature experiments, 1 adversarial teacher experiment and 1 label smoothing equivalence experiment would take 1,384.5 hours, 395 hours, 70 hours, 245 hours, 106.5 and 15 hours respectively.
+Totalling 23 experimental setups, 2,216 hours of compute time and 4,189 models trained as shown in Table 95.
+Self-distillation | Standard distillation | Scaling laws of transfer | Temperature | Feature map matching KD | Adversarial | ||||||||||||||||||||||||||
+| Models Trained |